### v1

# PO-Flow: Flow-based Generative Models for Sampling Potential Outcomes and Counterfactuals

Abstract
We propose PO-Flow, a novel continuous normalizing flow (CNF) framework for causal inference that jointly models potential outcomes and counterfactuals. Trained via flow matching, PO-Flow provides a unified framework for individualized potential outcome prediction, counterfactual predictions, and uncertainty-aware density learning. Among generative models, it is the first to enable density learning of potential outcomes without requiring explicit distributional assumptions (e.g., Gaussian mixtures), while also supporting counterfactual prediction conditioned on factual outcomes in general observational datasets. On benchmarks such as ACIC, IHDP, and IBM, it consistently outperforms prior methods across a range of causal inference tasks. Beyond that, PO-Flow succeeds in high-dimensional settings, including counterfactual image generation, demonstrating its broad applicability.
1 Introduction
Predicting potential outcomes (POs) lies at the core of causal inference. In a typical setting, each individual has two potential outcomes: $Y^{(1)}$ if they receive a treatment ($A=1$) and $Y^{(0)}$ if they do not ($A=0$). However, in the real world, only one outcome is observed, while the other potential outcome remains fundamentally unknown [28]. Accurately estimating both POs is crucial for decision-making in domains such as healthcare [41, 27, 50], where clinicians may wish to compare a patient’s likely outcomes under different treatment options [60, 20].
In the causal inference literature, the vast majority of methods focus on estimating the average treatment effect (ATE) and the conditional average treatment effect (CATE) at the population level [32, 25, 1, 22, 37, 42, 55, 67]. It is important to note that although some methods perform well in estimating population-level CATE, they may underperform in individualized potential outcome (PO) prediction, as accurately predicting POs in finite samples is challenging [52].
However, making decisions for individuals based on average treatment effects can be misleading. Therefore, accurately predicting individualized POs and learning their full distributions with uncertainty is a crucial task [69, 35]. Recent works [67, 44, 45] have explored generative models to sample POs and estimate empirical distributions. However, methods that learn the full density of POs without relying on explicit distributional assumptions (e.g., Gaussian mixtures) remain scarce.
In addition, another critical challenge is estimating the counterfactual outcome conditioned on the factual—that is, predicting what would have happened under the alternative treatment, given the observed outcome. This task is especially relevant in post-hoc clinical analyses, where practitioners ask, “Would the patient have fared better on another therapy?” [66, 29, 21].
Counterfactual prediction differs from standard PO estimation—it requires reasoning about outcomes under an alternative treatment conditioned on the observed outcome, introducing dependencies between observed and unobserved outcomes. Structural Causal Models (SCMs) [48] and modern extensions [46, 6, 13, 7, 65, 49, 9] offer the most widely used theoretical framework for counterfactual reasoning. SCMs methods typically assume known causal structures or aim to learn them explicitly. Yet, it may be possible to estimate counterfactuals directly without recovering the SCM [70]. In PO-Flow, we do not explicitly assume an SCM, but seek to directly estimate counterfactual outcomes.
We aim to develop a generative modeling framework that jointly supports population-level CATE estimation, individualized PO prediction with density modeling, and counterfactual outcome prediction without explicitly modeling an SCM—an area that has been relatively less explored.
1.1 Related Works
CATE Estimation: Numerous methods have been proposed for CATE estimation [32, 25, 1, 22, 37, 42, 55, 67]. Broadly, these fall into model-agnostic and model-specific approaches. Model-agnostic estimators, or meta-learners [37, 15, 16, 2, 54], typically follow one of two strategies [43]: (a) one-step methods, which fit regression models and compute CATE as the difference in predicted outcomes, and (b) two-step methods, which estimate nuisance functions to construct pseudo-outcomes, then regress them on covariates.
In contrast, model-specific estimators adapt machine learning techniques for CATE estimation, including methods such as [14, 55, 42, 34, 67]. More recently, neural networks have been widely applied to this task, with notable examples in [55, 42, 67]. However, most of these approaches mainly focus on estimating CATE rather than predicting individualized potential outcomes.
Individualized POs Prediction: Although some CATE methods can be adapted for potential outcomes (POs) prediction, they may perform suboptimally, particularly on finite samples, as they are not specifically optimized for accurate PO estimation. This limitation partly arises because many methods assume that CATE has a simpler functional structure than POs. Nevertheless, several CATE approaches are also designed with POs prediction as the target [37, 55, 67, 42, 45, 43], and these will serve as our primary baselines for experimental comparisons.
Counterfactual Outcomes Prediction: Structural Causal Models (SCMs) provide a well-established framework for counterfactual reasoning [48, 46, 6, 13, 7, 65, 49, 9], but practical applications often require specifying the full causal graph and structural equations, which may be difficult or infeasible for real-world data. Some recent works avoid explicitly modeling SCMs [39, 5, 44, 62], but typically address specific tasks and rely on additional information, such as temporal structure.
Some works [27, 34, 42, 38, 67] targeting CATE estimation describe their goal as counterfactual inference. However, in these works, "counterfactual" typically refers to the unobserved potential outcome under an alternative treatment, without conditioning on the observed factual outcome.
While SCMs offer theory-grounded counterfactual reasoning, general-purpose methods for individualized counterfactual prediction—without explicitly modeling SCMs or relying on auxiliary information—remain scarce. Although [67, 43] were not originally designed for counterfactuals, we adapt them by incorporating factual information into intermediate steps and use them as baselines.
Diffusion and Flow Models: Diffusion models [26, 57, 58] and flow-based generative models [12, 64, 40, 4] have demonstrated strong performance across a range of applications, including image generation, statistical sampling [18, 59, 61], and physics-informed modeling [8, 68, 51].
Recently, diffusion and flow-based generative models have been applied to causal inference tasks such as counterfactual image generation [53, 19], causal query answering [10], and inference over causal graphs or structural causal models [9, 33].
However, to our knowledge, only two generative algorithms have been proposed specifically for the potential outcomes (PO) task: the diffusion-based DiffPO [43] and the discrete normalizing flow-based INFs [45]. We conduct extensive comparisons with both methods in this work.
1.2 Main Contributions
We propose PO-Flow, a continuous normalizing flow (CNF) framework for potential outcomes and counterfactuals. Our main contributions are: (1) PO-Flow provides a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction—without explicitly modeling structural causal models (SCMs) or relying on auxiliary information for the counterfactual task; (2) PO-Flow enables learning the density of predicted potential outcomes, offering deeper insights into uncertainty; (3) PO-Flow outperforms state-of-the-art methods across a range of causal tasks on multiple benchmarks.
2 Preliminaries
2.1 Neural ODE and Continuous Normalizing Flow:
A Neural ODE is a continuous model where the trajectory of data is modeled as the solution of an ordinary differential equation (ODE). Formally, in $\mathbb{R}^{d}$, given an input $y_{0}=y(0)$ at time $t=0$, the transformation to the output $y_{1}=y(1)$ at time $t=1$ is governed by:
$\frac{dy(t)}{dt}=v_{\theta}(y(t),t),\quad t\in[0,1],$ (1)
where $v_{\theta}:\mathbb{R}^{d_{Y}+1}\to\mathbb{R}^{d_{Y}}$ represents the velocity field, which is parameterized by a neural network with parameter $\theta$. The time horizon is rescaled to $[0,1]$ without loss of generality.
Continuous Normalizing Flow (CNF) is a class of normalizing flows in which the transformation of a probability density is governed by a time-continuous Neural ODE. The marginal density of $y(t)$, denoted as $p(y,t)$, evolves according to the continuity equation implied by the ODE in Eq. (1). This continuity equation is given by:
$\partial_{t}p(y,t)+\nabla\cdot(p(y,t)v_{\theta}(y,t))=0,\quad t\in[0,1],$ (2)
[…]
When the Neural ODE is well-posed, it defines a continuous and invertible mapping from the initial sample $y_{0}$ to the terminal sample $y_{1}$. The inverse mapping from $y_{1}$ to $y_{0}$ can be computed by integrating the ODE (1) in reverse time. This framework allows setting $y_{0}\sim p(\cdot,0)$ as the data distribution and $y_{1}\sim p(\cdot,1)$ as the base (noise) distribution, typically chosen as $N(0,I)$.
For notational convenience, we define $p_{Y}(\cdot):=p(\cdot,0)$ as the data distribution and $q(\cdot):=p(\cdot,1)$ as the base (noise) distribution. Throughout this paper, we adopt the common choice $q(\cdot)=N(0,I)$.
2.2 Flow Matching
Flow Matching (FM) [40, 4] provides a simulation-free approach for training continuous normalizing flows by directly regressing the model velocity $v_{\theta}(y,t)$ onto a predefined target velocity field $u(y,t)$. The original FM objective is defined as:
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],y\sim p(\cdot,t)}\left[\|v_{\theta}(y,t)-u(y,t)\|^{2}\right],$ (3)
where $u(y,t)$ is an analytically specified target velocity field.
Linear Interpolant: Due to the intractability of $u(y,t)$ without conditioning on the initial data $y_{0}$, Conditional Flow Matching (CFM) is introduced. Conditioned on an observed sample $y_{0}\sim p_{Y}(\cdot)$, CFM employs a pre-specified interpolation function $I_{t}$ between two endpoints $y_{0}$ and $y_{1}\sim q(\cdot)$, where $q$ is typically $N(0,I)$. The interpolated path is defined as:
$\phi_{t}:=I_{t}(y_{0},y_{1}),\quad t\in[0,1].$ (4)
Here, $I_{t}$ can be analytically specified. Conditioned on $y_{0}$, the optimal and most information-preserving choice is the linear interpolation path [40]:
$\phi_{t}=(1-t)y_{0}+(t+\sigma_{\text{min}}(1-t))y_{1},$ (5)
where $y_{1}\sim N(0,I)$ and $\sigma_{\text{min}}$ is a small positive hyperparameter that ensures $p(\phi_{t},0)\sim N(y_{0},\sigma_{\text{min}}^{2}I)$. When $\sigma_{\text{min}}=0$, this reduces to the strict linear interpolant $\phi_{t}=(1-t)y_{0}+ty_{1}$.
Here, $\phi_{t}$ serves as a pre-specified reference trajectory that the ODE solution $y(t)$ should follow at time $t$. Under the linear interpolant defined in (5), the corresponding reference velocity field becomes:
$\frac{d\phi_{t}}{dt}=(1-\sigma_{\text{min}})y_{1}-y_{0}.$ (6)
Training loss: CFM trains the flow $v_{\theta}(y(t),t)$ by directly regressing it onto the reference velocity field (6). The training objective is given by:
$\mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],y_{0}\sim p_{Y}(\cdot),y_{1}\sim q(\cdot)}\|v_{\theta}(\phi_{t},t)-\frac{d\phi_{t}}{dt}\|^{2}.$ (7)
3 Problem Settings
Notations and Setups: We consider a setting where each individual is described by covariates $X\in\mathbb{R}^{d_{X}}$, a treatment assignment $A\in\{0,1\}$, and an outcome $Y^{(A)}\in\mathbb{R}^{d_{Y}}$ under treatment $A$. The observational dataset is $\mathcal{D}=\{(y^{(a)}_{i},x_{i},a_{i})\}_{i=1}^{n}\sim p_{Y,X,A}(\cdot)$, where $y_{i}^{(a)}$ is the observed potential outcome under treatment $a_{i}$, and $(x_{i},a_{i})\sim p_{X,A}(\cdot)$ are the associated covariates and treatment.
Potential Outcomes: In the potential outcomes framework, each individual has two potential outcomes: $Y^{(0)}$ (control) and $Y^{(1)}$ (treatment), but only one is observed based on assignment $A$. We define potential outcome prediction as estimating $Y^{(a)}\mid X,A=a$. For convenience, we denote $Y^{(a)}$ and $\hat{Y}^{(a)}$ as the true and predicted potential outcomes.
Counterfactual Outcomes: In contrast to potential outcomes, we define the counterfactual outcome as $Y_{\text{cf}}^{(1-a)}:=Y^{(1-a)}\mid X,A=a,Y^{(a)}$: the outcome had the individual received the opposite treatment, conditioned on its observed potential (factual) outcome. For convenience, we denote $Y_{\text{cf}}^{(1-a)}$ and $\hat{Y}_{\text{cf}}^{(1-a)}$ as the true and predicted counterfactual outcomes, respectively.
To ensure the identifiability of potential outcomes, we adopt the standard assumptions commonly used in the literature [43, 45, 67]:
Assumption 3.1.
(1) Consistency: If an individual receives treatment $A=a$, the observed outcome equals the corresponding potential outcome: $Y=Y^{(a)}$. (2) Unconfoundedness: All confounders are observed, i.e., $\{Y^{(0)},Y^{(1)}\}\perp\!\!\!\perp A\mid X$. (3) Overlap: Treatment assignment is probabilistic: $10<P(A=1\mid X=x)<$ 1, for all $x\in\mathcal{X}$.
Counterfactuals are rarely identifiable without strong assumptions [63, 47]. Estimating them can often be framed as an optimal transport problem [11, 17]. To constrain the space of joint transport mappings, our PO-Flow framework introduces the following assumption, applicable only to Section 4.2:
Assumption 3.2.
There exists a deterministic and invertible mapping $f:\mathbb{R}^{d_{Y}+d_{X}}\times\{0,1\}\rightarrow\mathbb{R}^{d_{Z}}$ that encodes the observed outcome as $Z_{Y}=f(Y^{(A)},X,A)$. A counterfactual latent representation can be obtained as $Z^{\prime}_{Y}=Z_{Y}+\delta$, where $\delta\sim\mathcal{N}(0,\Sigma(X))$ is an optional covariate-dependent noise term. The counterfactual is then recovered via $Y_{\text{cf}}^{(1-A)}=f^{-1}(Z^{\prime}_{Y},X,1-A)$.
4 PO-Flow: A Continuous Normalizing Flow for Potential Outcomes
We introduce PO-Flow, a continuous normalizing flow framework for modeling potential and counterfactual outcomes, with three main goals: 1. predicting individualized potential outcomes, 2. estimating counterfactual outcomes, and 3. learning the densities.
Our PO-Flow framework is built on Continuous Normalizing Flows (CNFs) and is trained using the Flow Matching (FM) objective. Our learning target is the trajectory of the sample $y(t)$ over $t\in[0,1]$. To condition the dynamics on both covariates $x$ and treatment $a$, our Neural ODE is defined as:
$\small\frac{dy(t)}{dt}=v_{\theta}(y(t),t,x,a),\quad t\in[0,1],$ (8)
where $v_{\theta}:\mathbb{R}^{d_{Y}+d_{X}+2}\rightarrow\mathbb{R}^{d_{Y}}$ is parametrized by a neural network.
Notation: Given a data point $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$, we define $y_{0}=y^{(a)}$ as the observed training sample at time $t=0$ in the Neural ODE, and $y_{1}\sim q(\cdot)$ as the noise sample at time $t=1$. Recall that we denote $\hat{y}^{(a)}$ and $\hat{y}_{\text{cf}}^{(1-a)}$ as the predicted potential and counterfactual outcomes, respectively.
Training loss: PO-Flow adopts the Conditional Flow Matching (CFM) objective for training. Under the updated definitions, our training loss retains the form of Equation (7) and is given by:
$\min_{\theta}\ \mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot),q(\cdot)}\|v_{\theta}(\phi_{t},t,x,a)-\frac{d\phi_{t}}{dt}\|^{2},$ (9)
where $\phi_{t}$ and $\frac{d\phi_{t}}{dt}$ are defined in (5) and (6), respectively.
Remark: Following prior work [23, 43], one may estimate propensity scores $w(x)=p(A=1\mid X=x)$ using a pre-trained classifier and apply inverse probability weighting $\frac{1}{w(x)}$ in the loss. However, we found this did not improve performance, but both implementations are included in our code.
Forward Process: Given the velocity field $v_{\theta}$, the forward process refers to pushing a factual outcome $y^{(a)}$ (at $t=0$) to a latent embedding $z_{y}$ (at $t=1$) via:
[…]
Through the forward process, $z_{y}$ encodes information from the factual outcome $y^{(a)}$ as well as its associated covariates $x$ and treatment $a$.
Reverse Process: The reverse process is initialized differently for interventional and counterfactual outcome prediction, as detailed in Sections 4.1 and 4.2. In general, given a noise sample $y_{1}\sim q(\cdot)$ at $t=1$, the reverse process is defined as:
[…]
Empirically, we use the Runge-Kutta method for numerical integration, with implementation details provided in Appendix B.4.
During training, we construct synthetic paths $\phi_{t}$ between $y_{0}$ and $y_{1}\sim q(\cdot)$ using the CFM loss, conditioning on $y_{0}$. However, Proposition 4.1 shows that the gradient of the CFM loss with respect to the model parameters is equivalent to the gradient of the original FM loss.
Proposition 4.1.
Assume that $p_{t}(y|x,a)>0$ for all $y\in\mathbb{R}^{d_{Y}}$ and $t\in[0,1]$. Then, up to a constant independent of $\theta$, the Conditional Flow Matching (CFM) loss and the original Flow Matching (FM) loss are equivalent. Hence,
$\nabla_{\theta}\mathcal{L}_{\text{FM}}(\theta)=\nabla_{\theta}\mathcal{L}_{\text{CFM}}(\theta).$
As a result of Proposition 4.1, the learned velocity field $v_{\theta}$ during the reverse process generates samples $\hat{y}_{0}$ independent of the individual samples $y_{0}$ used during training. Consequently, since $y_{1}\sim q(\cdot)$ is random, $v_{\theta}$ induces a distribution over outcomes conditioned on $(x,a)$.
4.1 PO-Flow for Potential Outcomes Prediction
  Require: Trained $v_{\theta}$
  Input: $(x,a)\sim p_{X,A}(\cdot)$
  1: $y_{1}\sim q(\cdot)$
[…]
  Return $\hat{y}^{(a)}:=\hat{y}_{0}$
Potential Outcomes Prediction
Algorithm 1
Potential Outcomes Prediction
The potential outcomes prediction in PO-Flow follows a simple reverse process, as defined in Equation (11). Starting from a noise sample $y_{1}\sim q(\cdot)$, we integrate the velocity field $v_{\theta}$ backward in time, conditioned on covariates $x$ and treatment $a$, to obtain a point estimate of the potential outcome $\hat{y}^{(a)}$. The procedure is summarized in Algorithm 1.
In addition to predicting point estimates of potential outcomes, PO-Flow is capable of learning the probability densities of the potential outcomes, as detailed in Section 4.3.
4.2 PO-Flow for Counterfactual Outcomes Prediction
  Require: Trained $v_{\theta}$
  Input: $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$
  1. Forward process (encoding):
[…]
         with $y(0)=y^{(a)}$.
  3. Uncertainty-aware perturbations:
  4.     $\delta\sim N(0,\Sigma(x))$
  5.     $z^{\prime}_{y}=z_{y}+\delta$
  6. Reverse process (decoding):
[…]
         with $y(1)=z^{\prime}_{y}$.
  Return: $\hat{y}_{\text{cf}}^{(1-a)}:=\hat{y}_{0}$
Counterfactual Prediction
Algorithm 2
Counterfactual Prediction
Under Assumption 3.2, the counterfactual outcome prediction in PO-Flow consists of both a forward and a reverse process. Given a factual observation $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$, the forward process (Lines 2–3 in Algorithm 2) encodes the information $(y^{(a)},x,a)$ into a latent representation centered at $z_{y}$. Lines 4–6 introduce uncertainty-aware perturbations to $z_{y}$, reflecting covariate-dependent uncertainty. Finally, the reverse process (Lines 7–8) decodes the latent $z_{y}^{\prime}$ back into the counterfactual outcome prediction $\hat{y}_{\text{cf}}^{(1-a)}$.
Figure 1: Illustration of counterfactual predictions. Left: two potential outcomes distribution. Right: Base (noise) distribution.
Because of the smooth and information-preserving nature of the PO-Flow path trained via flow matching, the encoded $z_{y}$ retains rich information from the factual outcome $y^{(a)}$, enabling accurate decoding of the counterfactual $\hat{y}_{\text{cf}}^{(1-a)}$. To our knowledge, this is the first work to explore counterfactual outcome prediction using flow or diffusion-based models.
As illustrated in Figure 1, diffusion models rely on noisy and stochastic forward processes, which degrade the information in $z_{y}$ and lead to reversed samples that resemble interventional potential outcomes rather than counterfactuals. In Section 5.4, we present counterfactuals using an adapted implementation of DiffPO [43] (originally not designed for counterfactual), compared with PO-Flow.
4.3 PO-Flow for Learning Potential Outcomes Densities
Another distinctive advantage of PO-Flow is its ability to learn the density of the Potential Outcomes (POs). We define $\hat{p}_{Y}(\cdot\mid x,a,y_{1})$ as the PO density learned by PO-Flow, conditioned on $(x,a)$ and a base sample $y_{1}$. The explicit form of the learned PO density is presented in Proposition 4.2, with the proof provided in Appendix A.
Proposition 4.2.
Given a noise sample $y_{1}\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=\log q(y_{1})+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)ds.$ (12)
Empirically, the divergence of the velocity field can be computed under the Hutchinson trace estimator (e.g., [31, 64]). More details are provided in Appendix B.5.
$\small\nabla\cdot v_{\theta}(y(t),t,x,a)\approx\mathbb{E}_{p(\epsilon)}\left[\epsilon\cdot\frac{v_{\theta}(y(t)+\sigma\epsilon,t,x,a)-v_{\theta}(y(t),t,x,a)}{\sigma}\right],$ (13)
where $p(\epsilon)$ is a distribution in $\mathbb{R}^{d_{Y}}$ satisfying $\mathbb{E}[\epsilon]=0$ and Cov$(\epsilon)=I$ (e.g., a standard Gaussian).
To our knowledge, PO-Flow is the first generative approach in causal inference to explicitly learn the probability density of predicted potential outcomes. Unlike prior works [44, 45] that generate PO samples and estimate empirical distributions, PO-Flow provides direct access to likelihoods, enabling deeper insights into uncertainty.
This density estimation allows PO-Flow to assess prediction confidence via log-likelihood, identify unlikely outcomes, and quantify uncertainty. For counterfactual prediction, one can perturb the latent code $z_{y}^{\prime}$ multiple times and select the most confident $\hat{y}_{\text{cf}}^{(1-a)}$ based on log-likelihood, effectively reducing variance while improving robustness.
5 Experiments
Baselines: Our baselines cover several categories: model-agnostic learners that learn the conditional distribution means (S-learner and T-learner [37]), representation-based conditional regressors (Counterfactual Regression (CFR) [55]), generative approaches (Causal Effect Variational Autoencoder (CEVAE) [42]; GAN for Individualized Treatment Effect (GANITE) [67]), and most recent generative approaches based on diffusion (DiffPO [43]) and discrete normalizing flows (INFs [45]).
Datasets: Counterfactual outcomes and unobserved potential outcomes (POs) are inaccessible in real-world data but can be simulated in semi-synthetic settings, as discussed in detail in Appendix B.1. Accordingly, we evaluate PO-Flow on standard synthetic benchmarks, including ACIC 2016, ACIC 2018, IHDP, and IBM.
The code and datasets are available at https://github.com/StatFusion/PO-Flow.
5.1 Prediction of Potential Outcomes
Table 1: Root mean squared error (RMSE) of estimated potential outcomes (POs) across different methods on the ACIC 2018, IHDP, and IBM datasets. Results are averaged over 10-fold cross-validation.
ACIC 2018 IHDP IBM
$\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$
[…]
DiffPO $0.72_{\pm.14}$ $0.79_{\pm.16}$ $1.33_{\pm.20}$ $1.40_{\pm.22}$ $1.82_{\pm.38}$ $1.85_{\pm.35}$
INFs $0.69_{\pm.16}$ $0.78_{\pm.17}$ $1.02_{\pm.10}$ $\mathbf{1.20_{\pm.12}}$ $1.52_{\pm.24}$ $1.57_{\pm.28}$
S-learner $\mathbf{0.54_{\pm.09}}$ $\mathbf{0.57_{\pm.11}}$ $1.31_{\pm.18}$ $1.44_{\pm.20}$ $1.01_{\pm.17}$ $1.16_{\pm.19}$
T-learner $1.57_{\pm.32}$ $1.71_{\pm.40}$ $1.45_{\pm.25}$ $1.49_{\pm.27}$ $1.89_{\pm.47}$ $1.96_{\pm.50}$
CEVAE $0.83_{\pm.17}$ $0.85_{\pm.17}$ $1.18_{\pm.15}$ $1.36_{\pm.18}$ $0.96_{\pm.12}$ $\mathbf{0.98_{\pm.12}}$
CFR $0.94_{\pm.10}$ $0.98_{\pm.12}$ $1.10_{\pm{.20}}$ $\mathbf{1.18_{\pm{.20}}}$ $2.09_{\pm.45}$ $2.17_{\pm.48}$
GANITE $0.88_{\pm.15}$ $0.97_{\pm.15}$ $1.60_{\pm.36}$ $1.67_{\pm.34}$ $2.48_{\pm.39}$ $2.59_{\pm.47}$
Evaluation Metrics: We evaluate the point estimates of potential outcomes (POs) using the root mean squared error (RMSE) between the estimated outcomes $\hat{y}_{i}$ and the true outcomes $y_{i}$, computed as $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$. $\text{RMSE}_{\text{in}}$ and $\text{RMSE}_{\text{out}}$ refer to the RMSE evaluated on the training and test sets, respectively.
Results: Table 1 shows that PO-Flow achieves the most accurate prediction of potential outcomes on the majority of datasets.
5.2 Estimation of CATEs
Conditional Average Treatment Effect (CATE): The CATE is defined as $\tau(x)=\mathbb{E}[y_{1}-y_{0}\mid X=x]$, representing the expected treatment effect conditioned on covariates $x$. It captures individual-level heterogeneity in causal effects.
Evaluation Metrics: We evaluate CATE using the precision in estimation of heterogeneous effect (PEHE), defined as the mean squared error between estimated and true effects: $\epsilon_{\text{PEHE}}=\frac{1}{n}\sum_{i=1}^{n}(\hat{\tau}(x_{i})-\tau(x_{i}))^{2}$. Table 5 reports the root PEHE for both training and test sets. PO-Flow achieves the most accurate CATE estimates on most datasets.
Table 2: Root Precision in Estimation of Heterogeneous Effect ($\sqrt{\epsilon_{\text{PEHE}}}$) across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets. Results are averaged over 10-fold cross-validation.
ACIC 2016 ACIC 2018 IHDP IBM
$\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$
[…]
DiffPO $0.91_{\pm.20}$ $1.12_{\pm.26}$ $0.07_{\pm.03}$ $\mathbf{0.08_{\pm.04}}$ $0.79_{\pm.16}$ $0.84_{\pm.18}$ $0.08_{\pm.03}$ $0.09_{\pm.03}$
INFs $1.37_{\pm.28}$ $1.42_{\pm.30}$ $0.33_{\pm.11}$ $0.37_{\pm.12}$ $0.90_{\pm.27}$ $0.95_{\pm.27}$ $0.16_{\pm.08}$ $0.18_{\pm.08}$
S-learner $3.81_{\pm.76}$ $3.87_{\pm.98}$ $\mathbf{0.06_{\pm.02}}$ $\mathbf{0.07_{\pm.03}}$ $0.92_{\pm.07}$ $0.93_{\pm.08}$ $0.06_{\pm.02}$ $0.07_{\pm.04}$
T-learner $5.02_{\pm.71}$ $5.15_{\pm.87}$ $1.21_{\pm.47}$ $1.32_{\pm.53}$ $1.24_{\pm.08}$ $1.31_{\pm.09}$ $0.60_{\pm.14}$ $0.69_{\pm.15}$
CEVAE $2.06_{\pm.61}$ $2.31_{\pm.74}$ $1.19_{\pm.15}$ $1.21_{\pm.16}$ $1.41_{\pm.10}$ $1.43_{\pm.10}$ $0.73_{\pm.15}$ $0.75_{\pm.16}$
CFR $2.80_{\pm.58}$ $2.74_{\pm.60}$ $0.77_{\pm.18}$ $0.80_{\pm.18}$ $0.73_{\pm.04}$ $0.78_{\pm.04}$ $1.05_{\pm.27}$ $1.08_{\pm.28}$
GANITE $3.02_{\pm.76}$ $3.11_{\pm.74}$ $0.55_{\pm.11}$ $0.60_{\pm.13}$ $1.90_{\pm.40}$ $2.40_{\pm.40}$ $0.77_{\pm.20}$ $0.80_{\pm.24}$
5.3 Learning the Full Potential Outcome Distributions
Beyond point estimation, PO-Flow offers two distinct advantages for modeling potential outcomes: (1) it enables individualized sampling from the conditional distribution $P(Y\mid X,A)$, generating a rich set of plausible outcomes for each individual; and (2) it supports learning the probability density of any generated outcome, providing deeper insights into predictive uncertainty.
5.3.1 Learning potential outcome densities
PO-Flow enables direct computation of potential outcome densities, as shown in Proposition 4.2. To our knowledge, it is the first generative method in causal potential outcomes to provide this capability.
Evaluation Metrics: To evaluate the learned densities, we compute the KL divergence between the PO-Flow estimate $\hat{p}(\cdot|x,a)$ and the ground-truth density $p_{\text{true}}(\cdot|x,a)$. Given that the semi-synthetic datasets assume $y^{(a)}\sim\mathcal{N}(\mu(a),1)$, the individual-level KL is $\text{KL}(\hat{Y}_{X}^{(a)}\|Y_{X}^{(a)})=\mathbb{E}_{\hat{y}\sim\hat{p}}\left[\log\frac{\hat{p}(\hat{y}|x,a)}{\mathcal{N}(\mu(\hat{y}),1)}\right]$, where $\hat{p}$ is computed via (12). Table 3 reports the average results.
Table 3: KL divergence of learned potential outcome distributions across datasets. Results are averaged over 10-fold cross-validation.
ACIC 2016 ACIC 2018 IHDP IBM
PO-Flow $\mathbf{0.14_{\pm.04}}$ $\mathbf{0.08_{\pm.02}}$ $\mathbf{0.09_{\pm.03}}$ $\mathbf{0.08_{\pm.03}}$
GP-ITE $1.73_{\pm.34}$ $0.73_{\pm.19}$ $1.42_{\pm.30}$ $0.68_{\pm.17}$
Baselines: A natural baseline with density estimation capability is Gaussian Processes for interventional treatment effects, as used in [3, 30].
Results: PO-Flow achieves much lower KL divergence than GP baselines on the ACIC 2018, IHDP, and IBM datasets, indicating more accurate density estimation of potential outcomes.
This property allows PO-Flow to assess prediction confidence via log-density and to identify unlikely outcomes. For counterfactual estimation, one can sample multiple candidates from a perturbed latent region and select the most likely one based on the model’s estimated density.
5.3.2 Empirical distribution estimations
Evaluation Metrics: To enable comparison of estimated potential outcome distributions across methods, we include empirical distribution estimates for all baselines. These are evaluated using the Wasserstein-1 distance, defined as: $W_{1}(\hat{p}(y|x,a),p(y|x,a))=\inf_{\gamma\in\Gamma(\hat{p},p)}\mathbb{E}_{(\hat{y},y\sim\gamma)}[\|\hat{y}-y\|_{1}]$. The results are shown in Table 4.
Table 4: In-sample and out-of-sample Wasserstein distances across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets. Results are averaged over 10-fold cross-validation.
ACIC 2016 ACIC 2018 IHDP IBM
$\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$
[…]
DiffPO $0.47_{\pm.17}$ $0.51_{\pm.18}$ $\mathbf{0.04_{\pm.02}}$ $\mathbf{0.13_{\pm.05}}$ $0.80_{\pm.15}$ $0.88_{\pm.17}$ $0.69_{\pm.14}$ $0.75_{\pm.14}$
INFs $0.59_{\pm.13}$ $0.67_{\pm.14}$ $0.08_{\pm.03}$ $0.16_{\pm.05}$ $0.32_{\pm.08}$ $\mathbf{0.40_{\pm.09}}$ $0.36_{\pm.08}$ $\mathbf{0.45_{\pm.09}}$
S-learner $1.01_{\pm.13}$ $1.02_{\pm.20}$ $0.35_{\pm.13}$ $0.49_{\pm.20}$ $0.71_{\pm.14}$ $0.85_{\pm.19}$ $0.46_{\pm.11}$ $0.55_{\pm.14}$
T-learner $1.22_{\pm.15}$ $1.48_{\pm.18}$ $1.04_{\pm.13}$ $1.00_{\pm.30}$ $0.83_{\pm.20}$ $0.92_{\pm.28}$ $1.21_{\pm.23}$ $1.67_{\pm.32}$
CEVAE $0.62_{\pm.17}$ $0.70_{\pm.21}$ $0.88_{\pm.18}$ $1.02_{\pm.23}$ $0.65_{\pm.14}$ $0.76_{\pm.20}$ $0.38_{\pm.07}$ $\mathbf{0.45_{\pm.07}}$
CFR $0.68_{\pm.13}$ $0.73_{\pm.15}$ $0.88_{\pm.12}$ $0.90_{\pm.15}$ $0.50_{\pm.17}$ $0.56_{\pm.19}$ $0.92_{\pm.20}$ $0.93_{\pm.21}$
GANITE $0.74_{\pm.20}$ $0.86_{\pm.23}$ $0.89_{\pm.24}$ $1.18_{\pm.25}$ $0.71_{\pm.23}$ $0.82_{\pm.29}$ $1.09_{\pm.27}$ $1.34_{\pm.30}$
Figure 2: Comparison between empirically estimated and the ground-truth POs distributions. Left: $P(y|x,a=0)$ on IHDP. Right: $P(y|x,a=1)$ on IHDP.
Understanding the distribution of potential outcomes (POs) is essential for reasoning about uncertainty in treatment decisions. As shown in Figure 2, PO distributions can shift differently across treatments for the same individual. PO-Flow provides not only point estimates but also distributional insight, supporting uncertainty-aware analysis. Section 5.3.1 further demonstrates its ability to estimate densities accurately. This enables practitioners to assess whether an outcome lies within a desired range and evaluate its likelihood (Proposition 4.2) under a given treatment, offering a clearer sense of confidence in the effect.
5.4 Counterfactual Outcome Estimation
Evaluation Metrics: We evaluate the predicted counterfactual outcomes using the root mean squared error (RMSE) between the estimated counterfactuals $\hat{y}_{i,\text{cf}}^{(1-a)}$ and the ground-truth counterfactuals $y_{i,\text{cf}}^{(1-a)}$, computed as $\text{RMSE}=\small\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i,\text{cf}}^{(1-a)}-\hat{y}_{i,\text{cf}}^{(1-a)})^{2}}$. The results are presented in Table 5. Here, $\text{RMSE}_{\text{in}}$ and $\text{RMSE}_{\text{out}}$ denote the RMSE evaluated on the training and test sets, respectively.
Baselines: To our knowledge, very few existing methods are specifically designed for counterfactual outcome prediction conditioned on factuals in general settings—without relying on structural causal models or temporal information. While methods like CFR include “counterfactual” in name, they target unobserved potential outcomes rather than factual-conditioned ones, though we still include them as valuable baselines. We also adapt DiffPO and GANITE for this task, as detailed in Appendix B.6.
Table 5: Root mean squared error (RMSE) of estimated counterfactual outcomes across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets. Results are averaged over 10-fold cross-validation.
ACIC 2016 ACIC 2018 IHDP IBM
$\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$
[…]
DiffPO $1.77_{\pm.20}$ $1.78_{\pm.20}$ $0.62_{\pm.14}$ $0.65_{\pm.15}$ $1.81_{\pm.22}$ $1.85_{\pm.23}$ $0.62_{\pm.13}$ $0.64_{\pm.16}$
CFR $3.15_{\pm.28}$ $3.38_{\pm.30}$ $0.99_{\pm.05}$ $1.01_{\pm.07}$ $\mathbf{1.63_{\pm.21}}$ $\mathbf{1.65_{\pm.25}}$ $0.86_{\pm.18}$ $0.87_{\pm.16}$
GANITE $3.41_{\pm.52}$ $3.67_{\pm.55}$ $0.68_{\pm.13}$ $0.70_{\pm.14}$ $2.48_{\pm.53}$ $2.72_{\pm.59}$ $1.84_{\pm.30}$ $2.02_{\pm.37}$
5.5 Counterfactual Image Generation
Figure 3: PO-Flow counterfactual image generation from female to male faces. $(a)$: Each row shows a factual outcome $y^{(a)}$ (1st column), its encoded latent $z_{y}$ (2nd column), and reverse-time decoding at $t=2/3$, $t=1/3$, and $t=0$ (3rd–5th columns), yielding the counterfactual $\hat{y}_{\text{cf}}^{(1-a)}$. $(b)$: Counterfactuals decoded from perturbed latents $z_{y}+\delta$, with $\delta\sim\mathcal{N}(0,0.2I)$.
Beyond low-dimensional datasets, PO-Flow also succeeds on high-dimensional counterfactual image generation. We train PO-Flow on the latent space of a pre-trained VAE using the CelebA dataset, predicting counterfactual male faces from original female faces. In this setting, $X\in\mathbb{R}^{6}$ encodes visual attributes (hair, eyebrows, bangs, lips), $a\in\{\text{male},\text{female}\}$ denotes gender, and $y\in\mathbb{R}^{512}$ is the VAE latent feature. Visual results are shown in Figures 3 and 6; detailed settings are in Appendix B.2.
5.6 Discussions and Comparisons
Training efficiency: PO-Flow exhibits high training efficiency, converging in approximately 300 gradient steps across multiple causal datasets, as shown in Appendix C.1.
Minimal architecture sensitivity: We find that PO-Flow is the least sensitive to neural network architecture and hyperparameters. Using the same NN structure across all causal datasets, PO-Flow consistently achieves state-of-the-art performance. In contrast, we find that generative approaches including DiffPO, INFs, and GANITE require careful hyperparameter tuning for each dataset.
Unique density estimation: Our continuous normalizing flow (CNF) backbone enables density learning of potential outcomes, as shown in Proposition 4.2 and Section 5.3.1, which, to our knowledge, is the first generative model in the field to achieve this.
Counterfactual: We take a step further by exploring counterfactual outcomes, enabled by the smooth PO-Flow mapping that preserves factual information during encoding. Unlike diffusion-based DiffPO, which has noisy forward processes, PO-Flow offers information-preserving forward processes, as shown in Figure 1. To our knowledge, it is among the first to explore empirical counterfactual prediction without relying on temporal data or explicitly modeling structural causal models.
However, one area for improving counterfactual prediction lies in modeling uncertainty more expressively. The current covariate-conditioned Gaussian perturbations may not fully capture counterfactual variation. Future work could explore more theoretically grounded counterfactual uncertainty.
6 Conclusions
We introduced PO-Flow, a continuous normalizing flow framework for modeling potential and counterfactual outcomes in causal inference. PO-Flow provides a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction—without explicitly modeling structural causal models (SCMs) or relying on auxiliary information for the counterfactual task.
Empirically, PO-Flow achieves state-of-the-art performance on benchmark datasets across multiple causal tasks. Among generative models, it is the first to accurately learn the density of potential outcomes, enabling principled uncertainty quantification. Furthermore, PO-Flow is efficient, robust, and extendable to high-dimensional settings such as counterfactual image generation. These results highlight PO-Flow as a unified and scalable solution for causal inference tasks.
Acknowledgment
D.I. acknowledges support from NSF (IIS-2212097), ARL (W911NF-2020-221), and ONR (N00014-23-C-1016). Y.X. acknowledges partial support from NSF DMS-2134037, CMMI-2112533, and the Coca-Cola Foundation. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsor(s).
Appendix A Proofs
Proposition 4.1. Assume that $p_{t}(y|x,a)>0$ for all $y\in\mathbb{R}^{d_{Y}}$ and $t\in[0,1]$. Then, up to a constant independent of $\theta$, the Conditional Flow Matching (CFM) loss and the original Flow Matching (FM) loss are equivalent. Hence,
$\nabla_{\theta}\mathcal{L}_{\text{FM}}(\theta)=\nabla_{\theta}\mathcal{L}_{\text{CFM}}(\theta).$
Proof:
We denote the conditional reference velocity field as $u(y,t,x,a|y_{0})=\frac{d\phi_{t}}{dt}=(1-\sigma_{\text{min}})y_{1}-y_{0}$, as defined in (6). We further define the marginal (unconditional) reference velocity field as:
[…]
where $p(y,t,x,a|y_{0})$ and $p(y,t,x,a)$ denote the probability densities induced by $u(y,t,x,a|y_{0})$ and $u(y,t,x,a)$, respectively. Next, the original flow matching loss, unconditioned on $y_{0}$, is defined as:
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y_{t},X,A}(\cdot)}\|v_{\theta}(y_{t},t,x,a)-u(y_{t},t,x,a)\|^{2}.$ (15)
In the main text, we chose $\phi_{t}$ as the linear interpolant between $y_{0}$ and $y_{1}$, which results in a constant reference velocity $\frac{d\phi_{t}}{dt}$ independent of $t$. This leads to the simplified CFM loss presented in Eq. (9). However, the general CFM loss, corresponding to an arbitrary reference velocity field $u(y,t\mid y_{0})$ conditioned on $y_{0}$, is defined as:
$\mathcal{L}_{\text{CFM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})}\|v_{\theta}(\cdot)-u_{t}(\cdot|y_{0})\|^{2}$ (16)
Next, we have that:
$\|v_{\theta}(y,t,x,a)-u(y,t,x,a)\|^{2}$ $=\|v_{\theta}(\cdot)\|^{2}-2\langle v_{\theta}(\cdot),u(\cdot)\rangle+\|u(\cdot)\|^{2}$ (17)
$\|v_{\theta}(y,t,x,a)-u(y,t,x,a|y_{0})\|^{2}$ $=\|v_{\theta}(\cdot)\|^{2}-2\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle+\|u(\cdot|y_{0})\|^{2}.$
Besides, the term $\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y_{t},X,A}(\cdot)}\|v_{\theta}(\cdot)\|^{2}$ can be equivalently expressed in conditional form as:
[…]
$=\int\|v_{\theta}(\cdot)\|^{2}p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\,dy\,dx\,da\,dt$
$=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})}\|v_{\theta}(\cdot)\|^{2}.$
Furthermore, by the law of total expectation, we have:
[…]
$=\int\left\langle v_{\theta}(\cdot),\int u(\cdot|y_{0})p_{Y_{t},X,A}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\right\rangle\,dy\,dx\,da\,dt$
$=\int\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle p_{Y_{t},X,A}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\,dy\,dx\,da\,dt$
$=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})}\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle.$
Finally, because $u_{t}(\cdot\mid y_{0})$ does not depend on $\theta$, substituting Eqs. (18) and (19) into Eq. (15) and differentiating with respect to $\theta$ yields:
$\nabla_{\theta}\mathcal{L}_{\text{FM}}(\theta)=\nabla_{\theta}\mathcal{L}_{\text{CFM}}(\theta).$ (20)
Proposition 4.2. Given a noise sample $y_{1}\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=\log q(y_{1})+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)dt.$
Proof:
Samples from Continuous Normalizing Flows (CNFs) evolve according to the following Neural ODE:
$\frac{dy(t)}{dt}=v_{\theta}(y(t),t,x,a),\quad t\in[0,1],$ (21)
which induces a corresponding evolution of the sample density governed by the Liouville continuity equation:
$\partial_{t}p(y,t)+\nabla\cdot\big{(}p(y,t)v_{\theta}(y,t,x,a)\big{)}=0.$ (22)
Here, $p(y,t)$ denotes the time-dependent probability density of the samples $y$.
Next, we have that the dynamics of the density $p(\cdot)$ governed by the velocity field $v_{\theta}(y,t,x,a)$ is given by:
$\frac{d}{dt}\log p(y(t),t)$ $=\frac{\nabla p(y(t),t)\cdot\partial_{t}y(t)+\partial_{t}p(y(t),t)}{p(y(t),t)}$ (23)
$=\frac{\nabla p\cdot\partial_{t}y+\partial_{t}p}{p}\Big{|}_{(y(t),t)}$ (24)
[…]
$=\frac{\nabla p\cdot v_{\theta}-(\nabla p\cdot v_{\theta}+p\nabla\cdot v_{\theta})}{p}\Big{|}_{(y(t),t)}$ (26)
$=-\nabla\cdot v_{\theta}.$ (27)
Starting from an initial sample $y_{1}\sim q(\cdot)$ and integrating from $t=1$ to $t=0$, we have:
$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=\log q(y_{1})+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)dt.$ (28)
Besides, if $q(\cdot)$ is selected as the commonly used $N(0,I)$, we have:
$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=-\frac{d_{Y}}{2}\log 2\pi-\frac{1}{2}\|y_{1}\|^{2}+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)dt.$ (29)
Appendix B Experimental Details
B.1 Synthetic Causal Datasets
Counterfactual outcomes are unobservable in real-world data but can be simulated using synthetic datasets. In our experiments, we evaluate on widely used benchmarks, including the Atlantic Causal Inference Conference datasets (ACIC 2016 and ACIC 2018), the Infant Health and Development Program (IHDP) [24, 42], and the IBM Causal Inference Benchmark [56].
The ACIC 2016 dataset contains 4,802 samples with 82 covariates each ($d_{X}=82$), ACIC 2018 has 1,000 samples with 177 covariates, IHDP includes 747 samples with 25 covariates, and the IBM dataset consists of 1,000 samples with 177 covariates. Each dataset provides three synthetic values per individual: the factual outcome $y^{(a)}$, the assigned treatment $a$, and the mean potential outcomes under both treatments, $\mu_{0}$ (for $a=0$) and $\mu_{1}$ (for $a=1$).
Moreover, under the following dataset synthesis assumptions: (1) fully specified outcome model $Y^{(A)}=f(X,A)+\epsilon$; (2) shared noise across treatment arms for each individual; and (3) invertibility of the structural function such that, given $X$, $A$, and $Y^{(A)}$, one can solve for $\epsilon$ — we can synthesize the counterfactual outcome through the following procedure:
1. Abduction: infer the unobserved noise term $\epsilon^{\prime}$ from the factual outcome $Y^{(A)}$, covariates $X$, and treatment $A$;
2. Action: intervene by changing the treatment assignment to $1-A$;
3. Prediction: use the inferred noise $\epsilon^{\prime}$ and new treatment to compute the counterfactual outcome via $Y^{(1-A)}=f(X,1-A)+\epsilon^{\prime}$.
All datasets are generated using diverse simulation mechanisms, and we provide the specific versions used along with our code. For example, the IHDP dataset is generated using 25 covariates and a coefficient matrix $W$, following the data-generating process:
$X\sim\text{Real-world distribution }p_{X}(\cdot),$ (30)
$A\sim\text{Real-world Treatment assignment based on}\ X,$
$Y=A(X\beta-\omega)+(1-A)\exp((X+W)\beta)+\epsilon,\quad\epsilon\sim N(0,1).$
where $\beta$, $W$, and $\omega$ are parameters defined by the specific simulation setting.
B.2 Counterfactual Image Experiments
To intuitively visualize PO-Flow’s counterfactual performance and evaluate its effectiveness on high-dimensional data, we train PO-Flow on the latent space of a pre-trained Variational Autoencoder (VAE) [36] using the CelebA dataset.
The CelebFaces Attributes Dataset (CelebA) comprises 202,599 face images, each with a resolution of $178\times 218$ pixels. Each image is annotated with 40 binary facial attributes. In our experiments, we selected the male/female attribute as the treatment variable $A\in\{\text{male},\text{female}\}$, and defined covariates $X\in\mathbb{R}^{6}$ to represent features such as hair, eyebrows, bangs, and lips.
The VAE is pre-trained on the full CelebA dataset with a 512-dimensional latent space. Each $178\times 218$ face image is then encoded into this latent space, yielding the training target $Y\in\mathbb{R}^{512}$. PO-Flow is subsequently trained on the triplet $(Y,X,A)$ to generate counterfactual male face representations given original female faces.
B.3 Neural Network Structure and Implementation Information
The neural network used in experiments on causal datasets (except the CelebA image task) takes as input $y(t)$, $x$, $t$, and $a$. The model is composed of the following components: (1) an outcome embedding layer that projects $y(t)$ to the same dimension as the concatenated $(x,a)$; (2) a FiLM layer that applies feature-wise affine transformations to the embedded $y(t)$, conditioned on $(x,a)$; (3) two residual blocks, each consisting of a gated two-layer MLP conditioned on $(x,a)$, with residual averaging and skip connections; and (4) a projection layer that maps the final hidden representation to two velocity vectors $v_{0}$ and $v_{1}$, corresponding to treatments 0 and 1.
The neural network used for the counterfactual image generation task has an architecture composed of the following components: (1) a conditioning MLP that encodes the pair $(t,a)$ into a hidden feature vector via two linear layers and a ReLU activation; (2) an input layer that takes the concatenation of $y(t)$ and $(t,a)$, followed by a ReLU-activated linear transformation; (3) four residual blocks, each composed of two linear layers with a ReLU nonlinearity, where the conditioning feature is injected additively at each block; and (4) an output linear layer that maps the final hidden representation back to the latent outcome dimension.
All experiments are conducted on a V100 GPU with 16 GB of memory. The training time for causal datasets (excluding CelebA counterfactual face image generation) depends on the dataset and batch sizes. For example, training on the ACIC 2018 dataset with a batch size of 256 typically takes 2–4 seconds per epoch. Log-density evaluation, as discussed in Section 4.3, involves divergence integration and requires significantly more time than other test metrics.
B.4 Runge-Kutta numerical integration method
We apply the classical 4th-order Runge-Kutta method to solve the Neural ODE in the forward process of PO-Flow. Given a factual outcome $y^{(a)}$, we aim to compute the latent embedding $z_{y}=y(1)$ by integrating the velocity field $v_{\theta}$ from $t=0$ to $t=1$, with initial condition $y(0)=y^{(a)}$. Let $h$ denote the step size, and define discrete time steps $t_{k}=kh$, for $k=0,1,\dots,T$, such that $Th=1$. The update rule at each step is:
$\small y(t_{k+1})=y(t_{k})+\frac{h}{6}\left(k_{1}+2k_{2}+2k_{3}+k_{4}\right),$
where
$k_{1}$ $=v_{\theta}(y(t_{k}),t_{k},x,a),\ k_{2}=v_{\theta}\left(y(t_{k})+\frac{h}{2}k_{1},t_{k}+\frac{h}{2},x,a\right),$
$k_{3}$ $=v_{\theta}\left(y(t_{k})+\frac{h}{2}k_{2},t_{k}+\frac{h}{2},x,a\right),\ k_{4}=v_{\theta}\left(y(t_{k})+hk_{3},t_{k}+h,x,a\right),$
with $y(0)=y^{(a)}$ and $y(1)=z_{y}$. This produces a numerically integrated trajectory from the observed sample to its latent representation.
B.5 Hutchinson trace estimator
The computations of the log-density of the POs involve the calculation of $\nabla\cdot v_{k}(x,t)$, i.e., the divergence of the velocity field represented by a neural network. This may be computed by brute force using reverse-mode automatic differentiation, which is much slower and less stable in high dimensions.
We can express $\nabla\cdot v_{\theta}(y,t,x,a)=\mathbb{E}_{\epsilon\sim p(\epsilon)}\left[\epsilon^{T}J_{v}(y)\epsilon\right]$, where $J_{v}(y)$ is the Jacobian of $v_{\theta}(y,t,x,a)$ at $y$. Given a fixed $\epsilon$, we have $J_{v}(y)\epsilon=\lim_{\sigma\to 0}\frac{v_{\theta}(y+\sigma\epsilon,t,x,a)-v_{\theta}(y,t,x,a)}{\sigma}$, which is the directional derivative of $v_{\theta}$ along the direction $\epsilon$. Therefore, for a sufficiently small $\sigma>0$, we can propose the following Hutchinson estimator  [31, 64]:
$\small\nabla\cdot v_{\theta}(y(t),t,x,a)\approx\mathbb{E}_{p(\epsilon)}\left[\epsilon\cdot\frac{v_{\theta}(y(t)+\sigma\epsilon,t,x,a)-v_{\theta}(y(t),t,x,a)}{\sigma}\right],$ (31)
where $p(\epsilon)$ is a distribution in $\mathbb{R}^{d_{Y}}$ satisfying $\mathbb{E}[\epsilon]=0$ and Cov$(\epsilon)=I$ (e.g., a standard Gaussian). This approximation becomes exact as $\sigma\to 0$.
B.6 Adapted Implementation Details for DiffPO and GANITE Counterfactuals
DiffPO was not originally designed for counterfactual tasks. However, since DiffPO uses a diffusion model as its base, we are able to adapt the algorithm to incorporate an "encoding" and "decoding" structure similar to that of PO-Flow. Our adapted implementation of their algorithm for counterfactual estimation is summarized in Algorithm 3.
  Input: $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$
  1: $z_{0}\leftarrow y^{(a)}$
  2: Forward Process:
  3:     for $t=0,\cdots,T-1$:
  4:       $z_{t+1}\leftarrow\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}z_{t}+f_{\theta}(z_{t},t|x,a)\left(\sqrt{1-\alpha_{t+1}}-\sqrt{\frac{\alpha_{t+1}(1-\alpha_{t})}{\alpha_{t}}}\right)$
  5:     $z_{y}\leftarrow z_{T}$
  6: Reverse Process:
  7:     Follow their official implementation in Section 5.3
  8: Return $\hat{y}^{(1-a)}:=\hat{y}_{0}$
Algorithm 3 Adapted DiffPO for Counterfactual
Similarly, GANITE was not originally designed for counterfactual inference. It consists of a generator and a discriminator, with the final output targeting CATE estimation. However, the generator also produces a proxy counterfactual given the factual input, which we use to evaluate its counterfactual estimation performance.
Appendix C Additional Experimental Results
C.1 RMSE Convergence
PO-Flow demonstrates rapid convergence across datasets, requiring only around 250 training iterations (not epochs) with a batch size of 200 per iteration. This highlights the model’s strong training efficiency.
C.2 Additional counterfactual images
Figure 6: PO-Flow counterfactual image generation from female to male faces. $(a)$: Each row shows a factual outcome $y^{(a)}$ (1st column), its encoded latent $z_{y}$ (2nd column), and reverse-time decoding at $t=2/3$, $t=1/3$, and $t=0$ (3rd–5th columns), yielding the counterfactual $\hat{y}^{(1-a)}$. $(b)$: Counterfactuals decoded from perturbed latents $z_{y}+\delta$, with $\delta\sim\mathcal{N}(0,0.2I)$.

### v2

@@
 Abstract
-We propose PO-Flow, a novel continuous normalizing flow (CNF) framework for causal inference that jointly models potential outcomes and counterfactuals.
-Trained via flow matching, PO-Flow provides a unified framework for individualized potential outcome prediction, counterfactual predictions, and uncertainty-aware density learning.
-Among generative models, it is the first to enable density learning of potential outcomes without requiring explicit distributional assumptions (e.g., Gaussian mixtures), while also supporting counterfactual prediction conditioned on factual outcomes in general observational datasets.
-On benchmarks such as ACIC, IHDP, and IBM, it consistently outperforms prior methods across a range of causal inference tasks.
-Beyond that, PO-Flow succeeds in high-dimensional settings, including counterfactual image generation, demonstrating its broad applicability.
+Predicting potential and counterfactual outcomes from observational data is central to clinical decision-making, where physicians must weigh treatments for an individual patient rather than relying solely on average effects at the population level.
+We propose PO-Flow, a continuous normalizing flow (CNF) framework for causal inference that jointly models potential outcomes and counterfactuals.
+Trained via flow matching, PO-Flow provides a unified approach to average treatment effect estimation, individualized potential outcome prediction, and counterfactual prediction.
+Besides, PO-Flow directly learns the densities of potential outcomes, enabling likelihood-based evaluation of predictions.
+Furthermore, PO-Flow explores counterfactual outcome generation conditioned on the observed factual in general observational datasets, with a supporting recovery result under certain assumptions.
+PO-Flow outperforms modern baselines across diverse datasets and causal tasks in the potential outcomes framework.
 1 Introduction
@@
 However, in the real world, only one outcome is observed, while the other potential outcome remains fundamentally unknown [25].
-Accurately estimating both POs is crucial for decision-making in domains such as healthcare [41, 27, 50], where clinicians may wish to compare a patient’s likely outcomes under different treatment options [60, 20].
-In the causal inference literature, the vast majority of methods focus on estimating the average treatment effect (ATE) and the conditional average treatment effect (CATE) at the population level [32, 25, 1, 22, 37, 42, 55, 67].
-It is important to note that although some methods perform well in estimating population-level CATE, they may underperform in individualized potential outcome (PO) prediction, as accurately predicting POs in finite samples is challenging [52].
+Accurately estimating both POs is crucial for decision-making in healthcare [37, 24, 46], where clinicians may wish to compare a patient’s likely outcomes under different treatment options rather than rely solely on average effects [56, 17].
+In the Potential Outcomes (POs) literature, the vast majority of methods focus on estimating the conditional average treatment effect (CATE): $\mathbb{E}[Y^{(1)}-Y^{(0)}|X]$, at the population level [28, 22, 1, 19, 33, 38, 50, 61].
+It is important to note that although some methods perform well in estimating population-level CATE, they may underperform in individualized POs prediction, as accurately predicting POs in finite samples is challenging [47].
 However, making decisions for individuals based on average treatment effects can be misleading.
-Therefore, accurately predicting individualized POs and learning their full distributions with uncertainty is a crucial task [69, 35].
+Therefore, accurately predicting individualized POs and learning their full distributions $p(Y^{(A)}|X,A)$ is a crucial task [63, 31].
 Recent works [61, 40, 41] have explored generative models to sample POs and estimate empirical distributions.
 However, methods that learn the full density of POs without relying on explicit distributional assumptions (e.g., Gaussian mixtures) remain scarce.
-In addition, another critical challenge is estimating the counterfactual outcome conditioned on the factual—that is, predicting what would have happened under the alternative treatment, given the observed outcome.
-This task is especially relevant in post-hoc clinical analyses, where practitioners ask, “Would the patient have fared better on another therapy?” [66, 29, 21].
-Counterfactual prediction differs from standard PO estimation—it requires reasoning about outcomes under an alternative treatment conditioned on the observed outcome, introducing dependencies between observed and unobserved outcomes.
+In addition, another critical challenge is estimating the counterfactual outcome conditioned on the factual: $Y_{\mathrm{CF}}^{(1-A)}:=Y^{(1-A)}|X,Y^{(A)}$, i.e., predicting what would have happened under the alternative treatment, given the observed outcome $Y^{(A)}$.
+This task is especially relevant in post-hoc clinical analyses, where practitioners ask, “Would the patient have fared better on another therapy?” [60, 18].
+Counterfactual prediction differs from standard PO estimation—it requires reasoning about outcomes under an alternative treatment conditioned on the observed factual outcome, introducing dependencies between observed and unobserved outcomes.
 Structural Causal Models (SCMs) [44] and modern extensions [42, 6, 12, 7, 59, 45, 8] offer the most widely used theoretical framework for counterfactual reasoning.
 SCMs methods typically assume known causal structures or aim to learn them explicitly.
-Yet, it may be possible to estimate counterfactuals directly without recovering the SCM [70].
+However, it may be possible to estimate counterfactuals directly without recovering the SCM [64].
 In PO-Flow, we do not explicitly assume an SCM, but seek to directly estimate counterfactual outcomes.
-We aim to develop a generative modeling framework that jointly supports population-level CATE estimation, individualized PO prediction with density modeling, and counterfactual outcome prediction without explicitly modeling an SCM—an area that has been relatively less explored.
+Despite recent progress, two key gaps remain in the potential outcomes literature.
+First, counterfactual outcome estimation conditioned on the observed factual outcome has received little attention, though it is crucial for post-hoc analyses.
+Second, beyond counterfactuals, most methods are specialized—some target CATE estimation, others aim at potential outcome prediction or distributional modeling—but few provide a unified and scalable algorithmic framework that can jointly address these tasks.
+1.1 Main Contributions
+We propose PO-Flow, a continuous normalizing flow (CNF) framework for potential outcomes and counterfactuals.
+Our main contributions are: (1) Enable counterfactual outcome generation conditioned on the observed factual, together with a supporting counterfactual recovery result (Corollary 4.5) under certain assumptions; (2) Provide a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction; (3) Learn the density of potential outcomes, supporting principled uncertainty quantification; (4) Achieve superior performance over state-of-the-art methods across diverse causal tasks and benchmark datasets.
 1.2 Related Works
@@
 Broadly, these fall into model-agnostic and model-specific approaches.
-Model-agnostic estimators, or meta-learners [37, 15, 16, 2, 54], typically follow one of two strategies [43]: (a) one-step methods, which fit regression models and compute CATE as the difference in predicted outcomes, and (b) two-step methods, which estimate nuisance functions to construct pseudo-outcomes, then regress them on covariates.
+Model-agnostic estimators, or meta-learners [33, 15, 14, 2, 49], typically follow two strategies [39]: (a) one-step methods, which fit regression models and compute CATE as the difference in predicted outcomes, and (b) two-step methods, which estimate nuisance functions to construct pseudo-outcomes, then regress them on covariates.
 In contrast, model-specific estimators adapt machine learning techniques for CATE estimation, including methods such as [13, 50, 38, 30, 61].
@@
 This limitation partly arises because many methods assume that CATE has a simpler functional structure than POs.
-Nevertheless, several CATE approaches are also designed with POs prediction as the target [37, 55, 67, 42, 45, 43], and these will serve as our primary baselines for experimental comparisons.
+Nevertheless, several CATE approaches are also designed with POs prediction as the target [33, 50, 61, 38, 41, 39].
 Counterfactual Outcomes Prediction: Structural Causal Models (SCMs) provide a well-established framework for counterfactual reasoning [44, 42, 6, 12, 7, 59, 45, 8], but practical applications often require specifying the full causal graph and structural equations, which may be difficult or infeasible for real-world data.
@@
 While SCMs offer theory-grounded counterfactual reasoning, general-purpose methods for individualized counterfactual prediction—without explicitly modeling SCMs or relying on auxiliary information—remain scarce.
-Although [67, 43] were not originally designed for counterfactuals, we adapt them by incorporating factual information into intermediate steps and use them as baselines.
-Diffusion and Flow Models: Diffusion models [26, 57, 58] and flow-based generative models [12, 64, 40, 4] have demonstrated strong performance across a range of applications, including image generation, statistical sampling [18, 59, 61], and physics-informed modeling [8, 68, 51].
+Although [61, 39] were not originally designed for counterfactuals, we can adapt them by incorporating factual information into intermediate steps and use them as baselines.
+Diffusion and Flow Models: Diffusion models [23, 52, 54] and flow-based generative models [10, 58, 36, 4] have demonstrated strong performance across a range of applications.
 Recently, diffusion and flow-based generative models have been applied to causal inference tasks such as counterfactual image generation [48, 16], causal query answering [9], and inference over causal graphs or structural causal models [8, 29].
-However, to our knowledge, only two generative algorithms have been proposed specifically for the potential outcomes (PO) task: the diffusion-based DiffPO [43] and the discrete normalizing flow-based INFs [45].
+To our knowledge, two generative algorithms have been proposed specifically for the potential outcomes (PO) task: the diffusion-based DiffPO [39] and the discrete normalizing flow-based INFs [41].
 We conduct extensive comparisons with both methods in this work.
-1.2 Main Contributions
-We propose PO-Flow, a continuous normalizing flow (CNF) framework for potential outcomes and counterfactuals.
-Our main contributions are: (1) PO-Flow provides a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction—without explicitly modeling structural causal models (SCMs) or relying on auxiliary information for the counterfactual task; (2) PO-Flow enables learning the density of predicted potential outcomes, offering deeper insights into uncertainty; (3) PO-Flow outperforms state-of-the-art methods across a range of causal tasks on multiple benchmarks.
 2 Preliminaries
@@
 Formally, in $\mathbb{R}^{d}$, given an input $y_{0}=y(0)$ at time $t=0$, the transformation to the output $y_{1}=y(1)$ at time $t=1$ is governed by:
-$\frac{dy(t)}{dt}=v_{\theta}(y(t),t),\quad t\in[0,1],$ (1)
-where $v_{\theta}:\mathbb{R}^{d_{Y}+1}\to\mathbb{R}^{d_{Y}}$ represents the velocity field, which is parameterized by a neural network with parameter $\theta$.
+$\small\frac{dy(t)}{dt}=v_{\theta}(y(t),t),\quad t\in[0,1],$ (1)
+where $v_{\theta}:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ represents the velocity field, which is parameterized by a neural network with parameter $\theta$.
 The time horizon is rescaled to $[0,1]$ without loss of generality.
@@
 This continuity equation is given by:
-$\partial_{t}p(y,t)+\nabla\cdot(p(y,t)v_{\theta}(y,t))=0,\quad t\in[0,1],$ (2)
-[…]
+$\small\partial_{t}p(y,t)+\nabla\cdot(p(y,t)v_{\theta}(y,t))=0,\quad t\in[0,1],$ (2)
+where $\nabla\cdot$ denotes the divergence operator.
 When the Neural ODE is well-posed, it defines a continuous and invertible mapping from the initial sample $y_{0}$ to the terminal sample $y_{1}$.
@@
 The original FM objective is defined as:
-$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],y\sim p(\cdot,t)}\left[\|v_{\theta}(y,t)-u(y,t)\|^{2}\right],$ (3)
+$\small\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],y\sim p(\cdot,t)}\left[\|v_{\theta}(y,t)-u(y,t)\|^{2}\right],$ (3)
 where $u(y,t)$ is an analytically specified target velocity field.
@@
 The interpolated path is defined as:
-$\phi_{t}:=I_{t}(y_{0},y_{1}),\quad t\in[0,1].$ (4)
+$\small\phi(y_{0},y_{1};t):=I_{t}(y_{0},y_{1}),\quad t\in[0,1].$ (4)
 Here, $I_{t}$ can be analytically specified. Conditioned on $y_{0}$, the optimal and most information-preserving choice is the linear interpolation path [36]:
-$\phi_{t}=(1-t)y_{0}+(t+\sigma_{\text{min}}(1-t))y_{1},$ (5)
-where $y_{1}\sim N(0,I)$ and $\sigma_{\text{min}}$ is a small positive hyperparameter that ensures $p(\phi_{t},0)\sim N(y_{0},\sigma_{\text{min}}^{2}I)$.
-When $\sigma_{\text{min}}=0$, this reduces to the strict linear interpolant $\phi_{t}=(1-t)y_{0}+ty_{1}$.
-Here, $\phi_{t}$ serves as a pre-specified reference trajectory that the ODE solution $y(t)$ should follow at time $t$.
+$\small\phi(y_{0},y_{1};t)=(1-t)y_{0}+(t+\sigma_{\text{min}}(1-t))y_{1},$ (5)
+where $y_{1}\sim N(0,I)$ and $\sigma_{\text{min}}$ is a small positive hyperparameter that ensures $p(\phi,0)\sim N(y_{0},\sigma_{\text{min}}^{2}I)$.
+When $\sigma_{\text{min}}=0$, this reduces to the strict linear interpolant $\phi=(1-t)y_{0}+ty_{1}$.
+Here, $\phi$ serves as a pre-specified reference trajectory that the ODE solution $y(t)$ should follow at time $t$.
 Under the linear interpolant defined in (5), the corresponding reference velocity field becomes:
-$\frac{d\phi_{t}}{dt}=(1-\sigma_{\text{min}})y_{1}-y_{0}.$ (6)
+$\small\frac{d\phi}{dt}=(1-\sigma_{\text{min}})y_{1}-y_{0}.$ (6)
 Training loss: CFM trains the flow $v_{\theta}(y(t),t)$ by directly regressing it onto the reference velocity field (6). The training objective is given by:
-$\mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],y_{0}\sim p_{Y}(\cdot),y_{1}\sim q(\cdot)}\|v_{\theta}(\phi_{t},t)-\frac{d\phi_{t}}{dt}\|^{2}.$ (7)
+$\small\mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],y_{0}\sim p_{Y}(\cdot),y_{1}\sim q(\cdot)}\|v_{\theta}(\phi,t)-\frac{d\phi}{dt}\|^{2}.$ (7)
 3 Problem Settings
-Notations and Setups: We consider a setting where each individual is described by covariates $X\in\mathbb{R}^{d_{X}}$, a treatment assignment $A\in\{0,1\}$, and an outcome $Y^{(A)}\in\mathbb{R}^{d_{Y}}$ under treatment $A$.
+Notations and Setups: We consider a setting where each individual is described by covariates $X\in\mathbb{R}^{d_{X}}$, a treatment assignment $A\in\{0,1\}$, and a potential outcome $Y^{(A)}\in\mathbb{R}$ under treatment $A$.
 The observational dataset is $\mathcal{D}=\{(y^{(a)}_{i},x_{i},a_{i})\}_{i=1}^{n}\sim p_{Y,X,A}(\cdot)$, where $y_{i}^{(a)}$ is the observed potential outcome under treatment $a_{i}$, and $(x_{i},a_{i})\sim p_{X,A}(\cdot)$ are the associated covariates and treatment.
@@
 For convenience, we denote $Y^{(a)}$ and $\hat{Y}^{(a)}$ as the true and predicted potential outcomes.
-Counterfactual Outcomes: In contrast to potential outcomes, we define the counterfactual outcome as $Y_{\text{cf}}^{(1-a)}:=Y^{(1-a)}\mid X,A=a,Y^{(a)}$: the outcome had the individual received the opposite treatment, conditioned on its observed potential (factual) outcome.
+Counterfactual Outcomes: In contrast to potential outcomes, we define the counterfactual outcome as $Y_{\text{cf}}^{(1-a)}:=Y^{(1-a)}\mid X,A=a,Y^{(a)}$, which is the outcome had the individual received the opposite treatment, conditioned on its observed potential (factual) outcome.
 For convenience, we denote $Y_{\text{cf}}^{(1-a)}$ and $\hat{Y}_{\text{cf}}^{(1-a)}$ as the true and predicted counterfactual outcomes, respectively.
-To ensure the identifiability of potential outcomes, we adopt the standard assumptions commonly used in the literature [43, 45, 67]:
+To ensure the identifiability of potential outcomes (in Section 4.1), we adopt the standard assumptions commonly used in the literature [39, 41, 61]:
 Assumption 3.1.
@@
 (2) Unconfoundedness: All confounders are observed, i.e., $\{Y^{(0)},Y^{(1)}\}\perp\!\!\!\perp A\mid X$.
-(3) Overlap: Treatment assignment is probabilistic: $10<P(A=1\mid X=x)<$ 1, for all $x\in\mathcal{X}$.
-Counterfactuals are rarely identifiable without strong assumptions [63, 47].
-Estimating them can often be framed as an optimal transport problem [11, 17].
-To constrain the space of joint transport mappings, our PO-Flow framework introduces the following assumption, applicable only to Section 4.2:
-Assumption 3.2.
-There exists a deterministic and invertible mapping $f:\mathbb{R}^{d_{Y}+d_{X}}\times\{0,1\}\rightarrow\mathbb{R}^{d_{Z}}$ that encodes the observed outcome as $Z_{Y}=f(Y^{(A)},X,A)$.
-A counterfactual latent representation can be obtained as $Z^{\prime}_{Y}=Z_{Y}+\delta$, where $\delta\sim\mathcal{N}(0,\Sigma(X))$ is an optional covariate-dependent noise term.
-The counterfactual is then recovered via $Y_{\text{cf}}^{(1-A)}=f^{-1}(Z^{\prime}_{Y},X,1-A)$.
+(3) Overlap: Treatment assignment is probabilistic: $0<P(A=1\mid X=x)<1$, for all $x\in\mathcal{X}$.
+In Section 4.2, we will introduce additional assumptions on counterfactual outcomes estimation in PO-Flow and provide the corresponding theoretical analyses for counterfactual recovery.
 4 PO-Flow: A Continuous Normalizing Flow for Potential Outcomes
@@
 predicting individualized potential outcomes, 2.
-estimating counterfactual outcomes, and 3.
+predicting counterfactual outcomes, and 3.
 learning the densities.
@@
 Our learning target is the trajectory of the sample $y(t)$ over $t\in[0,1]$.
-To condition the dynamics on both covariates $x$ and treatment $a$, our Neural ODE is defined as:
-$\small\frac{dy(t)}{dt}=v_{\theta}(y(t),t,x,a),\quad t\in[0,1],$ (8)
-where $v_{\theta}:\mathbb{R}^{d_{Y}+d_{X}+2}\rightarrow\mathbb{R}^{d_{Y}}$ is parametrized by a neural network.
+To condition the dynamics on both covariates $x$ and treatment $a$, the Neural ODE is defined as:
+$\small\frac{dy(t)}{dt}=v_{\theta}(y(t),t;x,a),\quad t\in[0,1],$ (8)
+where $v_{\theta}$ is parametrized by a neural network, with $y(0)$ following the data distribution and $y(1)\sim q(\cdot)$.
 Notation: Given a data point $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$, we define $y_{0}=y^{(a)}$ as the observed training sample at time $t=0$ in the Neural ODE, and $y_{1}\sim q(\cdot)$ as the noise sample at time $t=1$.
@@
 Training loss: PO-Flow adopts the Conditional Flow Matching (CFM) objective for training. Under the updated definitions, our training loss retains the form of Equation (7) and is given by:
-$\min_{\theta}\ \mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot),q(\cdot)}\|v_{\theta}(\phi_{t},t,x,a)-\frac{d\phi_{t}}{dt}\|^{2},$ (9)
-where $\phi_{t}$ and $\frac{d\phi_{t}}{dt}$ are defined in (5) and (6), respectively.
-Remark: Following prior work [23, 43], one may estimate propensity scores $w(x)=p(A=1\mid X=x)$ using a pre-trained classifier and apply inverse probability weighting $\frac{1}{w(x)}$ in the loss.
-However, we found this did not improve performance, but both implementations are included in our code.
-Forward Process: Given the velocity field $v_{\theta}$, the forward process refers to pushing a factual outcome $y^{(a)}$ (at $t=0$) to a latent embedding $z_{y}$ (at $t=1$) via:
-[…]
-Through the forward process, $z_{y}$ encodes information from the factual outcome $y^{(a)}$ as well as its associated covariates $x$ and treatment $a$.
-Reverse Process: The reverse process is initialized differently for interventional and counterfactual outcome prediction, as detailed in Sections 4.1 and 4.2.
-In general, given a noise sample $y_{1}\sim q(\cdot)$ at $t=1$, the reverse process is defined as:
-[…]
-Empirically, we use the Runge-Kutta method for numerical integration, with implementation details provided in Appendix B.4.
-During training, we construct synthetic paths $\phi_{t}$ between $y_{0}$ and $y_{1}\sim q(\cdot)$ using the CFM loss, conditioning on $y_{0}$.
-However, Proposition 4.1 shows that the gradient of the CFM loss with respect to the model parameters is equivalent to the gradient of the original FM loss.
-Proposition 4.1.
-Assume that $p_{t}(y|x,a)>0$ for all $y\in\mathbb{R}^{d_{Y}}$ and $t\in[0,1]$.
-Then, up to a constant independent of $\theta$, the Conditional Flow Matching (CFM) loss and the original Flow Matching (FM) loss are equivalent.
-Hence,
-$\nabla_{\theta}\mathcal{L}_{\text{FM}}(\theta)=\nabla_{\theta}\mathcal{L}_{\text{CFM}}(\theta).$
-As a result of Proposition 4.1, the learned velocity field $v_{\theta}$ during the reverse process generates samples $\hat{y}_{0}$ independent of the individual samples $y_{0}$ used during training.
-Consequently, since $y_{1}\sim q(\cdot)$ is random, $v_{\theta}$ induces a distribution over outcomes conditioned on $(x,a)$.
+$\min_{\theta}\ \mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot),q(\cdot)}\|v_{\theta}(\phi,t;x,a)-\frac{d\phi}{dt}\|^{2},$ (9)
+where $\phi$ and $\frac{d\phi}{dt}$ are defined in (5) and (6), respectively, with $y_{0}=y^{(a)}$ and $y_{1}\sim q(\cdot)$.
+Remark: Following prior works [20, 39], one may estimate propensity scores $w(x)=p(A=1\mid X=x)$ with a pre-trained classifier and apply inverse weighting $\frac{1}{w(x)}$ in the loss.
+We found this provided little performance improvement, but both implementations are included in our code.
+Forward Process: We treat the forward process as an encoding operation, denoted by the function $Z:=\Phi_{\theta}(Y^{(A)};\,X,A)$.
+Given the velocity field $v_{\theta}$, the forward process refers to pushing a factual outcome $y^{(a)}$ (at $t=0$) to a latent embedding $z$ (at $t=1$) via:
+$z=\Phi_{\theta}(y^{(a)};\,x,a)=y^{(a)}+\int_{0}^{1}v_{\theta}(y(t),t;x,a)dt,\ \text{with}\ y(0)=y^{(a)}.$ (10)
+Through the forward process, $z$ encodes information from the factual outcome $y^{(a)}$ as well as its associated covariates $x$ and treatment $a$.
+Reverse Process: We treat the reverse process as a decoding operation.
+It is initialized differently for potential outcome (Section 4.1) and counterfactual prediction (Section 4.2).
+In general, at $t=1$, given a latent representation $z$, which can be either sampled from $q(\cdot)=N(0,1)$ or obtained by encoding a factual sample, the reverse process is defined as:
+$\hat{y}_{0}=\Phi_{\theta}^{-1}(z;x,a)=z-\int_{0}^{1}v_{\theta}(y(t),t;x,a)dt,\ \text{with}\ y(1)=z.$ (11)
+Empirically, we use the Runge-Kutta method (Appendix B.3) for numerical integration.
+During training, we construct synthetic paths $\phi$ between $y_{0}$ and $y_{1}\sim q(\cdot)$ using the CFM loss conditioned on $y_{0}$.
+As shown in Proposition A.1 (Appendix), the gradient of the CFM loss is equivalent to that of the original FM loss.
+Consequently, the learned velocity field $v_{\theta}$ generates $\hat{y}_{0}$ independent of the specific training samples $y_{0}$, and since $y_{1}\sim q(\cdot)$ is random, $v_{\theta}$ induces a distribution over outcomes conditioned on $(x,a)$.
 4.1 PO-Flow for Potential Outcomes Prediction
@@
 Input: $(x,a)\sim p_{X,A}(\cdot)$
-  1: $y_{1}\sim q(\cdot)$
-[…]
-  Return $\hat{y}^{(a)}:=\hat{y}_{0}$
+ 1: $z\sim q(\cdot)$
+ 2: $\hat{y}_{0}=z-\int_{0}^{1}v_{\theta}(y(t),t;x,a)\,dt$
+Return: $\hat{y}^{(a)}:=\hat{y}_{0}$
 Potential Outcomes Prediction
@@
 Potential Outcomes Prediction
-The potential outcomes prediction in PO-Flow follows a simple reverse process, as defined in Equation (11).
-Starting from a noise sample $y_{1}\sim q(\cdot)$, we integrate the velocity field $v_{\theta}$ backward in time, conditioned on covariates $x$ and treatment $a$, to obtain a point estimate of the potential outcome $\hat{y}^{(a)}$.
-The procedure is summarized in Algorithm 1.
-In addition to predicting point estimates of potential outcomes, PO-Flow is capable of learning the probability densities of the potential outcomes, as detailed in Section 4.3.
+To sample potential outcomes $\hat{y}^{(a)}$ for an individual $(x,a)$, the procedure follows a simple reverse process, as defined in Equation (11).
+Starting from a noise sample $z\sim q(\cdot)$, we integrate the velocity field $v_{\theta}$ backward in time, conditioned on covariates $x$ and treatment $a$, to obtain a point estimate of the potential outcome $\hat{y}^{(a)}$.
+The steps are summarized in Algorithm 1.
+In addition to predicting point estimates of potential outcomes, PO-Flow is capable of learning the densities of the potential outcomes, as detailed in Section 4.3.
 4.2 PO-Flow for Counterfactual Outcomes Prediction
+The counterfactual outcome prediction consists of both a forward and a reverse process.
+Given a factual $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$, the forward process (Lines 1–2 in Alg.
+2) encodes $(y^{(a)},x,a)$ into a latent representation $z$.
+Subsequently, the reverse process (Lines 3–5) decodes $z$ under the counterfactual treatment 1a1-a, yielding the counterfactual prediction $\hat{y}_{\text{cf}}^{(1-a)}$.
+The process of counterfactual prediction is also illustrated in Figure 1.
 Require: Trained $v_{\theta}$
-  Input: $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$
-  1. Forward process (encoding):
-[…]
+Input: Factual sample $\small(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$
+ 1. Forward process {abduction}:
+3 2.  $z=y^{(a)}+\int_{0}^{1}v_{\theta}(y(t),t;x,a)\,dt,$
    with $y(0)=y^{(a)}$.
-  3. Uncertainty-aware perturbations:
-  4.     $\delta\sim N(0,\Sigma(x))$
-  5.     $z^{\prime}_{y}=z_{y}+\delta$
-  6. Reverse process (decoding):
-[…]
-         with $y(1)=z^{\prime}_{y}$.
+ 3. Reverse process {action and prediction}:
+ 4.  Set intervened action to 1a1-a;
+4 5.  $\hat{y}_{0}=z-\int_{0}^{1}v_{\theta}(y(t),t;x,1-a)\,dt,$
+   with $y(1)=z$.
 Return: $\hat{y}_{\text{cf}}^{(1-a)}:=\hat{y}_{0}$
@@
 Counterfactual Prediction
-Under Assumption 3.2, the counterfactual outcome prediction in PO-Flow consists of both a forward and a reverse process.
-Given a factual observation $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$, the forward process (Lines 2–3 in Algorithm 2) encodes the information $(y^{(a)},x,a)$ into a latent representation centered at $z_{y}$.
-Lines 4–6 introduce uncertainty-aware perturbations to $z_{y}$, reflecting covariate-dependent uncertainty.
-Finally, the reverse process (Lines 7–8) decodes the latent $z_{y}^{\prime}$ back into the counterfactual outcome prediction $\hat{y}_{\text{cf}}^{(1-a)}$.
 Figure 1: Illustration of counterfactual predictions. Left: two potential outcomes distribution. Right: base (noise) distribution.
-Because of the smooth and information-preserving nature of the PO-Flow path trained via flow matching, the encoded $z_{y}$ retains rich information from the factual outcome $y^{(a)}$, enabling accurate decoding of the counterfactual $\hat{y}_{\text{cf}}^{(1-a)}$.
-To our knowledge, this is the first work to explore counterfactual outcome prediction using flow or diffusion-based models.
-As illustrated in Figure 1, diffusion models rely on noisy and stochastic forward processes, which degrade the information in $z_{y}$ and lead to reversed samples that resemble interventional potential outcomes rather than counterfactuals.
-In Section 5.4, we present counterfactuals using an adapted implementation of DiffPO [43] (originally not designed for counterfactual), compared with PO-Flow.
+Because of the smooth and information-preserving nature of the continuous normalizing flow path, the encoded $z$ retains rich information from the factual outcome $y^{(a)}$, enabling accurate decoding of the counterfactual $\hat{y}_{\text{cf}}^{(1-a)}$ under counterfactual treatment 1a1-a.
+To our knowledge, this work is among the early attempts at counterfactual prediction in the Potential Outcomes (POs) literature.
+In the following, we present a supporting result on the counterfactual recovery properties of PO-Flow under monotone structural causal models (SCMs).
+While PO-Flow does not explicitly aim to model the true SCM during training or inference, we show that if the data are generated from monotone SCMs, then PO-Flow provides recovery guarantees under certain conditions.
+We denote the SCM for an individual as $Y^{(A)}=f_{X}(A,U)$, where $U$ denotes the exogenous noise.
+The encoding function is given by $Z:=\Phi_{\theta}(Y^{(A)};\,X,A)$, which corresponds to Equation (10).
+We begin by introducing the following assumptions:
+Assumption 4.1.
+(A1) $U\perp\!\!\!\perp A$.
+(A2) The structural causal model $f_{X}(A,U)$ is monotone in $U$ for each action $A\in\{0,1\}$.
+(A3) For every $x$, the encoded latent $Z=\Phi_{\theta}(Y^{(A)};\,X,A)$ induced by the flow is statistically independent of $A$, i.e., $Z\perp\!\!\!\perp A\ |\ X=x$.
+Remark 4.2.
+In the POs framework, covariates $x$ are multivariate while the potential outcome is univariate, $y\in\mathbb{R}$.
+Therefore, (A2) is invoked for the univariate case.
+It is automatically satisfied under an additive structural causal model, $Y^{(A)}=f_{X}(A,U):=f_{X}^{*}(A)+U$.
+Under certain identifiability conditions, it also holds for non-linear models [62, 55].
+The following theoretical results hold only under monotone SCM settings.
+(A3) assumes the encoded representation $Z=\Phi_{\theta}(Y^{(A)};\,X,A)$ in PO-Flow is statistically independent of treatment $A$, though not necessarily functionally invariant to $A$.
+In the infinite-data limit with exact training, the continuous normalizing flow (Equation (8)) maps each treatment-conditioned outcome $Y$ to the same base distribution $q(Z)=\mathcal{N}(0,1)$, so $p_{\theta}(Z\mid X,A)=q(Z)$ and (A3) holds exactly.
+In finite samples, deviations may arise; we assess this empirically in Appendix D.1.
+Under the Assumption 4.1, we present a result for PO-Flow when the true SCM is monotone:
+Proposition 4.3 (Encoded $Z$ as a function of the exogenous noise $U$).
+Let Assumption 4.1 hold.
+Assume that the exogenous noise $U\sim\mathrm{Unif}[0,1]$.
+Then there exists a continuously differentiable bijection function $\psi_{X}:\;\mathcal{U}\;\longrightarrow\;\mathcal{Z}$ that is independent of the treatment assignment $A$, i.e.,
+$\small Z\;=\;\Phi_{\theta}(Y^{(A)};X,A)\;=\;\Phi_{\theta}(f_{X}(A,U);X,A)\;=\;\psi_{X}\bigl(U\bigr)$ (12)
+Remark 4.4.
+We can transform the exogenous noise $U\sim\mathrm{Unif}[0,1]$ assumed in Proposition 12 into other distributions, such as Gaussian.
+If $Z\sim\mathcal{N}(0,1)$ with CDF $F(Z)$, then $U=F(Z)\sim\mathrm{Unif}[0,1]$, and any structural assignment $f(\cdot,U)$ can be written equivalently as $\tilde{f}(\cdot,Z)=f(\cdot,F(Z))$.
+Therefore, the conclusions of Proposition 12 also hold with other noise distributions.
+Following Proposition 12, we have the following result on counterfactual recovery under monotone SCMs:
+Corollary 4.5 (Counterfactual recovery under monotone SCMs).
+Assume the conditions in Assumption 4.1 hold.
+Consider a factual sample $\bigl(Y^{(A)},X,A\bigr)$ given by the structural equation $Y^{(A)}=f_{X}\bigl(A,U\bigr)$, and let its encoded representation be $Z=\Phi_{\theta}(Y^{(A)};X,A)$.
+Apply the intervention $\mathrm{do}\bigl(A=\gamma\bigr)$.
+Assume that the true counterfactual is given by $Y_{\mathrm{CF}}^{(\gamma)}=f_{X}(\gamma,U)$.
+The decoder then recovers the counterfactual almost surely:
+$\hat{Y}_{\mathrm{CF}}^{\gamma}:=\Phi_{\theta}^{-1}\bigl(Z;\ X,\gamma\bigr)=Y_{\mathrm{CF}}^{\gamma}.$ (13)
+Remark 4.6.
+One may notice that Assumptions (A1) and (A2) mirror those in Bijective Generation Mechanisms (BGM) [43].
+BGM establishes class-level identifiability: if a learned mechanism lies in the BGM class and matches the observational distribution ($\hat{f}_{X}(A,U)\stackrel{{\scriptstyle d}}{{=}}f_{X}(A,U)$), then counterfactuals are identifiable.
+By contrast, our Prop.
+4.3 shows that PO-Flow’s encoder is bijective in $U$, placing our model within the BGM class.
+We then use an alternative constructive route (building on the intermediate step in Prop.
+4.3) to obtain Cor.
+4.5 under (A1)–(A3).
+Alternatively, one could invoke the BGM result after Prop.
+4.3 to reach Cor.
+4.5; see Appendix C.
 4.3 PO-Flow for Learning Potential Outcomes Densities
-Another distinctive advantage of PO-Flow is its ability to learn the density of the Potential Outcomes (POs).
-We define $\hat{p}_{Y}(\cdot\mid x,a,y_{1})$ as the PO density learned by PO-Flow, conditioned on $(x,a)$ and a base sample $y_{1}$.
-The explicit form of the learned PO density is presented in Proposition 4.2, with the proof provided in Appendix A.
+Another advantage of PO-Flow is its ability to learn the density of the Potential Outcomes (POs).
+We define $p_{\theta,Y|X,A}(\cdot\mid x,a,z)$ as the PO density computed by PO-Flow, conditioned on $(x,a)$ and a base sample $z$.
+The explicit form of this density, mapping from latent to data space, is presented in Proposition 4.7, with the proof provided in Appendix A.
 Proposition 4.7.
-Given a noise sample $y_{1}\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
-$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=\log q(y_{1})+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)ds.$ (12)
-Empirically, the divergence of the velocity field can be computed under the Hutchinson trace estimator (e.g., [31, 64]). More details are provided in Appendix B.5.
-$\small\nabla\cdot v_{\theta}(y(t),t,x,a)\approx\mathbb{E}_{p(\epsilon)}\left[\epsilon\cdot\frac{v_{\theta}(y(t)+\sigma\epsilon,t,x,a)-v_{\theta}(y(t),t,x,a)}{\sigma}\right],$ (13)
-where $p(\epsilon)$ is a distribution in $\mathbb{R}^{d_{Y}}$ satisfying $\mathbb{E}[\epsilon]=0$ and Cov$(\epsilon)=I$ (e.g., a standard Gaussian).
-To our knowledge, PO-Flow is the first generative approach in causal inference to explicitly learn the probability density of predicted potential outcomes.
-Unlike prior works [44, 45] that generate PO samples and estimate empirical distributions, PO-Flow provides direct access to likelihoods, enabling deeper insights into uncertainty.
-This density estimation allows PO-Flow to assess prediction confidence via log-likelihood, identify unlikely outcomes, and quantify uncertainty.
-For counterfactual prediction, one can perturb the latent code $z_{y}^{\prime}$ multiple times and select the most confident $\hat{y}_{\text{cf}}^{(1-a)}$ based on log-likelihood, effectively reducing variance while improving robustness.
+Given a noise sample $z\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
+$\log p_{\theta,Y\mid X,A}\bigl(y(0)\mid x,a,z\bigr)=\log q(z)\;+\;\int_{0}^{1}\nabla_{y}\!\cdot\,v_{\theta}\bigl(y(t),t;\,x,a\bigr)\,dt.$ (14)
+Empirically, the divergence of the velocity field can be computed much more efficiently under the Hutchinson trace estimator (e.g., [27, 58]). More details are provided in Appendix B.4.
+Unlike prior works [40, 41] that generate PO samples and approximate empirical distributions, PO-Flow provides direct likelihood evaluation.
+This allows principled uncertainty quantification, supports ranking outcomes by plausibility (Table 4), enables information-theoretic analyses such as KL divergence (Table 3), and facilitates tasks like anomaly detection via quantile-based thresholds.
 5 Experiments
-Baselines: Our baselines cover several categories: model-agnostic learners that learn the conditional distribution means (S-learner and T-learner [37]), representation-based conditional regressors (Counterfactual Regression (CFR) [55]), generative approaches (Causal Effect Variational Autoencoder (CEVAE) [42]; GAN for Individualized Treatment Effect (GANITE) [67]), and most recent generative approaches based on diffusion (DiffPO [43]) and discrete normalizing flows (INFs [45]).
+Baselines: Our baselines cover several categories: model-agnostic learners that learn the conditional distribution means (S-learner [33], T-learner [33], and DR-Learner [32]), representation-based conditional regressors (Counterfactual Regression (CFR) [50]), deep generative methods (Causal Effect Variational Autoencoder (CEVAE) [38]; GAN for Individualized Treatment Effect (GANITE) [61]), and most recent generative approaches based on diffusion (DiffPO [39]) and discrete normalizing flows (INFs [41]).
 Datasets: Counterfactual outcomes and unobserved potential outcomes (POs) are inaccessible in real-world data but can be simulated in semi-synthetic settings, as discussed in detail in Appendix B.1.
-Accordingly, we evaluate PO-Flow on standard synthetic benchmarks, including ACIC 2016, ACIC 2018, IHDP, and IBM.
-The code and datasets are available at https://github.com/StatFusion/PO-Flow.
+Accordingly, we evaluate PO-Flow on synthetic benchmarks, including ACIC 2016, ACIC 2018, IHDP, and IBM.
+The code and datasets are included in the supplementary materials for anonymous submission.
 5.1 Prediction of Potential Outcomes
@@
 $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$
-[…]
+PO-Flow $\mathbf{0.53_{\pm.10}}$ $0.60_{\pm.12}$ $\mathbf{0.98_{\pm.11}}$ $\mathbf{1.19_{\pm.13}}$ $\mathbf{0.90_{\pm.09}}$ $\mathbf{0.97_{\pm.11}}$
 DiffPO $0.72_{\pm.14}$ $0.79_{\pm.16}$ $1.33_{\pm.20}$ $1.40_{\pm.22}$ $1.82_{\pm.38}$ $1.85_{\pm.35}$
 INFs $0.69_{\pm.16}$ $0.78_{\pm.17}$ $1.02_{\pm.10}$ $\mathbf{1.20_{\pm.12}}$ $1.52_{\pm.24}$ $1.57_{\pm.28}$
-S-learner $\mathbf{0.54_{\pm.09}}$ $\mathbf{0.57_{\pm.11}}$ $1.31_{\pm.18}$ $1.44_{\pm.20}$ $1.01_{\pm.17}$ $1.16_{\pm.19}$
+S-learner $\mathbf{0.54_{\pm.09}}$ $\mathbf{0.57_{\pm.11}}$ $1.31_{\pm.18}$ $1.44_{\pm.20}$ $01_{\pm.17} 1.1$$16_{\pm.19}$
 T-learner $1.57_{\pm.32}$ $1.71_{\pm.40}$ $1.45_{\pm.25}$ $1.49_{\pm.27}$ $1.89_{\pm.47}$ $1.96_{\pm.50}$
 CEVAE $0.83_{\pm.17}$ $0.85_{\pm.17}$ $1.18_{\pm.15}$ $1.36_{\pm.18}$ $0.96_{\pm.12}$ $\mathbf{0.98_{\pm.12}}$
-CFR $0.94_{\pm.10}$ $0.98_{\pm.12}$ $1.10_{\pm{.20}}$ $\mathbf{1.18_{\pm{.20}}}$ $2.09_{\pm.45}$ $2.17_{\pm.48}$
+[…]
 GANITE $0.88_{\pm.15}$ $0.97_{\pm.15}$ $1.60_{\pm.36}$ $1.67_{\pm.34}$ $2.48_{\pm.39}$ $2.59_{\pm.47}$
-Evaluation Metrics: We evaluate the point estimates of potential outcomes (POs) using the root mean squared error (RMSE) between the estimated outcomes $\hat{y}_{i}$ and the true outcomes $y_{i}$, computed as $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$.
-$\text{RMSE}_{\text{in}}$ and $\text{RMSE}_{\text{out}}$ refer to the RMSE evaluated on the training and test sets, respectively.
+Evaluation Metrics: We evaluate the point estimates of potential outcomes (POs) using the root mean squared error (RMSE) between the estimated outcomes $\hat{y}_{i}^{(a)}$ and the true outcomes $y_{i}^{(a)}$, computed as $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}^{(a)}-\hat{y}_{i}^{(a)})^{2}}$, where $\text{RMSE}_{\text{in}}$ and $\text{RMSE}_{\text{out}}$ refer to the RMSE evaluated on the training and test sets, respectively.
 Results: Table 1 shows that PO-Flow achieves the most accurate prediction of potential outcomes on the majority of datasets.
+Besides, we also report in Table 4 the RMSE of the most likely POs selected by PO-Flow using the highest log-density (as computed in Proposition 4.7).
 5.2 Estimation of CATEs
-Conditional Average Treatment Effect (CATE): The CATE is defined as $\tau(x)=\mathbb{E}[y_{1}-y_{0}\mid X=x]$, representing the expected treatment effect conditioned on covariates $x$.
+Conditional Average Treatment Effect (CATE): The CATE is defined as $\tau(x)=\mathbb{E}[y^{(a=1)}-y^{(a=0)}\mid X=x]$, representing the expected treatment effect conditioned on covariates $x$.
 It captures individual-level heterogeneity in causal effects.
@@
 Table 2: Root Precision in Estimation of Heterogeneous Effect ($\sqrt{\epsilon_{\text{PEHE}}}$) across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets.
-Results are averaged over 10-fold cross-validation.
 ACIC 2016 ACIC 2018 IHDP IBM
 $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$
+PO-Flow $\mathbf{0.76_{\pm.08}}$ $\mathbf{0.82_{\pm.09}}$ $\mathbf{0.05_{\pm.01}}$ $\mathbf{0.07_{\pm.02}}$ $\mathbf{0.41_{\pm.08}}$ $\mathbf{0.45_{\pm.10}}$ $\mathbf{0.04_{\pm.01}}$ $\mathbf{0.04_{\pm.01}}$
 […]
-DiffPO $0.91_{\pm.20}$ $1.12_{\pm.26}$ $0.07_{\pm.03}$ $\mathbf{0.08_{\pm.04}}$ $0.79_{\pm.16}$ $0.84_{\pm.18}$ $0.08_{\pm.03}$ $0.09_{\pm.03}$
 INFs $1.37_{\pm.28}$ $1.42_{\pm.30}$ $0.33_{\pm.11}$ $0.37_{\pm.12}$ $0.90_{\pm.27}$ $0.95_{\pm.27}$ $0.16_{\pm.08}$ $0.18_{\pm.08}$
-S-learner $3.81_{\pm.76}$ $3.87_{\pm.98}$ $\mathbf{0.06_{\pm.02}}$ $\mathbf{0.07_{\pm.03}}$ $0.92_{\pm.07}$ $0.93_{\pm.08}$ $0.06_{\pm.02}$ $0.07_{\pm.04}$
+S-learner $3.81_{\pm.76}$ $3.87_{\pm.98}$ $\mathbf{0.06_{\pm.02}}$ $\mathbf{0.07_{\pm.03}}$ $0.92_{\pm.07}$ $93_{\pm.08} 0.0$$6_{\pm.02} 0.0$$07_{\pm.04}$
 T-learner $5.02_{\pm.71}$ $5.15_{\pm.87}$ $1.21_{\pm.47}$ $1.32_{\pm.53}$ $1.24_{\pm.08}$ $1.31_{\pm.09}$ $0.60_{\pm.14}$ $0.69_{\pm.15}$
+DR-learner $3.04_{\pm.19}$ $3.51_{\pm.23}$ $0.43_{\pm.08}$ $0.57_{\pm.10}$ $1.03_{\pm.09}$ $1.28_{\pm.15}$ $0.12_{\pm.03}$ $0.19_{\pm.05}$
 CEVAE $2.06_{\pm.61}$ $2.31_{\pm.74}$ $1.19_{\pm.15}$ $1.21_{\pm.16}$ $1.41_{\pm.10}$ $1.43_{\pm.10}$ $0.73_{\pm.15}$ $0.75_{\pm.16}$
@@
 5.3 Learning the Full Potential Outcome Distributions
-Beyond point estimation, PO-Flow offers two distinct advantages for modeling potential outcomes: (1) it enables individualized sampling from the conditional distribution $P(Y\mid X,A)$, generating a rich set of plausible outcomes for each individual; and (2) it supports learning the probability density of any generated outcome, providing deeper insights into predictive uncertainty.
+Beyond point estimation, PO-Flow offers two distinct advantages for modeling potential outcomes: (1) it enables individualized sampling from the conditional distribution $P(Y\mid X,A)$, generating a rich set of plausible outcomes for each individual; and (2) it supports learning the density of any generated outcome, providing deeper insights into predictive uncertainty.
 5.3.1 Learning potential outcome densities
 PO-Flow enables direct computation of potential outcome densities, as shown in Proposition 4.7.
-To our knowledge, it is the first generative method in causal potential outcomes to provide this capability.
-Evaluation Metrics: To evaluate the learned densities, we compute the KL divergence between the PO-Flow estimate $\hat{p}(\cdot|x,a)$ and the ground-truth density $p_{\text{true}}(\cdot|x,a)$.
-Given that the semi-synthetic datasets assume $y^{(a)}\sim\mathcal{N}(\mu(a),1)$, the individual-level KL is $\text{KL}(\hat{Y}_{X}^{(a)}\|Y_{X}^{(a)})=\mathbb{E}_{\hat{y}\sim\hat{p}}\left[\log\frac{\hat{p}(\hat{y}|x,a)}{\mathcal{N}(\mu(\hat{y}),1)}\right]$, where $\hat{p}$ is computed via (12).
-Table 3 reports the average results.
+To evaluate the learned densities, we compute the KL divergence between the PO-Flow estimate $p_{\theta,Y|X,A}(\cdot|x,a)$ and the ground-truth density $p_{\text{true}}(\cdot|x,a)$.
+Given that the semi-synthetic datasets assume $y^{(a)}\sim\mathcal{N}(\mu(a),1)$, the individual-level KL is $\text{KL}(\hat{Y}_{X}^{(a)}\|Y_{X}^{(a)})=\mathbb{E}_{\hat{y}\sim p_{\theta,Y|X,A}}\left[\log\frac{p_{\theta,Y|X,A}(\hat{y}|x,a)}{\mathcal{N}(\hat{y};\mu(a),1)}\right]$, where $p_{\theta,Y\mid X,A}$ is computed via (14).
 Table 3: KL divergence of learned potential outcome distributions across datasets. Results are averaged over 10-fold cross-validation.
@@
 Results: PO-Flow achieves much lower KL divergence than GP baselines on the ACIC 2018, IHDP, and IBM datasets, indicating more accurate density estimation of potential outcomes.
-This property allows PO-Flow to assess prediction confidence via log-density and to identify unlikely outcomes.
-For counterfactual estimation, one can sample multiple candidates from a perturbed latent region and select the most likely one based on the model’s estimated density.
+This property enables PO-Flow to identify most likely potential outcomes based on their log-densities.
+Table 4 reports the RMSE of POs selected by PO-Flow with the highest log-density, illustrating one application of this property.
+Table 4: Root mean squared error (RMSE) of estimated potential outcomes (POs) across different methods on the ACIC 2018, IHDP, and IBM datasets.
+For PO-Flow, results are based on the most likely PO selected from 100 simulated outcomes using log-density.
+ACIC 2018 IHDP IBM
+$\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$
+PO-Flow $\mathbf{0.42_{\pm.07}}$ $\mathbf{0.54_{\pm.10}}$ $\mathbf{0.96_{\pm.09}}$ $\mathbf{1.05_{\pm.10}}$ $\mathbf{0.84_{\pm.06}}$ $\mathbf{0.90_{\pm.08}}$
+DiffPO $0.72_{\pm.14}$ $0.79_{\pm.16}$ $1.33_{\pm.20}$ $1.40_{\pm.22}$ $1.82_{\pm.38}$ $1.85_{\pm.35}$
+INFs $0.69_{\pm.16}$ $0.78_{\pm.17}$ $1.02_{\pm.10}$ $1.20_{\pm.12}$ $1.52_{\pm.24}$ $1.57_{\pm.28}$
+S-learner $0.54_{\pm.09}$ $0.57_{\pm.11}$ $1.31_{\pm.18}$ $1.44_{\pm.20}$ $01_{\pm.17} 1.1$$16_{\pm.19}$
+T-learner $1.57_{\pm.32}$ $1.71_{\pm.40}$ $1.45_{\pm.25}$ $1.49_{\pm.27}$ $1.89_{\pm.47}$ $1.96_{\pm.50}$
+CEVAE $0.83_{\pm.17}$ $0.85_{\pm.17}$ $1.18_{\pm.15}$ $1.36_{\pm.18}$ $0.96_{\pm.12}$ $0.98_{\pm.12}$
+CFR $0.94_{\pm.10}$ $0.98_{\pm.12}$ $10_{\pm{.20}} 1.1$$18_{\pm{.20}}$ $2.09_{\pm.45}$ $2.17_{\pm.48}$
+GANITE $0.88_{\pm.15}$ $0.97_{\pm.15}$ $1.60_{\pm.36}$ $1.67_{\pm.34}$ $2.48_{\pm.39}$ $2.59_{\pm.47}$
+Compared with the averaged RMSE over all POs generated by PO-Flow in Table 1, selecting the most likely POs based on log-density yields smaller RMSEs, providing more confident and closer estimates.
 5.3.2 Empirical distribution estimations
 Evaluation Metrics: To enable comparison of estimated potential outcome distributions across methods, we include empirical distribution estimates for all baselines.
-These are evaluated using the Wasserstein-1 distance, defined as: $W_{1}(\hat{p}(y|x,a),p(y|x,a))=\inf_{\gamma\in\Gamma(\hat{p},p)}\mathbb{E}_{(\hat{y},y\sim\gamma)}[\|\hat{y}-y\|_{1}]$.
+These are evaluated using the Wasserstein-1 distance, defined as: $W_{1}(p_{\theta,Y|X,A}(y|x,a),p_{Y|X,A}(y|x,a))=\inf_{\gamma\in\Gamma(p_{\theta},p)}\mathbb{E}_{(\hat{y},y\sim\gamma)}[\|\hat{y}-y\|_{1}]$.
 The results are shown in Table 5.
@@
 $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$
+PO-Flow $\mathbf{0.36_{\pm.11}}$ $\mathbf{0.42_{\pm.12}}$ $\mathbf{0.05_{\pm.03}}$ $\mathbf{0.12_{\pm.04}}$ $\mathbf{0.30_{\pm.05}}$ $\mathbf{0.41_{\pm.06}}$ $\mathbf{0.33_{\pm.07}}$ $\mathbf{0.44_{\pm.08}}$
 […]
-DiffPO $0.47_{\pm.17}$ $0.51_{\pm.18}$ $\mathbf{0.04_{\pm.02}}$ $\mathbf{0.13_{\pm.05}}$ $0.80_{\pm.15}$ $0.88_{\pm.17}$ $0.69_{\pm.14}$ $0.75_{\pm.14}$
-INFs $0.59_{\pm.13}$ $0.67_{\pm.14}$ $0.08_{\pm.03}$ $0.16_{\pm.05}$ $0.32_{\pm.08}$ $\mathbf{0.40_{\pm.09}}$ $0.36_{\pm.08}$ $\mathbf{0.45_{\pm.09}}$
+INFs $0.59_{\pm.13}$ $67_{\pm.14} 0.0$$08_{\pm.03}$ $0.16_{\pm.05}$ $0.32_{\pm.08}$ $\mathbf{0.40_{\pm.09}}$ $0.36_{\pm.08}$ $\mathbf{0.45_{\pm.09}}$
 S-learner $1.01_{\pm.13}$ $1.02_{\pm.20}$ $0.35_{\pm.13}$ $0.49_{\pm.20}$ $0.71_{\pm.14}$ $0.85_{\pm.19}$ $0.46_{\pm.11}$ $0.55_{\pm.14}$
@@
 GANITE $0.74_{\pm.20}$ $0.86_{\pm.23}$ $0.89_{\pm.24}$ $1.18_{\pm.25}$ $0.71_{\pm.23}$ $0.82_{\pm.29}$ $1.09_{\pm.27}$ $1.34_{\pm.30}$
-Figure 2: Comparison between empirically estimated and the ground-truth POs distributions. Left: $P(y|x,a=0)$ on IHDP. Right: $P(y|x,a=1)$ on IHDP.
-Understanding the distribution of potential outcomes (POs) is essential for reasoning about uncertainty in treatment decisions.
-As shown in Figure 2, PO distributions can shift differently across treatments for the same individual.
-PO-Flow provides not only point estimates but also distributional insight, supporting uncertainty-aware analysis.
-Section 5.3.1 further demonstrates its ability to estimate densities accurately.
-This enables practitioners to assess whether an outcome lies within a desired range and evaluate its likelihood (Proposition 4.2) under a given treatment, offering a clearer sense of confidence in the effect.
+Understanding the distribution of potential outcomes (POs) is essential for uncertainty-aware treatment decisions, as outcomes for the same individual may vary substantially across treatments.
+PO-Flow provides not only point estimates but also full distributional insight, enabling practitioners to assess whether outcomes fall within desired ranges and to evaluate their likelihood (Proposition 4.7).
 5.4 Counterfactual Outcome Estimation
-Evaluation Metrics: We evaluate the predicted counterfactual outcomes using the root mean squared error (RMSE) between the estimated counterfactuals $\hat{y}_{i,\text{cf}}^{(1-a)}$ and the ground-truth counterfactuals $y_{i,\text{cf}}^{(1-a)}$, computed as $\text{RMSE}=\small\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i,\text{cf}}^{(1-a)}-\hat{y}_{i,\text{cf}}^{(1-a)})^{2}}$.
-The results are presented in Table 5.
-Here, $\text{RMSE}_{\text{in}}$ and $\text{RMSE}_{\text{out}}$ denote the RMSE evaluated on the training and test sets, respectively.
-Baselines: To our knowledge, very few existing methods are specifically designed for counterfactual outcome prediction conditioned on factuals in general settings—without relying on structural causal models or temporal information.
+Evaluation Metrics: We assess counterfactual predictions using the RMSE between model-estimated $\hat{y}_{\text{cf}}^{(1-a)}$ and ground-truth $y_{\text{cf}}^{(1-a)}$. Results are in Table 6.
+Baselines: To our knowledge, very few existing methods are specifically designed for counterfactual outcome prediction conditioned on factuals in general settings—without explicitly modeling structural causal models or relying on temporal information.
 While methods like CFR include “counterfactual” in name, they target unobserved potential outcomes rather than factual-conditioned ones, though we still include them as valuable baselines.
-We also adapt DiffPO and GANITE for this task, as detailed in Appendix B.6.
+We also adapted DiffPO and GANITE for this task, as detailed in Appendix B.5.
 Table 6: Root mean squared error (RMSE) of estimated counterfactual outcomes across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets.
@@
 $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$
-[…]
+PO-Flow $\mathbf{1.50_{\pm.15}}$ $\mathbf{1.53_{\pm.17}}$ $\mathbf{0.04_{\pm.01}}$ $\mathbf{0.05_{\pm.02}}$ $\mathbf{1.63_{\pm.09}}$ $1.69_{\pm.10}$ $\mathbf{0.04_{\pm.01}}$ $\mathbf{0.04_{\pm.02}}$
 DiffPO $1.77_{\pm.20}$ $1.78_{\pm.20}$ $0.62_{\pm.14}$ $0.65_{\pm.15}$ $1.81_{\pm.22}$ $1.85_{\pm.23}$ $0.62_{\pm.13}$ $0.64_{\pm.16}$
-CFR $3.15_{\pm.28}$ $3.38_{\pm.30}$ $0.99_{\pm.05}$ $1.01_{\pm.07}$ $\mathbf{1.63_{\pm.21}}$ $\mathbf{1.65_{\pm.25}}$ $0.86_{\pm.18}$ $0.87_{\pm.16}$
+CFR $15_{\pm.28} 3.3$$38_{\pm.30}$ $0.99_{\pm.05}$ $1.01_{\pm.07}$ $\mathbf{1.63_{\pm.21}}$ $\mathbf{1.65_{\pm.25}}$ $0.86_{\pm.18}$ $0.87_{\pm.16}$
 GANITE $3.41_{\pm.52}$ $3.67_{\pm.55}$ $0.68_{\pm.13}$ $0.70_{\pm.14}$ $2.48_{\pm.53}$ $2.72_{\pm.59}$ $1.84_{\pm.30}$ $2.02_{\pm.37}$
-5.5 Counterfactual Image Generation
-Figure 3: PO-Flow counterfactual image generation from female to male faces.
-$(a)$: Each row shows a factual outcome $y^{(a)}$ (1st column), its encoded latent $z_{y}$ (2nd column), and reverse-time decoding at $t=2/3$, $t=1/3$, and $t=0$ (3rd–5th columns), yielding the counterfactual $\hat{y}_{\text{cf}}^{(1-a)}$.
-$(b)$: Counterfactuals decoded from perturbed latents $z_{y}+\delta$, with $\delta\sim\mathcal{N}(0,0.2I)$.
-Beyond low-dimensional datasets, PO-Flow also succeeds on high-dimensional counterfactual image generation.
-We train PO-Flow on the latent space of a pre-trained VAE using the CelebA dataset, predicting counterfactual male faces from original female faces.
-In this setting, $X\in\mathbb{R}^{6}$ encodes visual attributes (hair, eyebrows, bangs, lips), $a\in\{\text{male},\text{female}\}$ denotes gender, and $y\in\mathbb{R}^{512}$ is the VAE latent feature.
-Visual results are shown in Figures 3 and 6; detailed settings are in Appendix B.2.
 5.5 Discussions and Comparisons
-Training efficiency: PO-Flow exhibits high training efficiency, converging in approximately 300 gradient steps across multiple causal datasets, as shown in Appendix C.1.
-Minimal architecture sensitivity: We find that PO-Flow is the least sensitive to neural network architecture and hyperparameters.
-Using the same NN structure across all causal datasets, PO-Flow consistently achieves state-of-the-art performance.
-In contrast, we find that generative approaches including DiffPO, INFs, and GANITE require careful hyperparameter tuning for each dataset.
-Unique density estimation: Our continuous normalizing flow (CNF) backbone enables density learning of potential outcomes, as shown in Proposition 4.2 and Section 5.3.1, which, to our knowledge, is the first generative model in the field to achieve this.
+Training efficiency: PO-Flow exhibits high training efficiency, converging in approximately 300 gradient steps across multiple causal datasets (Appendix D.3), and demonstrates notable advantages in both training and sampling time compared to other baselines (Table 9).
+Density estimation for principled uncertainty quantification: Our continuous normalizing flow (CNF) backbone enables density learning of potential outcomes, as detailed in Section 5.3.1.
+This property allows us to generate multiple potential outcomes for each individual and to identify the most likely ones based on their log-densities.
+The empirical evaluation of this procedure is presented in Table 4.
 Counterfactual: We take a step further by exploring counterfactual outcomes, enabled by the smooth PO-Flow mapping that preserves factual information during encoding.
-Unlike diffusion-based DiffPO, which has noisy forward processes, PO-Flow offers information-preserving forward processes, as shown in Figure 1.
-To our knowledge, it is among the first to explore empirical counterfactual prediction without relying on temporal data or explicitly modeling structural causal models.
-However, one area for improving counterfactual prediction lies in modeling uncertainty more expressively.
-The current covariate-conditioned Gaussian perturbations may not fully capture counterfactual variation.
-Future work could explore more theoretically grounded counterfactual uncertainty.
+Under (A1)–(A3), we include a supporting recovery result for this architecture (Proposition 12; Corollary 4.5).
+It is among the first to explore counterfactual outcomes in the potential outcomes framework without relying on temporal data or explicitly modeling structural causal models.
 6 Conclusions
 We introduced PO-Flow, a continuous normalizing flow framework for modeling potential and counterfactual outcomes in causal inference.
-PO-Flow provides a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction—without explicitly modeling structural causal models (SCMs) or relying on auxiliary information for the counterfactual task.
-Empirically, PO-Flow achieves state-of-the-art performance on benchmark datasets across multiple causal tasks.
-Among generative models, it is the first to accurately learn the density of potential outcomes, enabling principled uncertainty quantification.
-Furthermore, PO-Flow is efficient, robust, and extendable to high-dimensional settings such as counterfactual image generation.
-These results highlight PO-Flow as a unified and scalable solution for causal inference tasks.
+PO-Flow provides a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction in potential outcomes—without explicitly modeling structural causal models (SCMs) or relying on auxiliary information.
+To our knowledge, this constitutes one of the earliest attempts to address counterfactual outcome prediction in the potential outcomes literature, a crucial yet underexplored direction.
+Theoretically, we include a supporting recovery result for the proposed architecture (Corollary 4.5).
+Empirically, PO-Flow attains state-of-the-art performance on benchmark datasets across standard causal tasks, while additionally supporting counterfactual outcome prediction and log-density estimation for principled uncertainty quantification.
+In addition, PO-Flow is both computationally efficient and robust compared to modern baselines.
+These results highlight PO-Flow as a unified and scalable solution for potential-outcomes–based causal tasks, supporting reliable decision-making in domains such as healthcare.
 Acknowledgment
@@
 Appendix A Proofs
-Proposition 4.1.
-Assume that $p_{t}(y|x,a)>0$ for all $y\in\mathbb{R}^{d_{Y}}$ and $t\in[0,1]$.
+Proposition A.1.
+Assume that $p(y|x,a)>0$ for all $y$ and $t\in[0,1]$.
 Then, up to a constant independent of $\theta$, the Conditional Flow Matching (CFM) loss and the original Flow Matching (FM) loss are equivalent.
@@
 Proof:
-We denote the conditional reference velocity field as $u(y,t,x,a|y_{0})=\frac{d\phi_{t}}{dt}=(1-\sigma_{\text{min}})y_{1}-y_{0}$, as defined in (6).
+We denote the conditional reference velocity field as $u(y,t,x,a|y_{0})=\frac{d\phi}{dt}=(1-\sigma_{\text{min}})y_{1}-y_{0}$, as defined in (6).
 We further define the marginal (unconditional) reference velocity field as:
-[…]
+$u(y,t,x,a)=\int u(y,t,x,a|y_{0})\frac{p(y,t,x,a|y_{0})p_{Y}(y_{0})}{p(y,t,x,a)}dy_{0},$ (15)
 where $p(y,t,x,a|y_{0})$ and $p(y,t,x,a)$ denote the probability densities induced by $u(y,t,x,a|y_{0})$ and $u(y,t,x,a)$, respectively.
 Next, the original flow matching loss, unconditioned on $y_{0}$, is defined as:
-$\mathcal{L}_{\text{FM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y_{t},X,A}(\cdot)}\|v_{\theta}(y_{t},t,x,a)-u(y_{t},t,x,a)\|^{2}.$ (15)
-In the main text, we chose $\phi_{t}$ as the linear interpolant between $y_{0}$ and $y_{1}$, which results in a constant reference velocity $\frac{d\phi_{t}}{dt}$ independent of $t$.
+$\mathcal{L}_{\text{FM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot)}\|v_{\theta}(y,t,x,a)-u(y,t,x,a)\|^{2}.$ (16)
+In the main text, we chose $\phi$ as the linear interpolant between $y_{0}$ and $y_{1}$, which results in a constant reference velocity $\frac{d\phi}{dt}$ independent of $t$.
 This leads to the simplified CFM loss presented in Eq.
@@
 However, the general CFM loss, corresponding to an arbitrary reference velocity field $u(y,t\mid y_{0})$ conditioned on $y_{0}$, is defined as:
-$\mathcal{L}_{\text{CFM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})}\|v_{\theta}(\cdot)-u_{t}(\cdot|y_{0})\|^{2}$ (16)
+$\mathcal{L}_{\text{CFM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y,X,A|Y_{0}}(\cdot|y_{0})}\|v_{\theta}(\cdot)-u(\cdot|y_{0})\|^{2}$ (17)
 Next, we have that:
@@
 $\|v_{\theta}(y,t,x,a)-u(y,t,x,a|y_{0})\|^{2}$ $=\|v_{\theta}(\cdot)\|^{2}-2\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle+\|u(\cdot|y_{0})\|^{2}.$
-Besides, the term $\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y_{t},X,A}(\cdot)}\|v_{\theta}(\cdot)\|^{2}$ can be equivalently expressed in conditional form as:
-[…]
-$=\int\|v_{\theta}(\cdot)\|^{2}p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\,dy\,dx\,da\,dt$
-$=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})}\|v_{\theta}(\cdot)\|^{2}.$
+Besides, the term $\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot)}\|v_{\theta}(\cdot)\|^{2}$ can be equivalently expressed in conditional form as:
+$\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot)}\|v_{\theta}(\cdot)\|^{2}$ $=\int\|v_{\theta}(\cdot)\|^{2}p_{Y,X,A}(\cdot)\,dy\,dx\,da\,dt$ (19)
+$=\int\|v_{\theta}(\cdot)\|^{2}p_{Y,X,A|Y_{0}}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\,dy\,dx\,da\,dt$
+$=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y,X,A|Y_{0}}(\cdot|y_{0})}\|v_{\theta}(\cdot)\|^{2}.$
 Furthermore, by the law of total expectation, we have:
-[…]
-$=\int\left\langle v_{\theta}(\cdot),\int u(\cdot|y_{0})p_{Y_{t},X,A}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\right\rangle\,dy\,dx\,da\,dt$
-$=\int\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle p_{Y_{t},X,A}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\,dy\,dx\,da\,dt$
-$=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})}\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle.$
-Finally, because $u_{t}(\cdot\mid y_{0})$ does not depend on $\theta$, substituting Eqs. (18) and (19) into Eq. (15) and differentiating with respect to $\theta$ yields:
+$\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot)}\langle v_{\theta}(\cdot),u(\cdot)\rangle$ $=\int\left\langle v_{\theta}(\cdot),\frac{\int u(\cdot|y_{0})p_{Y,X,A}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}}{p_{Y,X,A}(\cdot)}\right\rangle p_{Y,X,A}(\cdot)\,dy\,dx\,da\,dt$ (20)
+$=\int\left\langle v_{\theta}(\cdot),\int u(\cdot|y_{0})p_{Y,X,A}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\right\rangle\,dy\,dx\,da\,dt$
+$=\int\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle p_{Y,X,A}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\,dy\,dx\,da\,dt$
+$=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y,X,A|Y_{0}}(\cdot|y_{0})}\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle.$
+Finally, because $u(\cdot\mid y_{0})$ does not depend on $\theta$, substituting Eqs. (19) and (20) into Eq. (16) and differentiating with respect to $\theta$ yields:
 $\nabla_{\theta}\mathcal{L}_{\text{FM}}(\theta)=\nabla_{\theta}\mathcal{L}_{\text{CFM}}(\theta).$ (21)
-Proposition 4.2. Given a noise sample $y_{1}\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
-$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=\log q(y_{1})+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)dt.$
+Proposition 12.
+(Encoded $Z$ as a function of the exogenous noise $U$) Let Assumption 4.1 hold.
+Assume that the exogenous noise $U\sim\mathrm{Unif}[0,1]$.
+Then there exists a continuously differentiable bijection function $\psi_{X}:\;\mathcal{U}\;\longrightarrow\;\mathcal{Z}$ that is independent of the treatment assignment $A$, i.e.,
+$Z\;=\;\Phi_{\theta}(Y^{(A)};\ X,A)\;=\;\Phi_{\theta}(f_{X}(A,U);X,A)\;=\;\psi_{X}\bigl(U\bigr)\quad\text{a.s.}$ (22)
+Proof. Based on the structural causal model, we have $Y^{(A)}=f_{X}(A,U)$. We define
+$m_{X,A}(U)\;:=Z=\Phi_{\theta}(Y^{(A)};\,X,A)=\;\Phi_{\theta}\bigl(f_{X}(A,U);\ X,A\bigr),$
+as a function that may depend on the action $A$. Therefore, our target becomes proving that $m_{X,A}(U)$ is functionally invariant to $A$.
+By (A3) in Assumption 4.1, we have $Z\perp\!\!\!\perp A\ |\ X$. Therefore,
+$p_{\theta,Z\mid X,A}(z)=p_{\theta,Z}(z).$ (23)
+In continuous normalizing flows, $\Phi_{\theta}$ is invertible with respect to $Y^{(A)}$ due to the invertibility of the underlying neural ODE.
+In 1-D, this implies that $\Phi_{\theta}$ is monotonic in $Y^{(A)}=f_{X}(A,U)$.
+Without loss of generality, we assume that $\Phi_{\theta}$ is strictly increasing in $f_{X}$.
+Moreover, by (A2), since $f_{X}(\cdot,U)$ is strictly monotone in $U$.
+Without loss of generality, we assume strictly increasing.
+Then it follows by the composition rule that $m_{X,A}(U)$ is strictly increasing in $U$ and hence bijective on $[0,1]$.
+Since $Z=\Phi_{\theta}(Y^{(A)};\,X,A)=m_{X,A}(U)$, we may apply change of variables formula w.r.t. $U$:
+$p_{\theta,Z|X,A}(z)\;=\;p_{U}\!\bigl(m_{X,A}^{-1}(z)\bigr)\;\left|\frac{d}{dz}m_{X,A}^{-1}\!\bigl(z\bigr)\right|=1\cdot\frac{d}{dz}m_{X,A}^{-1}\!\bigl(z\bigr),$ (24)
+where the last equation follows from the uniform distribution of $U$ and the fact that $p_{Z}>0$.
+Then based on (23), we have:
+$p_{\theta,Z|X,A=1}(z)=\frac{d}{dz}m_{X,1}^{-1}\!\bigl(z\bigr)=\frac{d}{dz}m_{X,0}^{-1}\!\bigl(z\bigr)=p_{\theta,Z|X,A=0}(z).$ (25)
+It follows that:
+$m_{X,A}^{-1}(z)=\int^{z}\frac{d}{dt}m_{X,A}^{-1}\!\bigl(t\bigr)dt=\int^{z}c(t)dt+c_{A},$ (26)
+where $c(t)=\frac{d}{dt}m_{X,1}^{-1}\!\bigl(t\bigr)=\frac{d}{dt}m_{X,0}^{-1}\!\bigl(t\bigr)$ is independent of $A$, and $c_{A}$ is a constant for each $A$.
+By re-inverting (26), we have:
+$m_{X,A}(u)=(m_{X,A}^{-1})^{-1}(u)=\Bigl(\underbrace{\int^{\bullet}c(t)\,\mathrm{d}t}_{:=H(\bullet)}+c_{A}\Bigr)^{-1}(u)=H^{-1}\bigl(u-c_{A}\bigr).$ (27)
+Since $m_{X,A}$ is a bijection $0,1]\to\text{supp}(Z)=\{z\in\mathbb{R}:p_{Z}(z)>0\}$, we have:
+$m_{X,A}(0)=\inf\ \text{supp}(Z),\ \ m_{X,A}(1)=\sup\ \text{supp}(Z),\ \text{for both}\ A.$ (28)
+Besides, since the support of $Z$ does not depend on $A$ because of (23), we have that
+$\inf\ \text{supp}(Z)=m_{X,0}(0)=H^{-1}(-c_{0})=H^{-1}(-c_{1})=m_{X,1}(0).$ (29)
+Therefore, $c_{0}=c_{1}:=c$ is a constant that does not depend on $A$.
+Therefore, we can write
+$\psi_{X}(U)\;:=\;H^{-1}\bigl(U-c\bigr)\;=\;m_{X,0}(U)\;=\;m_{X,1}(U).$ (30)
+As a result, we conclude that
+$Z\;=\;\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr)\;=\;\psi_{X}\bigl(U\bigr)\quad\text{a.s.},$ (31)
+is a function that only depends on the exogenous noise $U$ (and covariates $X$), and is invariant to the treatment $A$.
+Corollary 4.5 (Counterfactual recovery) Assume the conditions in Assumption 4.1 hold.
+Consider a factual sample $\bigl(Y^{(A)},X,A\bigr)$ given by the structural equation $Y^{(A)}=f_{X}\bigl(A,U\bigr)$, and let its encoded representation be $Z=\Phi_{\theta}(Y^{(A)};X,A)$.
+Apply the intervention $\mathrm{do}\bigl(A=\gamma\bigr)$.
+Assume that the true counterfactual is given by $Y_{\mathrm{CF}}^{(\gamma)}=f_{X}(\gamma,U)$.
+The decoder then recovers the counterfactual almost surely:
+$\hat{Y}_{\mathrm{CF}}^{\gamma}:=\Phi_{\theta}^{-1}\bigl(Z;\ X,\gamma\bigr)=Y_{\mathrm{CF}}^{\gamma}.$ (32)
+Proof. Due to the deterministic and invertible nature of the continuous normalizing flow when the conditioning input is fixed, we have,
+$\Phi_{\theta}^{-1}(\Phi_{\theta}(Y;\,X,A);X,A)=Y.$ (33)
+From Proposition 4.3, we have $Z=\;\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr)\;=\Phi_{\theta}(f_{X}(A,U);\ X,A\bigr)\;=\;\psi_{X}\bigl(U\bigr)$ does not depend on $A$.
+Therefore, fixing $U$ as the same and plugging in $\mathrm{do}\bigl(A=\gamma\bigr)$, we have:
+$\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr)=\Phi_{\theta}\bigr(f_{X}(A,U);\ X,A\bigr)=\psi_{X}\bigl(U\bigr)=\Phi_{\theta}\bigr(f_{X}(\gamma,U);\ X,\gamma\bigr)=\Phi_{\theta}(Y_{\mathrm{CF}}^{\gamma};\ X,\gamma),$ (34)
+or more simply,
+$\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr)=\Phi_{\theta}(Y_{\mathrm{CF}}^{\gamma};\ X,\gamma),$ (35)
+we further have,
+$\Phi_{\theta}^{-1}(\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr);\ X,\gamma)=\Phi_{\theta}^{-1}(\Phi_{\theta}(Y_{\mathrm{CF}}^{\gamma};\ X,\gamma);\ X,\gamma).$ (36)
+Noting that the L.H.S. of (36) is exactly the counterfactual procedures (encoding-decoding) of PO-Flow, and by (33), the R.H.S. of (36) equals to $Y_{\mathrm{CF}}^{\gamma}$. Therefore,
+$\hat{Y}_{\mathrm{CF}}^{\gamma}:=\Phi_{\theta}^{-1}\bigl(Z;\ X,\gamma\bigr)=Y_{\mathrm{CF}}^{\gamma}.$ (37)
+Proposition 4.7. Given a noise sample $z\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
+$\log p_{\theta,Y\mid X,A}\bigl(y(0)\mid x,a,z\bigr)=\log q(z)\;+\;\int_{0}^{1}\nabla_{y}\!\cdot\,v_{\theta}\bigl(y(t),t;\,x,a\bigr)\,dt.$
 Proof:
 Samples from Continuous Normalizing Flows (CNFs) evolve according to the following Neural ODE:
-$\frac{dy(t)}{dt}=v_{\theta}(y(t),t,x,a),\quad t\in[0,1],$ (21)
+$\frac{dy(t)}{dt}=v_{\theta}(y(t),t;x,a),\quad t\in[0,1],$ (38)
 which induces a corresponding evolution of the sample density governed by the Liouville continuity equation:
-$\partial_{t}p(y,t)+\nabla\cdot\big{(}p(y,t)v_{\theta}(y,t,x,a)\big{)}=0.$ (22)
-Here, $p(y,t)$ denotes the time-dependent probability density of the samples $y$.
-Next, we have that the dynamics of the density $p(\cdot)$ governed by the velocity field $v_{\theta}(y,t,x,a)$ is given by:
-$\frac{d}{dt}\log p(y(t),t)$ $=\frac{\nabla p(y(t),t)\cdot\partial_{t}y(t)+\partial_{t}p(y(t),t)}{p(y(t),t)}$ (23)
-$=\frac{\nabla p\cdot\partial_{t}y+\partial_{t}p}{p}\Big{|}_{(y(t),t)}$ (24)
+$\partial p_{\theta}(y,t)+\nabla\cdot\big(p_{\theta}(y,t)v_{\theta}(y,t;x,a)\big)=0.$ (39)
+Here, $p_{\theta}(y,t)$ denotes the time-dependent probability density of the samples $y$.
+Next, we have that the dynamics of the density $p_{\theta}(\cdot)$ governed by the velocity field $v_{\theta}(y,t,x,a)$ is given by:
+$\frac{d}{dt}\log p_{\theta}(y(t),t)$ $=\frac{\nabla p_{\theta}(y(t),t)\cdot\partial_{t}y(t)+\partial_{t}p_{\theta}(y(t),t)}{p_{\theta}(y(t),t)}$ (40)
+$=\frac{\nabla p_{\theta}\cdot\partial_{t}y+\partial_{t}p_{\theta}}{p_{\theta}}\Big|_{(y(t),t)}$ (41)
 […]
-$=\frac{\nabla p\cdot v_{\theta}-(\nabla p\cdot v_{\theta}+p\nabla\cdot v_{\theta})}{p}\Big{|}_{(y(t),t)}$ (26)
+$=\frac{\nabla p_{\theta}\cdot v_{\theta}-(\nabla p_{\theta}\cdot v_{\theta}+p_{\theta}\nabla\cdot v_{\theta})}{p_{\theta}}\Big|_{(y(t),t)}$ (43)
 $=-\nabla\cdot v_{\theta}.$ (44)
-Starting from an initial sample $y_{1}\sim q(\cdot)$ and integrating from $t=1$ to $t=0$, we have:
-$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=\log q(y_{1})+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)dt.$ (28)
+Starting from an initial sample $z\sim q(\cdot)$ and integrating from $t=1$ to $t=0$, we have:
+$\log p_{\theta,Y\mid X,A}\bigl(y(0)\mid x,a,z\bigr)=\log q(z)\;+\;\int_{0}^{1}\nabla_{y}\!\cdot\,v_{\theta}\bigl(y(t),t;\,x,a\bigr)\,dt.$ (45)
 Besides, if $q(\cdot)$ is selected as the commonly used $N(0,I)$, we have:
-$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=-\frac{d_{Y}}{2}\log 2\pi-\frac{1}{2}\|y_{1}\|^{2}+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)dt.$ (29)
+$\log p_{\theta,Y\mid X,A}\bigl(y(0)\mid x,a,z\bigr)=-\frac{d_{Y}}{2}\log 2\pi-\frac{1}{2}\|z\|^{2}+\int_{0}^{1}\nabla_{y}\cdot v_{\theta}(y(t),t;x,a)dt.$ (46)
 Appendix B Experimental Details
@@
 Each dataset provides three synthetic values per individual: the factual outcome $y^{(a)}$, the assigned treatment $a$, and the mean potential outcomes under both treatments, $\mu_{0}$ (for $a=0$) and $\mu_{1}$ (for $a=1$).
-Moreover, under the following dataset synthesis assumptions: (1) fully specified outcome model $Y^{(A)}=f(X,A)+\epsilon$; (2) shared noise across treatment arms for each individual; and (3) invertibility of the structural function such that, given $X$, $A$, and $Y^{(A)}$, one can solve for $\epsilon$ — we can synthesize the counterfactual outcome through the following procedure:
+Moreover, under the following dataset synthesis assumptions: (1) fully specified outcome model $Y^{(A)}=f(X,A)+\epsilon$; (2) shared noise across treatment arms for each individual; and (3) invertibility of the structural function such that, given $X$, AA, and $Y^{(A)}$, one can solve for $\epsilon$ — we can synthesize the counterfactual outcome through the following procedure:
 1. Abduction: infer the unobserved noise term $\epsilon^{\prime}$ from the factual outcome $Y^{(A)}$, covariates $X$, and treatment $A$;
-2. Action: intervene by changing the treatment assignment to $1-A$;
+2. Action: intervene by changing the treatment assignment to 1A1-A;
 3. Prediction: use the inferred noise $\epsilon^{\prime}$ and new treatment to compute the counterfactual outcome via $Y^{(1-A)}=f(X,1-A)+\epsilon^{\prime}$.
@@
 where $\beta$, $W$, and $\omega$ are parameters defined by the specific simulation setting.
-B.2 Counterfactual Image Experiments
-To intuitively visualize PO-Flow’s counterfactual performance and evaluate its effectiveness on high-dimensional data, we train PO-Flow on the latent space of a pre-trained Variational Autoencoder (VAE) [36] using the CelebA dataset.
-The CelebFaces Attributes Dataset (CelebA) comprises 202,599 face images, each with a resolution of $178\times 218$ pixels.
-Each image is annotated with 40 binary facial attributes.
-In our experiments, we selected the male/female attribute as the treatment variable $A\in\{\text{male},\text{female}\}$, and defined covariates $X\in\mathbb{R}^{6}$ to represent features such as hair, eyebrows, bangs, and lips.
-The VAE is pre-trained on the full CelebA dataset with a 512-dimensional latent space.
-Each $178\times 218$ face image is then encoded into this latent space, yielding the training target $Y\in\mathbb{R}^{512}$.
-PO-Flow is subsequently trained on the triplet $(Y,X,A)$ to generate counterfactual male face representations given original female faces.
-B.3 Neural Network Structure and Implementation Information
-The neural network used in experiments on causal datasets (except the CelebA image task) takes as input $y(t)$, $x$, $t$, and $a$.
+B.2 Neural Network Structure and Implementation Information
+The neural network used in experiments on causal datasets takes as input $y(t)$, $x$, $t$, and $a$.
 The model is composed of the following components: (1) an outcome embedding layer that projects $y(t)$ to the same dimension as the concatenated $(x,a)$; (2) a FiLM layer that applies feature-wise affine transformations to the embedded $y(t)$, conditioned on $(x,a)$; (3) two residual blocks, each consisting of a gated two-layer MLP conditioned on $(x,a)$, with residual averaging and skip connections; and (4) a projection layer that maps the final hidden representation to two velocity vectors $v_{0}$ and $v_{1}$, corresponding to treatments 0 and 1.
-The neural network used for the counterfactual image generation task has an architecture composed of the following components: (1) a conditioning MLP that encodes the pair $(t,a)$ into a hidden feature vector via two linear layers and a ReLU activation; (2) an input layer that takes the concatenation of $y(t)$ and $(t,a)$, followed by a ReLU-activated linear transformation; (3) four residual blocks, each composed of two linear layers with a ReLU nonlinearity, where the conditioning feature is injected additively at each block; and (4) an output linear layer that maps the final hidden representation back to the latent outcome dimension.
-All experiments are conducted on a V100 GPU with 16 GB of memory.
-The training time for causal datasets (excluding CelebA counterfactual face image generation) depends on the dataset and batch sizes.
+All experiments are conducted on an A100 GPU with 16 GB of memory.
+The training time for causal datasets depends on the dataset and batch sizes.
 For example, training on the ACIC 2018 dataset with a batch size of 256 typically takes 2–4 seconds per epoch.
 Log-density evaluation, as discussed in Section 4.3, involves divergence integration and requires significantly more time than other test metrics.
-B.4 Runge-Kutta numerical integration method
+B.3 Runge-Kutta numerical integration method
 We apply the classical 4th-order Runge-Kutta method to solve the Neural ODE in the forward process of PO-Flow.
-Given a factual outcome $y^{(a)}$, we aim to compute the latent embedding $z_{y}=y(1)$ by integrating the velocity field $v_{\theta}$ from $t=0$ to $t=1$, with initial condition $y(0)=y^{(a)}$.
+Given a factual outcome $y^{(a)}$, we aim to compute the latent embedding $z=y(1)$ by integrating the velocity field $v_{\theta}$ from $t=0$ to $t=1$, with initial condition $y(0)=y^{(a)}$.
 Let $h$ denote the step size, and define discrete time steps $t_{k}=kh$, for $k=0,1,\dots,T$, such that $Th=1$.
@@
 $k_{3}$ $=v_{\theta}\left(y(t_{k})+\frac{h}{2}k_{2},t_{k}+\frac{h}{2},x,a\right),\ k_{4}=v_{\theta}\left(y(t_{k})+hk_{3},t_{k}+h,x,a\right),$
-with $y(0)=y^{(a)}$ and $y(1)=z_{y}$. This produces a numerically integrated trajectory from the observed sample to its latent representation.
-B.5 Hutchinson trace estimator
+with $y(0)=y^{(a)}$ and $y(1)=z$. This produces a numerically integrated trajectory from the observed sample to its latent representation.
+B.4 Hutchinson trace estimator
 The computations of the log-density of the POs involve the calculation of $\nabla\cdot v_{k}(x,t)$, i.e., the divergence of the velocity field represented by a neural network.
@@
 $\small\nabla\cdot v_{\theta}(y(t),t,x,a)\approx\mathbb{E}_{p(\epsilon)}\left[\epsilon\cdot\frac{v_{\theta}(y(t)+\sigma\epsilon,t,x,a)-v_{\theta}(y(t),t,x,a)}{\sigma}\right],$ (48)
-where $p(\epsilon)$ is a distribution in $\mathbb{R}^{d_{Y}}$ satisfying $\mathbb{E}[\epsilon]=0$ and Cov$(\epsilon)=I$ (e.g., a standard Gaussian). This approximation becomes exact as $\sigma\to 0$.
-B.6 Adapted Implementation Details for DiffPO and GANITE Counterfactuals
+[…]
+B.5 Adapted Implementation Details for DiffPO and GANITE Counterfactuals
 DiffPO was not originally designed for counterfactual tasks.
@@
  3:    for $t=0,\cdots,T-1$:
-  4:       $z_{t+1}\leftarrow\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}z_{t}+f_{\theta}(z_{t},t|x,a)\left(\sqrt{1-\alpha_{t+1}}-\sqrt{\frac{\alpha_{t+1}(1-\alpha_{t})}{\alpha_{t}}}\right)$
-  5:     $z_{y}\leftarrow z_{T}$
+ 4:     $z_{t+1}\leftarrow\sqrt{\frac{\alpha_{t+1}}{\alpha}}z+f_{\theta}(z,t|x,a)\left(\sqrt{1-\alpha_{t+1}}-\sqrt{\frac{\alpha_{t+1}(1-\alpha)}{\alpha}}\right)$
+ 5:    $z\leftarrow z_{T}$
  6: Reverse Process:
@@
 However, the generator also produces a proxy counterfactual given the factual input, which we use to evaluate its counterfactual estimation performance.
-Appendix C Additional Experimental Results
-C.1 RMSE Convergence
+Figure 2: Illustration of counterfactual predictions. Left: two potential outcomes distribution. Right: Base (noise) distribution.
+As shown in Figure 2, standard diffusion models use noisy and stochastic forward processes that may degrade the information in $z$, producing reverse samples that resemble interventional potential outcomes rather than counterfactuals.
+In practice, probability flow ODEs [53, 11] or DDIM [52] can be applied to obtain a velocity field equivalent to the ODE derived from the trained diffusion score/noise, under which diffusion and flow matching are theoretically equivalent.
+Appendix C Comparisons to BGM settings
+Bijective Generation Mechanisms (BGM).
+The BGM framework [43] shows that if the true structural mechanism $f$ is in the BGM class (bijective/strictly monotone in the exogenous noise), then any learned mechanism $\hat{f}$ that (i) is also in the BGM class and (ii) matches the observed distribution—i.e., for the same parents value $X$ and action $A$,
+$\hat{f}_{X}(A,U)\ \stackrel{{\scriptstyle d}}{{=}}\ f_{X}(A,U)\quad(\text{equivalently }(\hat{f}(X,\cdot))_{\#}P_{U}=(f(X,\cdot))_{\#}P_{U}),$
+yields the same counterfactuals as $f$. Their result is model-agnostic, at the class level.
+PO-Flow.
+Our findings are specific to a particular methodology and architecture, as this is a methodology paper where we analyze continuous normalizing flows (CNFs).
+With the additional assumption (A3) specific to PO-Flow, Proposition 4.3 shows that the PO-Flow encoded latent $Z$ is bijective in the exogenous noise $U$.
+Consequently, Corollary 4.5 proves that the encode–decode procedure recovers the true counterfactual.
+Relationships.
+Because of Proposition 4.3 (enabled by the CNF), PO-Flow implements a bijective-in-noise mechanism and is thus comparable to BGM in the sense that $\hat{f}$ is bijective in $U$.
+However, BGM additionally requires observational distribution matching to obtain class-level identifiability, whereas the proofs of Proposition 4.3 and Corollary 4.5 do not assume such matching.
+Alternative route. The counterfactual recovery result (Corollary 4.5) can alternatively be established by imposing an additional assumption:
+(A4) Observational matching. For each $(x,a)$, the PO-Flow–induced observational law matches the true one, i.e.,
+$(\hat{f}_{X}(a,\cdot))_{\#}P_{U}\;=\;(f_{X}(a,\cdot))_{\#}P_{U}.$
+Under (A1)–(A3), Proposition 4.3 establishes a continuously differentiable bijection $\psi_{X}:U\to Z$ (independent of $A$), so the induced mechanism $\hat{f}_{X}$ implemented by PO-Flow is bijective in the exogenous noise, i.e., PO-Flow lies in the BGM class.
+With the additional assumption (A4), we can therefore directly invoke the counterfactual recovery result of the BGM framework [43].
+Appendix D Additional Experimental Results
+D.1 Empirical Validation of Assumption (A3)
+We test (A3): $p_{\theta}(Z\mid X,A)=q(Z)=\mathcal{N}(0,1)$, using a two-sample Maximum Mean Discrepancy (MMD).
+Specifically, we test the joint factorization $p(Z,X,A)=q(Z)p(X,A)$ by computing the MMD between the empirical joint sample $\{(z_{i},x_{i},a_{i})\}$ and a synthetic joint sample $\{(z_{i}^{\prime},x_{i},a_{i})\}$, where $z_{i}^{\prime}\sim q=\mathcal{N}(0,1)$.
+We employ a product kernel $k((z,x,a),(z^{\prime},x^{\prime},a^{\prime}))=k_{Z}(z,z^{\prime})\cdot k_{X}(x,x^{\prime})\cdot k_{A}(a,a^{\prime})$, where $k_{Z}(z,z^{\prime})$ and $k_{X}(x,x^{\prime})$ are radial basis function (RBF) mixtures and $k_{A}(a,a^{\prime})=\mathbf{1}[a=a^{\prime}]$ is the indicator kernel for discrete actions:
+$k_{Z}(z,z^{\prime})$ $=\exp\!\left(-\frac{\|z-z^{\prime}\|^{2}}{2\sigma_{Z}^{2}}\right),$ (49)
+$k_{X}(x,x^{\prime})$ $=\exp\!\left(-\frac{\|x-x^{\prime}\|^{2}}{2\sigma_{X}^{2}}\right).$
+To set the bandwidths $\{\sigma_{Z}\}$ and $\{\sigma_{X}\}$, we first compute the pooled pairwise distances on the union of the observed and synthetic sets (e.g., $\tilde{Z}=Z\cup Z^{\prime}$).
+We then calculate the median of these distances, and define the bandwidths as $\tfrac{1}{2}\,\text{median}$.
+The empirical MMD is given as:
+$\widehat{\mathrm{MMD}}^{2}_{u}=\frac{1}{n(n-1)}\sum_{i\neq j}k(u_{i},u_{j})+\frac{1}{n(n-1)}\sum_{i\neq j}k(v_{i},v_{j})-\frac{2}{n(n-1)}\sum_{i\neq j}k(u_{i},v_{j}),$ (50)
+with $u_{i}=(z_{i},x_{i},a_{i})$, $v_{i}=(z_{i}^{\prime},x_{i},a_{i})$.
+For more reliable comparisons, we additionally sample two ground-truth sets, $\{(z_{i}^{\prime},x_{i},a_{i})\}$ and $\{(z_{i}^{\prime\prime},x_{i},a_{i})\}$, where both $z_{i}^{\prime}\sim\mathcal{N}(0,1)$ and $z_{i}^{\prime\prime}\sim\mathcal{N}(0,1)$.
+The test is conducted on each dataset considered in the paper, and the aggregated results are summarized in the table below.
+Table 7: Empirical validation of Assumption (A3), $p_{\theta}(Z\mid X,A)=q(Z)=\mathcal{N}(0,1)$.
+We report two-sample MMD values comparing the empirical joint distribution $\{(z_{i},x_{i},a_{i})\}$ with synthetic samples $\{(z_{i}^{\prime},x_{i},a_{i})\}$ across datasets.
+Lower MMD indicates closer agreement with the assumed factorization $p(Z,X,A)=q(Z)p(X,A)$.
+ACIC 2016 ACIC 2018 IHDP IBM
+PO-Flow $4.9\times 10^{-2}$ $7.2\times 10^{-3}$ $2.7\times 10^{-2}$ $6.8\times 10^{-2}$
+True $4.1\times 10^{-2}$ $5.9\times 10^{-3}$ $2.5\times 10^{-2}$ $5.3\times 10^{-2}$
+As shown in Table 7, PO-Flow achieves small two-sample joint MMD values, which are comparable to those obtained from the ground-truth independent samples $\{(z_{i}^{\prime},x_{i},a_{i})\}$ and $\{(z_{i}^{\prime\prime},x_{i},a_{i})\}$.
+This provides empirical support for the validity of Assumption (A3).
+D.2 Identifying the Most Likely Potential Outcomes via Log-Density
+As shown in Proposition 4.7 and Section 5.3.1, we can naturally select the most likely potential outcome based on log-density.
+Table 8 reports the RMSE of PO-Flow when using the most likely PO selected by the algorithm.
+Specifically, for each sample we simulate 100 POs with PO-Flow, select the one with the highest log-density, and compute the results accordingly.
+Table 8: Root mean squared error (RMSE) of estimated potential outcomes (POs) across different methods on the ACIC 2018, IHDP, and IBM datasets.
+For PO-Flow, results are based on the most likely PO selected from 100 simulated outcomes using log-density.
+All values are averaged over 10-fold cross-validation.
+ACIC 2018 IHDP IBM
+$\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$
+PO-Flow $\mathbf{0.42_{\pm.07}}$ $\mathbf{0.54_{\pm.10}}$ $\mathbf{0.96_{\pm.09}}$ $\mathbf{1.05_{\pm.10}}$ $\mathbf{0.84_{\pm.06}}$ $\mathbf{0.90_{\pm.08}}$
+DiffPO $0.72_{\pm.14}$ $0.79_{\pm.16}$ $1.33_{\pm.20}$ $1.40_{\pm.22}$ $1.82_{\pm.38}$ $1.85_{\pm.35}$
+INFs $0.69_{\pm.16}$ $0.78_{\pm.17}$ $1.02_{\pm.10}$ $1.20_{\pm.12}$ $1.52_{\pm.24}$ $1.57_{\pm.28}$
+S-learner $0.54_{\pm.09}$ $0.57_{\pm.11}$ $1.31_{\pm.18}$ $1.44_{\pm.20}$ $01_{\pm.17} 1.1$$16_{\pm.19}$
+T-learner $1.57_{\pm.32}$ $1.71_{\pm.40}$ $1.45_{\pm.25}$ $1.49_{\pm.27}$ $1.89_{\pm.47}$ $1.96_{\pm.50}$
+CEVAE $0.83_{\pm.17}$ $0.85_{\pm.17}$ $1.18_{\pm.15}$ $1.36_{\pm.18}$ $0.96_{\pm.12}$ $0.98_{\pm.12}$
+CFR $0.94_{\pm.10}$ $0.98_{\pm.12}$ $10_{\pm{.20}} 1.1$$18_{\pm{.20}}$ $2.09_{\pm.45}$ $2.17_{\pm.48}$
+GANITE $0.88_{\pm.15}$ $0.97_{\pm.15}$ $1.60_{\pm.36}$ $1.67_{\pm.34}$ $2.48_{\pm.39}$ $2.59_{\pm.47}$
+Compared with the averaged RMSE over all POs generated by PO-Flow in Table 1, selecting the most likely POs based on log-density yields smaller RMSEs, providing more confident and closer estimates.
+D.3 Convergence and Computational Costs
+(a) Convergence of MSE and RMSE for predicted potential outcomes over the training iterations on the ACIC 2018, IHDP, and IBM datasets.
+Figure 5(a) shows the convergence of RMSE on multiple datasets.
 PO-Flow demonstrates rapid convergence across datasets, requiring only around 250 training iterations (not epochs) with a batch size of 200 per iteration.
 This highlights the model’s strong training efficiency.
-C.2 Additional counterfactual images
-Figure 6: PO-Flow counterfactual image generation from female to male faces.
-$(a)$: Each row shows a factual outcome $y^{(a)}$ (1st column), its encoded latent $z_{y}$ (2nd column), and reverse-time decoding at $t=2/3$, $t=1/3$, and $t=0$ (3rd–5th columns), yielding the counterfactual $\hat{y}^{(1-a)}$.
-$(b)$: Counterfactuals decoded from perturbed latents $z_{y}+\delta$, with $\delta\sim\mathcal{N}(0,0.2I)$.
+Table 9: Comparison of model size, training time untill convergence, and sampling cost (for 500 samples).
+Methods # Params Training Time Sampling
+PO-Flow $32,394$ $\mathbf{0.96}$min $\mathbf{0.58}$s
+DiffPO $176,372$ 5.755.75min $4.98$s
+INFs $138,172$ 5.095.09min $0.74$s
+GANITE $167,381$ 4.934.93min $0.80$s
+CEVAE $97,864$ 4.644.64min $0.47$s
+In addition, Table 9 compares model size, training time to convergence, and sampling time (for 500 samples) across PO-Flow and other deep generative baselines.
+All experiments were conducted on a single A100 GPU.
+Notably, PO-Flow attains superior results with the fewest neural network parameters, leading to faster training and substantially reduced sampling time compared with diffusion-, discrete flow-, and GAN-based methods.

### v3

@@
-PO-Flow: Flow-based Generative Models for Sampling Potential Outcomes and Counterfactuals
+Flow-based Generative Modeling of Potential Outcomes and Counterfactuals
 Abstract
-Predicting potential and counterfactual outcomes from observational data is central to clinical decision-making, where physicians must weigh treatments for an individual patient rather than relying solely on average effects at the population level.
-We propose PO-Flow, a continuous normalizing flow (CNF) framework for causal inference that jointly models potential outcomes and counterfactuals.
-Trained via flow matching, PO-Flow provides a unified approach to average treatment effect estimation, individualized potential outcome prediction, and counterfactual prediction.
-Besides, PO-Flow directly learns the densities of potential outcomes, enabling likelihood-based evaluation of predictions.
-Furthermore, PO-Flow explores counterfactual outcome generation conditioned on the observed factual in general observational datasets, with a supporting recovery result under certain assumptions.
-PO-Flow outperforms modern baselines across diverse datasets and causal tasks in the potential outcomes framework.
-1 Introduction
-Predicting potential outcomes (POs) lies at the core of causal inference.
-In a typical setting, each individual has two potential outcomes: $Y^{(1)}$ if they receive a treatment ($A=1$) and $Y^{(0)}$ if they do not ($A=0$).
-However, in the real world, only one outcome is observed, while the other potential outcome remains fundamentally unknown [25].
-Accurately estimating both POs is crucial for decision-making in healthcare [37, 24, 46], where clinicians may wish to compare a patient’s likely outcomes under different treatment options rather than rely solely on average effects [56, 17].
-In the Potential Outcomes (POs) literature, the vast majority of methods focus on estimating the conditional average treatment effect (CATE): $\mathbb{E}[Y^{(1)}-Y^{(0)}|X]$, at the population level [28, 22, 1, 19, 33, 38, 50, 61].
-It is important to note that although some methods perform well in estimating population-level CATE, they may underperform in individualized POs prediction, as accurately predicting POs in finite samples is challenging [47].
-However, making decisions for individuals based on average treatment effects can be misleading.
-Therefore, accurately predicting individualized POs and learning their full distributions $p(Y^{(A)}|X,A)$ is a crucial task [63, 31].
-Recent works [61, 40, 41] have explored generative models to sample POs and estimate empirical distributions.
-However, methods that learn the full density of POs without relying on explicit distributional assumptions (e.g., Gaussian mixtures) remain scarce.
-In addition, another critical challenge is estimating the counterfactual outcome conditioned on the factual: $Y_{\mathrm{CF}}^{(1-A)}:=Y^{(1-A)}|X,Y^{(A)}$, i.e., predicting what would have happened under the alternative treatment, given the observed outcome $Y^{(A)}$.
-This task is especially relevant in post-hoc clinical analyses, where practitioners ask, “Would the patient have fared better on another therapy?” [60, 18].
-Counterfactual prediction differs from standard PO estimation—it requires reasoning about outcomes under an alternative treatment conditioned on the observed factual outcome, introducing dependencies between observed and unobserved outcomes.
-Structural Causal Models (SCMs) [44] and modern extensions [42, 6, 12, 7, 59, 45, 8] offer the most widely used theoretical framework for counterfactual reasoning.
-SCMs methods typically assume known causal structures or aim to learn them explicitly.
-However, it may be possible to estimate counterfactuals directly without recovering the SCM [64].
-In PO-Flow, we do not explicitly assume an SCM, but seek to directly estimate counterfactual outcomes.
-Despite recent progress, two key gaps remain in the potential outcomes literature.
-First, counterfactual outcome estimation conditioned on the observed factual outcome has received little attention, though it is crucial for post-hoc analyses.
-Second, beyond counterfactuals, most methods are specialized—some target CATE estimation, others aim at potential outcome prediction or distributional modeling—but few provide a unified and scalable algorithmic framework that can jointly address these tasks.
-1.1 Main Contributions
-We propose PO-Flow, a continuous normalizing flow (CNF) framework for potential outcomes and counterfactuals.
-Our main contributions are: (1) Enable counterfactual outcome generation conditioned on the observed factual, together with a supporting counterfactual recovery result (Corollary 4.5) under certain assumptions; (2) Provide a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction; (3) Learn the density of potential outcomes, supporting principled uncertainty quantification; (4) Achieve superior performance over state-of-the-art methods across diverse causal tasks and benchmark datasets.
-1.2 Related Works
-CATE Estimation: Numerous methods have been proposed for CATE estimation [28, 22, 1, 19, 33, 38, 50, 61].
-Broadly, these fall into model-agnostic and model-specific approaches.
-Model-agnostic estimators, or meta-learners [33, 15, 14, 2, 49], typically follow two strategies [39]: (a) one-step methods, which fit regression models and compute CATE as the difference in predicted outcomes, and (b) two-step methods, which estimate nuisance functions to construct pseudo-outcomes, then regress them on covariates.
-In contrast, model-specific estimators adapt machine learning techniques for CATE estimation, including methods such as [13, 50, 38, 30, 61].
-More recently, neural networks have been widely applied to this task, with notable examples in [50, 38, 61].
-However, most of these approaches mainly focus on estimating CATE rather than predicting individualized potential outcomes.
-Individualized POs Prediction: Although some CATE methods can be adapted for potential outcomes (POs) prediction, they may perform suboptimally, particularly on finite samples, as they are not specifically optimized for accurate PO estimation.
-This limitation partly arises because many methods assume that CATE has a simpler functional structure than POs.
-Nevertheless, several CATE approaches are also designed with POs prediction as the target [33, 50, 61, 38, 41, 39].
-Counterfactual Outcomes Prediction: Structural Causal Models (SCMs) provide a well-established framework for counterfactual reasoning [44, 42, 6, 12, 7, 59, 45, 8], but practical applications often require specifying the full causal graph and structural equations, which may be difficult or infeasible for real-world data.
-Some recent works avoid explicitly modeling SCMs [35, 5, 40, 57], but typically address specific tasks and rely on additional information, such as temporal structure.
-Some works [24, 30, 38, 34, 61] targeting CATE estimation describe their goal as counterfactual inference.
-However, in these works, "counterfactual" typically refers to the unobserved potential outcome under an alternative treatment, without conditioning on the observed factual outcome.
-While SCMs offer theory-grounded counterfactual reasoning, general-purpose methods for individualized counterfactual prediction—without explicitly modeling SCMs or relying on auxiliary information—remain scarce.
-Although [61, 39] were not originally designed for counterfactuals, we can adapt them by incorporating factual information into intermediate steps and use them as baselines.
-Diffusion and Flow Models: Diffusion models [23, 52, 54] and flow-based generative models [10, 58, 36, 4] have demonstrated strong performance across a range of applications.
-Recently, diffusion and flow-based generative models have been applied to causal inference tasks such as counterfactual image generation [48, 16], causal query answering [9], and inference over causal graphs or structural causal models [8, 29].
-To our knowledge, two generative algorithms have been proposed specifically for the potential outcomes (PO) task: the diffusion-based DiffPO [39] and the discrete normalizing flow-based INFs [41].
-We conduct extensive comparisons with both methods in this work.
-2 Preliminaries
-2.1 Neural ODE and Continuous Normalizing Flow:
-A Neural ODE is a continuous model where the trajectory of data is modeled as the solution of an ordinary differential equation (ODE).
-Formally, in $\mathbb{R}^{d}$, given an input $y_{0}=y(0)$ at time $t=0$, the transformation to the output $y_{1}=y(1)$ at time $t=1$ is governed by:
-$\small\frac{dy(t)}{dt}=v_{\theta}(y(t),t),\quad t\in[0,1],$ (1)
-where $v_{\theta}:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ represents the velocity field, which is parameterized by a neural network with parameter $\theta$.
-The time horizon is rescaled to $[0,1]$ without loss of generality.
-Continuous Normalizing Flow (CNF) is a class of normalizing flows in which the transformation of a probability density is governed by a time-continuous Neural ODE.
-The marginal density of $y(t)$, denoted as $p(y,t)$, evolves according to the continuity equation implied by the ODE in Eq.
-(1).
-This continuity equation is given by:
-$\small\partial_{t}p(y,t)+\nabla\cdot(p(y,t)v_{\theta}(y,t))=0,\quad t\in[0,1],$ (2)
-where $\nabla\cdot$ denotes the divergence operator.
-When the Neural ODE is well-posed, it defines a continuous and invertible mapping from the initial sample $y_{0}$ to the terminal sample $y_{1}$.
-The inverse mapping from $y_{1}$ to $y_{0}$ can be computed by integrating the ODE (1) in reverse time.
-This framework allows setting $y_{0}\sim p(\cdot,0)$ as the data distribution and $y_{1}\sim p(\cdot,1)$ as the base (noise) distribution, typically chosen as $N(0,I)$.
-For notational convenience, we define $p_{Y}(\cdot):=p(\cdot,0)$ as the data distribution and $q(\cdot):=p(\cdot,1)$ as the base (noise) distribution.
-Throughout this paper, we adopt the common choice $q(\cdot)=N(0,I)$.
-2.2 Flow Matching
-Flow Matching (FM) [36, 4] provides a simulation-free approach for training continuous normalizing flows by directly regressing the model velocity $v_{\theta}(y,t)$ onto a predefined target velocity field $u(y,t)$.
-The original FM objective is defined as:
-$\small\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],y\sim p(\cdot,t)}\left[\|v_{\theta}(y,t)-u(y,t)\|^{2}\right],$ (3)
-where $u(y,t)$ is an analytically specified target velocity field.
-Linear Interpolant: Due to the intractability of $u(y,t)$ without conditioning on the initial data $y_{0}$, Conditional Flow Matching (CFM) is introduced.
-Conditioned on an observed sample $y_{0}\sim p_{Y}(\cdot)$, CFM employs a pre-specified interpolation function $I_{t}$ between two endpoints $y_{0}$ and $y_{1}\sim q(\cdot)$, where $q$ is typically $N(0,I)$.
-The interpolated path is defined as:
-$\small\phi(y_{0},y_{1};t):=I_{t}(y_{0},y_{1}),\quad t\in[0,1].$ (4)
-Here, $I_{t}$ can be analytically specified. Conditioned on $y_{0}$, the optimal and most information-preserving choice is the linear interpolation path [36]:
-$\small\phi(y_{0},y_{1};t)=(1-t)y_{0}+(t+\sigma_{\text{min}}(1-t))y_{1},$ (5)
-where $y_{1}\sim N(0,I)$ and $\sigma_{\text{min}}$ is a small positive hyperparameter that ensures $p(\phi,0)\sim N(y_{0},\sigma_{\text{min}}^{2}I)$.
-When $\sigma_{\text{min}}=0$, this reduces to the strict linear interpolant $\phi=(1-t)y_{0}+ty_{1}$.
-Here, $\phi$ serves as a pre-specified reference trajectory that the ODE solution $y(t)$ should follow at time $t$.
-Under the linear interpolant defined in (5), the corresponding reference velocity field becomes:
-$\small\frac{d\phi}{dt}=(1-\sigma_{\text{min}})y_{1}-y_{0}.$ (6)
-Training loss: CFM trains the flow $v_{\theta}(y(t),t)$ by directly regressing it onto the reference velocity field (6). The training objective is given by:
-$\small\mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],y_{0}\sim p_{Y}(\cdot),y_{1}\sim q(\cdot)}\|v_{\theta}(\phi,t)-\frac{d\phi}{dt}\|^{2}.$ (7)
-3 Problem Settings
-Notations and Setups: We consider a setting where each individual is described by covariates $X\in\mathbb{R}^{d_{X}}$, a treatment assignment $A\in\{0,1\}$, and a potential outcome $Y^{(A)}\in\mathbb{R}$ under treatment $A$.
-The observational dataset is $\mathcal{D}=\{(y^{(a)}_{i},x_{i},a_{i})\}_{i=1}^{n}\sim p_{Y,X,A}(\cdot)$, where $y_{i}^{(a)}$ is the observed potential outcome under treatment $a_{i}$, and $(x_{i},a_{i})\sim p_{X,A}(\cdot)$ are the associated covariates and treatment.
-Potential Outcomes: In the potential outcomes framework, each individual has two potential outcomes: $Y^{(0)}$ (control) and $Y^{(1)}$ (treatment), but only one is observed based on assignment $A$.
-We define potential outcome prediction as estimating $Y^{(a)}\mid X,A=a$.
-For convenience, we denote $Y^{(a)}$ and $\hat{Y}^{(a)}$ as the true and predicted potential outcomes.
-Counterfactual Outcomes: In contrast to potential outcomes, we define the counterfactual outcome as $Y_{\text{cf}}^{(1-a)}:=Y^{(1-a)}\mid X,A=a,Y^{(a)}$, which is the outcome had the individual received the opposite treatment, conditioned on its observed potential (factual) outcome.
-For convenience, we denote $Y_{\text{cf}}^{(1-a)}$ and $\hat{Y}_{\text{cf}}^{(1-a)}$ as the true and predicted counterfactual outcomes, respectively.
-To ensure the identifiability of potential outcomes (in Section 4.1), we adopt the standard assumptions commonly used in the literature [39, 41, 61]:
-Assumption 3.1.
-(1) Consistency: If an individual receives treatment $A=a$, the observed outcome equals the corresponding potential outcome: $Y=Y^{(a)}$.
-(2) Unconfoundedness: All confounders are observed, i.e., $\{Y^{(0)},Y^{(1)}\}\perp\!\!\!\perp A\mid X$.
-(3) Overlap: Treatment assignment is probabilistic: $0<P(A=1\mid X=x)<1$, for all $x\in\mathcal{X}$.
-In Section 4.2, we will introduce additional assumptions on counterfactual outcomes estimation in PO-Flow and provide the corresponding theoretical analyses for counterfactual recovery.
-4 PO-Flow: A Continuous Normalizing Flow for Potential Outcomes
-We introduce PO-Flow, a continuous normalizing flow framework for modeling potential and counterfactual outcomes, with three main goals: 1.
-predicting individualized potential outcomes, 2.
-predicting counterfactual outcomes, and 3.
-learning the densities.
-Our PO-Flow framework is built on Continuous Normalizing Flows (CNFs) and is trained using the Flow Matching (FM) objective.
-Our learning target is the trajectory of the sample $y(t)$ over $t\in[0,1]$.
-To condition the dynamics on both covariates $x$ and treatment $a$, the Neural ODE is defined as:
-$\small\frac{dy(t)}{dt}=v_{\theta}(y(t),t;x,a),\quad t\in[0,1],$ (8)
-where $v_{\theta}$ is parametrized by a neural network, with $y(0)$ following the data distribution and $y(1)\sim q(\cdot)$.
-Notation: Given a data point $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$, we define $y_{0}=y^{(a)}$ as the observed training sample at time $t=0$ in the Neural ODE, and $y_{1}\sim q(\cdot)$ as the noise sample at time $t=1$.
-Recall that we denote $\hat{y}^{(a)}$ and $\hat{y}_{\text{cf}}^{(1-a)}$ as the predicted potential and counterfactual outcomes, respectively.
-Training loss: PO-Flow adopts the Conditional Flow Matching (CFM) objective for training. Under the updated definitions, our training loss retains the form of Equation (7) and is given by:
-$\min_{\theta}\ \mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot),q(\cdot)}\|v_{\theta}(\phi,t;x,a)-\frac{d\phi}{dt}\|^{2},$ (9)
-where $\phi$ and $\frac{d\phi}{dt}$ are defined in (5) and (6), respectively, with $y_{0}=y^{(a)}$ and $y_{1}\sim q(\cdot)$.
-Remark: Following prior works [20, 39], one may estimate propensity scores $w(x)=p(A=1\mid X=x)$ with a pre-trained classifier and apply inverse weighting $\frac{1}{w(x)}$ in the loss.
-We found this provided little performance improvement, but both implementations are included in our code.
-Forward Process: We treat the forward process as an encoding operation, denoted by the function $Z:=\Phi_{\theta}(Y^{(A)};\,X,A)$.
-Given the velocity field $v_{\theta}$, the forward process refers to pushing a factual outcome $y^{(a)}$ (at $t=0$) to a latent embedding $z$ (at $t=1$) via:
-$z=\Phi_{\theta}(y^{(a)};\,x,a)=y^{(a)}+\int_{0}^{1}v_{\theta}(y(t),t;x,a)dt,\ \text{with}\ y(0)=y^{(a)}.$ (10)
-Through the forward process, $z$ encodes information from the factual outcome $y^{(a)}$ as well as its associated covariates $x$ and treatment $a$.
-Reverse Process: We treat the reverse process as a decoding operation.
-It is initialized differently for potential outcome (Section 4.1) and counterfactual prediction (Section 4.2).
-In general, at $t=1$, given a latent representation $z$, which can be either sampled from $q(\cdot)=N(0,1)$ or obtained by encoding a factual sample, the reverse process is defined as:
-$\hat{y}_{0}=\Phi_{\theta}^{-1}(z;x,a)=z-\int_{0}^{1}v_{\theta}(y(t),t;x,a)dt,\ \text{with}\ y(1)=z.$ (11)
-Empirically, we use the Runge-Kutta method (Appendix B.3) for numerical integration.
-During training, we construct synthetic paths $\phi$ between $y_{0}$ and $y_{1}\sim q(\cdot)$ using the CFM loss conditioned on $y_{0}$.
-As shown in Proposition A.1 (Appendix), the gradient of the CFM loss is equivalent to that of the original FM loss.
-Consequently, the learned velocity field $v_{\theta}$ generates $\hat{y}_{0}$ independent of the specific training samples $y_{0}$, and since $y_{1}\sim q(\cdot)$ is random, $v_{\theta}$ induces a distribution over outcomes conditioned on $(x,a)$.
-4.1 PO-Flow for Potential Outcomes Prediction
+Predicting potential and counterfactual outcomes from observational data is central to individualized decision-making, particularly in clinical settings where treatment choices must be tailored to each patient rather than guided solely by population averages.
+We propose PO-Flow, a continuous normalizing flow (CNF) framework for causal inference that jointly models potential outcome distributions and factual-conditioned counterfactual outcomes.
+Trained via flow matching, PO-Flow provides a unified approach to individualized potential outcome prediction, conditional average treatment effect estimation, and counterfactual prediction.
+By encoding an observed factual outcome into a shared latent representation and decoding it under an alternative treatment, PO-Flow relates factual and counterfactual realizations at the individual level, rather than generating counterfactuals independently from marginal conditional distributions.
+In addition, PO-Flow supports likelihood-based evaluation of potential outcomes, enabling uncertainty-aware assessment of predictions.
+A supporting recovery guarantee is established under certain assumptions, and empirical results on benchmark datasets demonstrate strong performance across a range of causal inference tasks within the potential outcomes framework.
+I Introduction
+Predicting potential outcomes (POs) is central to causal inference and individualized decision-making.
+In the potential outcomes framework, each individual with covariates $X$ has two outcomes: $Y^{(1)}$ if treated ($A=1$) and $Y^{(0)}$ if untreated ($A=0)$, but only one factual outcome $Y^{(A)}$ is observed [12].
+In applications such as healthcare, practitioners often require patient-level comparisons across treatments, rather than relying solely on population-level averages [9].
+Most existing methods focus on estimating the conditional average treatment effect (CATE), $\mathbb{E}[Y^{(1)}-Y^{(0)}\mid X]$, using meta-learners and model-specific approaches such as representation learning and deep models [18, 7, 8, 2, 29].
+However, strong CATE performance does not necessarily imply accurate individualized potential outcome prediction in finite samples, since the assumptions are sufficient for CATE estimation can be weaker than those required for individual-level modeling [28].
+Accurately predicting individualized potential outcomes and learning their conditional distributions $p(Y^{(a)}\mid X)$ for each treatment $A=a$ is therefore a crucial task, particularly for uncertainty quantification and risk-sensitive decision-making [15].
+Recent work has explored generative models for sampling potential outcomes and estimating empirical distributions [38, 23, 24].
+Nevertheless, methods that learn the full density of potential outcomes within the potential outcomes framework, without relying on restrictive distributional assumptions (e.g., Gaussian mixtures), are still relatively limited in the literature.
+Beyond potential outcome prediction, an important but less-studied causal query is factual-conditioned counterfactual prediction,
+$Y_{\mathrm{cf}}^{(1-A)}:=Y^{(1-A)}\mid X,\,Y^{(A)},$
+which asks what would have happened under the alternative treatment, given the observed factual outcome.
+This differs from standard PO prediction $p(Y^{(A)}\mid X)$, which does not condition on the realized outcome.
+Conditioning on $Y^{(A)}$ induces dependence between factual and counterfactual outcomes and is especially relevant for post-hoc analyses (e.g., “would this patient have done better on another therapy?”) [37, 10].
+Structural causal models (SCMs) provide a canonical framework for such reasoning, but typically require specifying a causal graph and parametric structural equations [26, 5, 27].
+In contrast, within the potential outcomes literature, general-purpose methods for individualized factual-conditioned counterfactual prediction remain underexplored.
+[^1] Notably, some works [16, 21, 19, 38] use the term “counterfactual” to refer to the unobserved potential outcome under an alternative treatment, without conditioning on the observed factual outcome.
+This motivates the need for methods that support counterfactual reasoning without requiring a parametric SCM, while remaining compatible with the potential outcomes framework.
+Despite recent progress, two key gaps remain.
+First, counterfactual outcome estimation conditioned on the observed factual outcome has received limited attention, despite its importance for individualized and post-hoc analyses.
+Second, existing methods are often specialized—targeting CATE estimation, potential outcome prediction, or distributional modeling in isolation—rather than providing a unified and scalable framework that jointly addresses these tasks.
+In this paper, we propose PO-Flow, a continuous normalizing flow (CNF) trained via flow matching [20], which provides a unified generative framework for
+  • (i)
+    individualized potential outcome prediction,
+  • (ii)
+    conditional average treatment effect (CATE) estimation,
+  • (iii)
+    factual-conditioned counterfactual prediction via an abduction–action–prediction procedure enabled by invertible flows, with a supporting recovery result (Corollary V.6), and
+  • (iv)
+    potential outcome density learning for likelihood-based evaluation and uncertainty quantification.
+By encoding an observed factual outcome into a shared latent representation and decoding it under an alternative treatment, PO-Flow enables mapping an individual factual outcome to a corresponding counterfactual outcome at the realization level, rather than generating outcomes independently from marginal conditional distributions.
+II Preliminaries
+A Continuous Normalizing Flow (CNF) defines an invertible transformation between a data variable and a simple base variable via a time-dependent Neural Ordinary Differential Equation (ODE).
+Starting from the data $y(0)=y_{0}$ at time $t=0$, the trajectory $y(t)$ evolves according to
+$\frac{dy(t)}{dt}=v_{\theta}(y(t),t),\quad t\in[0,1],$ (1)
+where $v_{\theta}:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ denotes the velocity field parameterized by a neural network with parameters $\theta$.
+Under standard regularity conditions, the Neural ODE induces an invertible mapping from the data $y_{0}$ at $t=0$ to a base variable $y_{1}=z$ at $t=1$.
+The inverse mapping from $y_{1}$ to $y_{0}$ is obtained by integrating (1) backward in time.
+Throughout this paper, we adopt the standard choice $y_{1}=z\sim\mathcal{N}(0,I)$ at $t=1$, and denote the corresponding base density by $q(\cdot)=N(0,I)$.
+We train the CNF using conditional flow matching (CFM) [20, 4], which avoids explicit density simulation.
+Given a data sample $y_{0}$ and an independent base sample $y_{1}\sim q(\cdot)$, CFM defines a reference path $\phi_{t}=\phi(y_{0},y_{1};t)$ connecting $y_{0}$ and $y_{1}$.
+We adopt the linear interpolant
+$\phi_{t}=(1-t)\,y_{0}+t\,y_{1},$ (2)
+whose associated reference velocity is
+$\frac{d\phi_{t}}{dt}=y_{1}-y_{0}.$ (3)
+CFM trains the velocity field $v_{\theta}$ to match this reference velocity along the interpolation path by minimizing
+$\begin{split}&\mathcal{L}_{\mathrm{CFM}}(\theta)\\ =&~\mathbb{E}_{t\sim\mathcal{U}[0,1],\,y_{0}\sim p_{\text{data}},\,y_{1}\sim q(\cdot)}\left\|v_{\theta}(\phi_{t},t)-\frac{d\phi_{t}}{dt}\right\|^{2}.\end{split}$ (4)
+After training, samples are generated by drawing $y_{1}\sim q(\cdot)$ and integrating (1) backward in time to obtain $y_{0}$.
+III Problem Settings
+In the Potential Outcomes (POs) framework, each individual is characterized by covariates $X\in\mathbb{R}^{d_{X}}$, a binary treatment assignment $A\in\{0,1\}$, and an observed factual outcome $Y^{(A)}\in\mathbb{R}$.
+We observe an i.i.d.
+dataset
+$\mathcal{D}=\{(y_{i},x_{i},a_{i})\}_{i=1}^{n}\sim p_{Y,X,A},$
+where $y_{i}=y_{i}^{(a_{i})}$ denotes the realized potential outcome under the received treatment.
+Each individual has two potential outcomes, $Y^{(0)}$ and $Y^{(1)}$, but only one is observed.
+Potential outcome (PO) prediction targets the conditional distribution $p(Y\mid X,A=a)$, which equals $p(Y^{(a)}\mid X)$ under Assumption III.1 when $A=a$.
+We use $Y^{(a)}$ and $\hat{Y}^{(a)}$ to denote the true and predicted potential outcomes under treatment $a$, respectively.
+The counterfactual outcome asks what would have happened under the alternative treatment, given the observed factual outcome $Y^{(a)}$:
+$Y_{\mathrm{cf}}^{(1-a)}:=Y^{(1-a)}\mid X,\,Y^{(a)}.$
+We denote the true and predicted counterfactual outcomes by $Y_{\mathrm{cf}}^{(1-a)}$ and $\hat{Y}_{\mathrm{cf}}^{(1-a)}$, respectively.
+To ensure the identifiability of potential outcomes, we adopt the standard assumptions commonly used in the literature [22, 24, 38]:
+Assumption III.1.
+  1. 1.
+    Consistency: If an individual receives treatment $A=a$, then $Y=Y^{(a)}$.
+  2. 2.
+    Unconfoundedness: All confounders are observed, i.e., $\{Y^{(0)},Y^{(1)}\}\perp\!\!\!\perp A\mid X$.
+  3. 3.
+    Overlap: $0<P(A=1\mid X=x)<1$ for all $x\in\mathcal{X}$.
+Later, we will introduce additional assumptions for factual-conditioned counterfactual prediction in PO-Flow and establish the corresponding theoretical properties.
 Require: Trained $v_{\theta}$
@@
  1: $z\sim q(\cdot)$
- 2: $\hat{y}_{0}=z-\int_{0}^{1}v_{\theta}(y(t),t;x,a)\,dt$
+1 2: $\hat{y}_{0}=z-\int_{0}^{1}v_{\theta}(y(t),t;x,a)\,dt,$
+ with $y(1)=z$.
 Return: $\hat{y}^{(a)}:=\hat{y}_{0}$
-Potential Outcomes Prediction
-Algorithm 1
-Potential Outcomes Prediction
-To sample potential outcomes $\hat{y}^{(a)}$ for an individual $(x,a)$, the procedure follows a simple reverse process, as defined in Equation (11).
-Starting from a noise sample $z\sim q(\cdot)$, we integrate the velocity field $v_{\theta}$ backward in time, conditioned on covariates $x$ and treatment $a$, to obtain a point estimate of the potential outcome $\hat{y}^{(a)}$.
-The steps are summarized in Algorithm 1.
-In addition to predicting point estimates of potential outcomes, PO-Flow is capable of learning the densities of the potential outcomes, as detailed in Section 4.3.
-4.2 PO-Flow for Counterfactual Outcomes Prediction
-The counterfactual outcome prediction consists of both a forward and a reverse process.
-Given a factual $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$, the forward process (Lines 1–2 in Alg.
-2) encodes $(y^{(a)},x,a)$ into a latent representation $z$.
-Subsequently, the reverse process (Lines 3–5) decodes $z$ under the counterfactual treatment 1a1-a, yielding the counterfactual prediction $\hat{y}_{\text{cf}}^{(1-a)}$.
-The process of counterfactual prediction is also illustrated in Figure 1.
+Algorithm 1 Potential Outcomes Prediction
 Require: Trained $v_{\theta}$
@@
  3. Reverse process {action and prediction}:
- 4.  Set intervened action to 1a1-a;
+ 4.  Set intervened treatment to 1a1-a;
 2 5.  $\hat{y}_{0}=z-\int_{0}^{1}v_{\theta}(y(t),t;x,1-a)\,dt,$
@@
 Return: $\hat{y}_{\text{cf}}^{(1-a)}:=\hat{y}_{0}$
-Counterfactual Prediction
-Algorithm 2
-Counterfactual Prediction
-Figure 1: Illustration of counterfactual predictions. Left: two potential outcomes distribution. Right: base (noise) distribution.
-Because of the smooth and information-preserving nature of the continuous normalizing flow path, the encoded $z$ retains rich information from the factual outcome $y^{(a)}$, enabling accurate decoding of the counterfactual $\hat{y}_{\text{cf}}^{(1-a)}$ under counterfactual treatment 1a1-a.
-To our knowledge, this work is among the early attempts at counterfactual prediction in the Potential Outcomes (POs) literature.
-In the following, we present a supporting result on the counterfactual recovery properties of PO-Flow under monotone structural causal models (SCMs).
-While PO-Flow does not explicitly aim to model the true SCM during training or inference, we show that if the data are generated from monotone SCMs, then PO-Flow provides recovery guarantees under certain conditions.
-We denote the SCM for an individual as $Y^{(A)}=f_{X}(A,U)$, where $U$ denotes the exogenous noise.
-The encoding function is given by $Z:=\Phi_{\theta}(Y^{(A)};\,X,A)$, which corresponds to Equation (10).
-We begin by introducing the following assumptions:
-Assumption 4.1.
-(A1) $U\perp\!\!\!\perp A$.
-(A2) The structural causal model $f_{X}(A,U)$ is monotone in $U$ for each action $A\in\{0,1\}$.
-(A3) For every $x$, the encoded latent $Z=\Phi_{\theta}(Y^{(A)};\,X,A)$ induced by the flow is statistically independent of $A$, i.e., $Z\perp\!\!\!\perp A\ |\ X=x$.
-Remark 4.2.
-In the POs framework, covariates $x$ are multivariate while the potential outcome is univariate, $y\in\mathbb{R}$.
-Therefore, (A2) is invoked for the univariate case.
-It is automatically satisfied under an additive structural causal model, $Y^{(A)}=f_{X}(A,U):=f_{X}^{*}(A)+U$.
-Under certain identifiability conditions, it also holds for non-linear models [62, 55].
-The following theoretical results hold only under monotone SCM settings.
-(A3) assumes the encoded representation $Z=\Phi_{\theta}(Y^{(A)};\,X,A)$ in PO-Flow is statistically independent of treatment $A$, though not necessarily functionally invariant to $A$.
-In the infinite-data limit with exact training, the continuous normalizing flow (Equation (8)) maps each treatment-conditioned outcome $Y$ to the same base distribution $q(Z)=\mathcal{N}(0,1)$, so $p_{\theta}(Z\mid X,A)=q(Z)$ and (A3) holds exactly.
-In finite samples, deviations may arise; we assess this empirically in Appendix D.1.
-Under the Assumption 4.1, we present a result for PO-Flow when the true SCM is monotone:
-Proposition 4.3 (Encoded $Z$ as a function of the exogenous noise $U$).
-Let Assumption 4.1 hold.
-Assume that the exogenous noise $U\sim\mathrm{Unif}[0,1]$.
-Then there exists a continuously differentiable bijection function $\psi_{X}:\;\mathcal{U}\;\longrightarrow\;\mathcal{Z}$ that is independent of the treatment assignment $A$, i.e.,
-$\small Z\;=\;\Phi_{\theta}(Y^{(A)};X,A)\;=\;\Phi_{\theta}(f_{X}(A,U);X,A)\;=\;\psi_{X}\bigl(U\bigr)$ (12)
-Remark 4.4.
-We can transform the exogenous noise $U\sim\mathrm{Unif}[0,1]$ assumed in Proposition 12 into other distributions, such as Gaussian.
-If $Z\sim\mathcal{N}(0,1)$ with CDF $F(Z)$, then $U=F(Z)\sim\mathrm{Unif}[0,1]$, and any structural assignment $f(\cdot,U)$ can be written equivalently as $\tilde{f}(\cdot,Z)=f(\cdot,F(Z))$.
-Therefore, the conclusions of Proposition 12 also hold with other noise distributions.
-Following Proposition 12, we have the following result on counterfactual recovery under monotone SCMs:
-Corollary 4.5 (Counterfactual recovery under monotone SCMs).
-Assume the conditions in Assumption 4.1 hold.
-Consider a factual sample $\bigl(Y^{(A)},X,A\bigr)$ given by the structural equation $Y^{(A)}=f_{X}\bigl(A,U\bigr)$, and let its encoded representation be $Z=\Phi_{\theta}(Y^{(A)};X,A)$.
-Apply the intervention $\mathrm{do}\bigl(A=\gamma\bigr)$.
-Assume that the true counterfactual is given by $Y_{\mathrm{CF}}^{(\gamma)}=f_{X}(\gamma,U)$.
-The decoder then recovers the counterfactual almost surely:
-$\hat{Y}_{\mathrm{CF}}^{\gamma}:=\Phi_{\theta}^{-1}\bigl(Z;\ X,\gamma\bigr)=Y_{\mathrm{CF}}^{\gamma}.$ (13)
-Remark 4.6.
-One may notice that Assumptions (A1) and (A2) mirror those in Bijective Generation Mechanisms (BGM) [43].
-BGM establishes class-level identifiability: if a learned mechanism lies in the BGM class and matches the observational distribution ($\hat{f}_{X}(A,U)\stackrel{{\scriptstyle d}}{{=}}f_{X}(A,U)$), then counterfactuals are identifiable.
-By contrast, our Prop.
-4.3 shows that PO-Flow’s encoder is bijective in $U$, placing our model within the BGM class.
-We then use an alternative constructive route (building on the intermediate step in Prop.
-4.3) to obtain Cor.
-4.5 under (A1)–(A3).
-Alternatively, one could invoke the BGM result after Prop.
-4.3 to reach Cor.
-4.5; see Appendix C.
-4.3 PO-Flow for Learning Potential Outcomes Densities
-Another advantage of PO-Flow is its ability to learn the density of the Potential Outcomes (POs).
-We define $p_{\theta,Y|X,A}(\cdot\mid x,a,z)$ as the PO density computed by PO-Flow, conditioned on $(x,a)$ and a base sample $z$.
-The explicit form of this density, mapping from latent to data space, is presented in Proposition 4.7, with the proof provided in Appendix A.
-Proposition 4.7.
-Given a noise sample $z\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
-$\log p_{\theta,Y\mid X,A}\bigl(y(0)\mid x,a,z\bigr)=\log q(z)\;+\;\int_{0}^{1}\nabla_{y}\!\cdot\,v_{\theta}\bigl(y(t),t;\,x,a\bigr)\,dt.$ (14)
-Empirically, the divergence of the velocity field can be computed much more efficiently under the Hutchinson trace estimator (e.g., [27, 58]). More details are provided in Appendix B.4.
-Unlike prior works [40, 41] that generate PO samples and approximate empirical distributions, PO-Flow provides direct likelihood evaluation.
-This allows principled uncertainty quantification, supports ranking outcomes by plausibility (Table 4), enables information-theoretic analyses such as KL divergence (Table 3), and facilitates tasks like anomaly detection via quantile-based thresholds.
-5 Experiments
-Baselines: Our baselines cover several categories: model-agnostic learners that learn the conditional distribution means (S-learner [33], T-learner [33], and DR-Learner [32]), representation-based conditional regressors (Counterfactual Regression (CFR) [50]), deep generative methods (Causal Effect Variational Autoencoder (CEVAE) [38]; GAN for Individualized Treatment Effect (GANITE) [61]), and most recent generative approaches based on diffusion (DiffPO [39]) and discrete normalizing flows (INFs [41]).
-Datasets: Counterfactual outcomes and unobserved potential outcomes (POs) are inaccessible in real-world data but can be simulated in semi-synthetic settings, as discussed in detail in Appendix B.1.
-Accordingly, we evaluate PO-Flow on synthetic benchmarks, including ACIC 2016, ACIC 2018, IHDP, and IBM.
-The code and datasets are included in the supplementary materials for anonymous submission.
-5.1 Prediction of Potential Outcomes
-Table 1: Root mean squared error (RMSE) of estimated potential outcomes (POs) across different methods on the ACIC 2018, IHDP, and IBM datasets. Results are averaged over 10-fold cross-validation.
-ACIC 2018 IHDP IBM
-$\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$
-PO-Flow $\mathbf{0.53_{\pm.10}}$ $0.60_{\pm.12}$ $\mathbf{0.98_{\pm.11}}$ $\mathbf{1.19_{\pm.13}}$ $\mathbf{0.90_{\pm.09}}$ $\mathbf{0.97_{\pm.11}}$
-DiffPO $0.72_{\pm.14}$ $0.79_{\pm.16}$ $1.33_{\pm.20}$ $1.40_{\pm.22}$ $1.82_{\pm.38}$ $1.85_{\pm.35}$
-INFs $0.69_{\pm.16}$ $0.78_{\pm.17}$ $1.02_{\pm.10}$ $\mathbf{1.20_{\pm.12}}$ $1.52_{\pm.24}$ $1.57_{\pm.28}$
-S-learner $\mathbf{0.54_{\pm.09}}$ $\mathbf{0.57_{\pm.11}}$ $1.31_{\pm.18}$ $1.44_{\pm.20}$ $01_{\pm.17} 1.1$$16_{\pm.19}$
-T-learner $1.57_{\pm.32}$ $1.71_{\pm.40}$ $1.45_{\pm.25}$ $1.49_{\pm.27}$ $1.89_{\pm.47}$ $1.96_{\pm.50}$
-CEVAE $0.83_{\pm.17}$ $0.85_{\pm.17}$ $1.18_{\pm.15}$ $1.36_{\pm.18}$ $0.96_{\pm.12}$ $\mathbf{0.98_{\pm.12}}$
+Algorithm 2 Counterfactual Prediction
+IV Proposed Method
+This section presents PO-Flow, a continuous normalizing flow framework that jointly supports (i) individualized potential outcome prediction, (ii) factual-conditioned counterfactual prediction, and (iii) learning potential outcome densities.
+We model outcomes using a continuous normalizing flow whose velocity field $v_{\theta}$ is conditioned on covariates $x$ and treatment $a$:
+$\frac{dy(t)}{dt}=v_{\theta}\!\bigl(y(t),t;\,x,a\bigr),\qquad t\in[0,1].$ (5)
+Given a data point $(y^{(a)},x,a)\sim p_{Y,X,A}$, we set $y_{0}=y^{(a)}$ as the observed training sample at time $t=0$ and draw a base sample $y_{1}=z\sim q(\cdot)$ at $t=1$.
+Under this conditional formulation, the training objective retains the form of conditional flow matching (4) and is given by
+$\mathcal{L}_{\mathrm{CFM}}(\theta)=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ (y^{(a)},x,a)\sim p_{Y,X,A},\\ y_{1}\sim q(\cdot)\end{subarray}}\left\|v_{\theta}(\phi_{t},t;\,x,a)-\frac{d\phi_{t}}{dt}\right\|^{2},$ (6)
+where $\phi_{t}$ and $d\phi_{t}/dt$ are defined in (2) and (3), respectively.
+Remark IV.1.
+Following prior work [22], one may estimate propensity scores $w(x)=p(A=1\mid X=x)$ and apply inverse weighting $1/w(x)$ in the loss.
+We observed limited empirical improvement and therefore include both weighted and unweighted variants in our implementation.
+IV-A Encoding and Decoding via CNFs
+We interpret the forward process as an encoding operation, denoted by $Z:=\Phi_{\theta}(Y^{(A)};\,X,A)$:
+$z:=\Phi_{\theta}(y^{(a)};\,x,a)=y^{(a)}+\int_{0}^{1}v_{\theta}(y(t),t;\,x,a)\,dt,$ (7)
+with $y(0)=y^{(a)}$.
+Conversely, the reverse process serves as a decoding operation:
+$\hat{y}_{0}:=\Phi_{\theta}^{-1}(z;\,x,a)=z-\int_{0}^{1}v_{\theta}(y(t),t;\,x,a)\,dt,$ (8)
+with $y(1)=z$.
+To generate a potential outcome $\hat{y}^{(a)}$ for an individual with $(x,a)$, we sample $z\sim q(\cdot)$ and decode it using (8). Algorithm 1 summarizes the procedure.
+Counterfactual prediction follows an encode–decode procedure.
+Given a factual sample $(y^{(a)},x,a)$, the forward process (7) encodes it into $z$ which captures the factual information, and the reverse process (8) then decodes $z$ under the counterfactual treatment 1a1-a to obtain $\hat{y}_{\mathrm{cf}}^{(1-a)}$.
+Algorithm 2 summarizes the procedure.
+V Theoretical Properties
+We characterize theoretical properties of PO-Flow for (i) factual-conditioned counterfactual recovery under monotone SCMs and (ii) likelihood evaluation for potential outcomes.
+V-A Counterfactual Recovery under Monotone SCMs
+We assume the underlying data-generating process follows a structural causal model (SCM)
+$Y^{(A)}=f_{X}(A,U),$
+where $U$ denotes exogenous noise. PO-Flow defines an encoding map $Z:=\Phi_{\theta}(Y^{(A)};X,A)$ via (7).
+Assumption V.1.
+  1. (A1)
+    $U\perp\!\!\!\perp A$.
+  2. (A2)
+    For each $A\in\{0,1\}$, the SCM $f_{X}(A,U)$ is strictly monotone in $U$.
+  3. (A3)
+    The encoded latent is conditionally treatment-invariant: $Z\perp\!\!\!\perp A\mid X$.
+Remark V.2.
+In the potential outcomes setting, $X$ is multivariate while the outcome is univariate ($y\in\mathbb{R}$), so (A2) concerns monotonicity in the scalar noise $U$.
+It holds automatically under additive SCMs $Y^{(A)}=f_{X}^{*}(A)+U$, and can also hold under certain identifiability conditions for nonlinear models [39, 34].
+Remark V.3.
+Assumption (A3) posits that the encoding $Z=\Phi_{\theta}(Y^{(A)};X,A)$ is conditionally independent of treatment given $X$.
+In the infinite-data limit with exact training, the CNF maps each treatment-conditioned outcome distribution to the same base distribution $q(Z)$, yielding $p_{\theta}(Z\mid X,A)=q(Z)$ and (A3) is expected to hold.
+In finite samples, deviations may arise; we assess this empirically in Appendix D.1 of [35].
+Under Assumption V.1, we obtain the following result; see Appendix A of [35] for proofs.
+Proposition V.4 (Encoded $Z$ as a function of exogenous noise $U$).
+Let Assumptions III.1 and V.1 hold and assume $U\sim\mathrm{Unif}[0,1]$. Then there exists a continuously differentiable bijection $\psi_{X}$ independent of AA such that
+$Z=\Phi_{\theta}(Y^{(A)};X,A)=\Phi_{\theta}(f_{X}(A,U);X,A)=\psi_{X}(U).$
+Remark V.5.
+The assumption $U\sim\mathrm{Unif}[0,1]$ is without loss of generality: if $Z\sim\mathcal{N}(0,1)$ with CDF $F$, then $U=F(Z)\sim\mathrm{Unif}[0,1]$, and the SCM can be equivalently written in terms of $Z$.
+Following Proposition V.4, we have the following result on counterfactual recovery under monotone SCMs; see Appendix A of [35] for proofs.
+Corollary V.6 (Counterfactual recovery under monotone SCMs).
+Assume Assumptions III.1 and V.1 hold.
+Consider a factual sample $(Y^{(A)},X,A)$ generated by $Y^{(A)}=f_{X}(A,U)$ and its encoding $Z=\Phi_{\theta}(Y^{(A)};X,A)$.
+Under the intervention $\mathrm{do}(A=\gamma)$, the true counterfactual is $Y_{\mathrm{cf}}^{(\gamma)}=f_{X}(\gamma,U)$, and the PO-Flow decoder recovers it almost surely:
+$\hat{Y}_{\mathrm{cf}}^{(\gamma)}:=\Phi_{\theta}^{-1}(Z;X,\gamma)=Y_{\mathrm{cf}}^{(\gamma)}.$
+V-B Likelihood Evaluation for Potential Outcome Densities
+PO-Flow also enables explicit likelihood evaluation for potential outcomes via the change-of-variables formula for CNFs; see Appendix A of [35] for the proof.
+Proposition V.7.
+Given a base sample $z\sim q(\cdot)$ at $t=1$, the log-density of the decoded outcome at $t=0$ is
 […]
-GANITE $0.88_{\pm.15}$ $0.97_{\pm.15}$ $1.60_{\pm.36}$ $1.67_{\pm.34}$ $2.48_{\pm.39}$ $2.59_{\pm.47}$
-Evaluation Metrics: We evaluate the point estimates of potential outcomes (POs) using the root mean squared error (RMSE) between the estimated outcomes $\hat{y}_{i}^{(a)}$ and the true outcomes $y_{i}^{(a)}$, computed as $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}^{(a)}-\hat{y}_{i}^{(a)})^{2}}$, where $\text{RMSE}_{\text{in}}$ and $\text{RMSE}_{\text{out}}$ refer to the RMSE evaluated on the training and test sets, respectively.
-Results: Table 1 shows that PO-Flow achieves the most accurate prediction of potential outcomes on the majority of datasets.
-Besides, we also report in Table 4 the RMSE of the most likely POs selected by PO-Flow using the highest log-density (as computed in Proposition 4.7).
-5.2 Estimation of CATEs
-Conditional Average Treatment Effect (CATE): The CATE is defined as $\tau(x)=\mathbb{E}[y^{(a=1)}-y^{(a=0)}\mid X=x]$, representing the expected treatment effect conditioned on covariates $x$.
-It captures individual-level heterogeneity in causal effects.
-Evaluation Metrics: We evaluate CATE using the precision in estimation of heterogeneous effect (PEHE), defined as the mean squared error between estimated and true effects: $\epsilon_{\text{PEHE}}=\frac{1}{n}\sum_{i=1}^{n}(\hat{\tau}(x_{i})-\tau(x_{i}))^{2}$.
-Table 2 reports the root PEHE for both training and test sets.
+where $y=\Phi_{\theta}^{-1}(z;x,a)$.
+VI Experiments
+We compare PO-Flow with model-agnostic mean learners (S-, T-, and DR-learners [18, 17]), representation-based regression (CFR [30]), deep generative methods (CEVAE [21], GANITE [38]), and recent generative baselines based on diffusion (DiffPO [22]) and discrete normalizing flows (INFs [24]).
+For likelihood-based comparisons, we additionally include Gaussian-process baselines for interventional treatment effects (GP-ITE) [3, 13].
+Since counterfactuals and unobserved potential outcomes are unavailable in real-world data, we evaluate on semi-synthetic benchmarks: ACIC 2016/2018, IHDP, and IBM.
+All results are averaged over 10-fold cross-validation; in all tables, columns I and O denote in-sample (training) and out-of-sample (test) performance, respectively.
+We evaluate PO-Flow across five complementary tasks, each designed to assess a distinct capability of the proposed framework.
+These include (i) point prediction of potential outcomes, (ii) conditional average treatment effect estimation, (iii) learning potential outcome densities for likelihood-based evaluation, (iv) distributional fit of potential outcomes, and (v) factual-conditioned counterfactual outcome prediction.
+Together, these tasks reflect the core claims of PO-Flow and provide a comprehensive evaluation of its individual-level predictive, distributional, and counterfactual reasoning capabilities.
+VI-A Task 1: Potential Outcome Prediction
+We evaluate point prediction of potential outcomes using RMSE using the standard definition $\mathrm{RMSE}=[\frac{1}{n}\sum_{i=1}^{n}(y_{i}^{(a)}-\hat{y}_{i}^{(a)})^{2}]^{1/2}.$ Table I shows that PO-Flow achieves the strongest performance on the majority of datasets.
+TABLE I: RMSE of estimated potential outcomes (10-fold cross-validation). Parentheses report standard deviation. I/O denote in-/out-of-sample.
+ACIC18 IHDP IBM
+Method I O I O I O
+PO-Flow 0.53(0.10) 0.60(0.12) 0.98(0.11) 1.19(0.13) 0.90(0.09) 0.97(0.11)
+DiffPO 0.52(0.13) 0.59(0.15) 1.33(0.20) 1.40(0.22) 1.82(0.38) 1.85(0.35)
+INFs 0.69(0.16) 0.78(0.17) 1.02(0.10) 1.20(0.12) 1.52(0.24) 1.57(0.28)
+S-learner 0.54(0.09) 0.57(0.11) 1.31(0.18) 1.44(0.20) 1.01(0.17) 1.16(0.19)
+T-learner 1.57(0.32) 1.71(0.40) 1.45(0.25) 1.49(0.27) 1.89(0.47) 1.96(0.50)
+CEVAE 0.83(0.17) 0.85(0.17) 1.18(0.15) 1.36(0.18) 0.96(0.12) 0.98(0.12)
+CFR 0.94(0.10) 0.98(0.12) 1.10(0.20) 1.18(0.20) 2.09(0.45) 2.17(0.48)
+GANITE 0.88(0.15) 0.97(0.15) 1.60(0.36) 1.67(0.34) 2.48(0.39) 2.59(0.47)
+VI-B Task 2: Conditional Average Treatment Effect (CATE)
+CATE is defined as $\tau(x)=\mathbb{E}[y^{(1)}-y^{(0)}\mid X=x]$, representing the expected treatment effect conditioned on covariates $x$.
+We evaluate CATE using the precision in estimation of heterogeneous effect (PEHE) [1], defined as:
+$\epsilon_{\text{PEHE}}=\frac{1}{n}\sum_{i=1}^{n}(\hat{\tau}(x_{i})-\tau(x_{i}))^{2},$ (10)
+where $\hat{\tau}(x)=\mathbb{E}_{p_{\theta}(\cdot\mid x,1)}[Y]-\mathbb{E}_{p_{\theta}(\cdot\mid x,0)}[Y]$, approximated by Monte Carlo samples from PO-Flow.
+Table II reports the root PEHE for training and test sets.
 PO-Flow achieves the most accurate CATE estimates on most datasets.
-Table 2: Root Precision in Estimation of Heterogeneous Effect ($\sqrt{\epsilon_{\text{PEHE}}}$) across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets.
-ACIC 2016 ACIC 2018 IHDP IBM
-$\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$
-PO-Flow $\mathbf{0.76_{\pm.08}}$ $\mathbf{0.82_{\pm.09}}$ $\mathbf{0.05_{\pm.01}}$ $\mathbf{0.07_{\pm.02}}$ $\mathbf{0.41_{\pm.08}}$ $\mathbf{0.45_{\pm.10}}$ $\mathbf{0.04_{\pm.01}}$ $\mathbf{0.04_{\pm.01}}$
-[…]
-INFs $1.37_{\pm.28}$ $1.42_{\pm.30}$ $0.33_{\pm.11}$ $0.37_{\pm.12}$ $0.90_{\pm.27}$ $0.95_{\pm.27}$ $0.16_{\pm.08}$ $0.18_{\pm.08}$
-S-learner $3.81_{\pm.76}$ $3.87_{\pm.98}$ $\mathbf{0.06_{\pm.02}}$ $\mathbf{0.07_{\pm.03}}$ $0.92_{\pm.07}$ $93_{\pm.08} 0.0$$6_{\pm.02} 0.0$$07_{\pm.04}$
-T-learner $5.02_{\pm.71}$ $5.15_{\pm.87}$ $1.21_{\pm.47}$ $1.32_{\pm.53}$ $1.24_{\pm.08}$ $1.31_{\pm.09}$ $0.60_{\pm.14}$ $0.69_{\pm.15}$
-DR-learner $3.04_{\pm.19}$ $3.51_{\pm.23}$ $0.43_{\pm.08}$ $0.57_{\pm.10}$ $1.03_{\pm.09}$ $1.28_{\pm.15}$ $0.12_{\pm.03}$ $0.19_{\pm.05}$
-CEVAE $2.06_{\pm.61}$ $2.31_{\pm.74}$ $1.19_{\pm.15}$ $1.21_{\pm.16}$ $1.41_{\pm.10}$ $1.43_{\pm.10}$ $0.73_{\pm.15}$ $0.75_{\pm.16}$
-CFR $2.80_{\pm.58}$ $2.74_{\pm.60}$ $0.77_{\pm.18}$ $0.80_{\pm.18}$ $0.73_{\pm.04}$ $0.78_{\pm.04}$ $1.05_{\pm.27}$ $1.08_{\pm.28}$
-GANITE $3.02_{\pm.76}$ $3.11_{\pm.74}$ $0.55_{\pm.11}$ $0.60_{\pm.13}$ $1.90_{\pm.40}$ $2.40_{\pm.40}$ $0.77_{\pm.20}$ $0.80_{\pm.24}$
-5.3 Learning the Full Potential Outcome Distributions
-Beyond point estimation, PO-Flow offers two distinct advantages for modeling potential outcomes: (1) it enables individualized sampling from the conditional distribution $P(Y\mid X,A)$, generating a rich set of plausible outcomes for each individual; and (2) it supports learning the density of any generated outcome, providing deeper insights into predictive uncertainty.
-5.3.1 Learning potential outcome densities
-PO-Flow enables direct computation of potential outcome densities, as shown in Proposition 4.7.
-To evaluate the learned densities, we compute the KL divergence between the PO-Flow estimate $p_{\theta,Y|X,A}(\cdot|x,a)$ and the ground-truth density $p_{\text{true}}(\cdot|x,a)$.
-Given that the semi-synthetic datasets assume $y^{(a)}\sim\mathcal{N}(\mu(a),1)$, the individual-level KL is $\text{KL}(\hat{Y}_{X}^{(a)}\|Y_{X}^{(a)})=\mathbb{E}_{\hat{y}\sim p_{\theta,Y|X,A}}\left[\log\frac{p_{\theta,Y|X,A}(\hat{y}|x,a)}{\mathcal{N}(\hat{y};\mu(a),1)}\right]$, where $p_{\theta,Y\mid X,A}$ is computed via (14).
-Table 3: KL divergence of learned potential outcome distributions across datasets. Results are averaged over 10-fold cross-validation.
-ACIC 2016 ACIC 2018 IHDP IBM
-PO-Flow $\mathbf{0.14_{\pm.04}}$ $\mathbf{0.08_{\pm.02}}$ $\mathbf{0.09_{\pm.03}}$ $\mathbf{0.08_{\pm.03}}$
-GP-ITE $1.73_{\pm.34}$ $0.73_{\pm.19}$ $1.42_{\pm.30}$ $0.68_{\pm.17}$
-Baselines: A natural baseline with density estimation capability is Gaussian Processes for interventional treatment effects, as used in [3, 26].
-Results: PO-Flow achieves much lower KL divergence than GP baselines on the ACIC 2018, IHDP, and IBM datasets, indicating more accurate density estimation of potential outcomes.
-This property enables PO-Flow to identify most likely potential outcomes based on their log-densities.
-Table 4 reports the RMSE of POs selected by PO-Flow with the highest log-density, illustrating one application of this property.
-Table 4: Root mean squared error (RMSE) of estimated potential outcomes (POs) across different methods on the ACIC 2018, IHDP, and IBM datasets.
-For PO-Flow, results are based on the most likely PO selected from 100 simulated outcomes using log-density.
-ACIC 2018 IHDP IBM
-$\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$
-PO-Flow $\mathbf{0.42_{\pm.07}}$ $\mathbf{0.54_{\pm.10}}$ $\mathbf{0.96_{\pm.09}}$ $\mathbf{1.05_{\pm.10}}$ $\mathbf{0.84_{\pm.06}}$ $\mathbf{0.90_{\pm.08}}$
-DiffPO $0.72_{\pm.14}$ $0.79_{\pm.16}$ $1.33_{\pm.20}$ $1.40_{\pm.22}$ $1.82_{\pm.38}$ $1.85_{\pm.35}$
-INFs $0.69_{\pm.16}$ $0.78_{\pm.17}$ $1.02_{\pm.10}$ $1.20_{\pm.12}$ $1.52_{\pm.24}$ $1.57_{\pm.28}$
-S-learner $0.54_{\pm.09}$ $0.57_{\pm.11}$ $1.31_{\pm.18}$ $1.44_{\pm.20}$ $01_{\pm.17} 1.1$$16_{\pm.19}$
-T-learner $1.57_{\pm.32}$ $1.71_{\pm.40}$ $1.45_{\pm.25}$ $1.49_{\pm.27}$ $1.89_{\pm.47}$ $1.96_{\pm.50}$
-CEVAE $0.83_{\pm.17}$ $0.85_{\pm.17}$ $1.18_{\pm.15}$ $1.36_{\pm.18}$ $0.96_{\pm.12}$ $0.98_{\pm.12}$
-CFR $0.94_{\pm.10}$ $0.98_{\pm.12}$ $10_{\pm{.20}} 1.1$$18_{\pm{.20}}$ $2.09_{\pm.45}$ $2.17_{\pm.48}$
-GANITE $0.88_{\pm.15}$ $0.97_{\pm.15}$ $1.60_{\pm.36}$ $1.67_{\pm.34}$ $2.48_{\pm.39}$ $2.59_{\pm.47}$
-Compared with the averaged RMSE over all POs generated by PO-Flow in Table 1, selecting the most likely POs based on log-density yields smaller RMSEs, providing more confident and closer estimates.
-5.3.2 Empirical distribution estimations
-Evaluation Metrics: To enable comparison of estimated potential outcome distributions across methods, we include empirical distribution estimates for all baselines.
-These are evaluated using the Wasserstein-1 distance, defined as: $W_{1}(p_{\theta,Y|X,A}(y|x,a),p_{Y|X,A}(y|x,a))=\inf_{\gamma\in\Gamma(p_{\theta},p)}\mathbb{E}_{(\hat{y},y\sim\gamma)}[\|\hat{y}-y\|_{1}]$.
-The results are shown in Table 5.
-Table 5: In-sample and out-of-sample Wasserstein distances across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets. Results are averaged over 10-fold cross-validation.
-ACIC 2016 ACIC 2018 IHDP IBM
-$\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$
-PO-Flow $\mathbf{0.36_{\pm.11}}$ $\mathbf{0.42_{\pm.12}}$ $\mathbf{0.05_{\pm.03}}$ $\mathbf{0.12_{\pm.04}}$ $\mathbf{0.30_{\pm.05}}$ $\mathbf{0.41_{\pm.06}}$ $\mathbf{0.33_{\pm.07}}$ $\mathbf{0.44_{\pm.08}}$
-[…]
-INFs $0.59_{\pm.13}$ $67_{\pm.14} 0.0$$08_{\pm.03}$ $0.16_{\pm.05}$ $0.32_{\pm.08}$ $\mathbf{0.40_{\pm.09}}$ $0.36_{\pm.08}$ $\mathbf{0.45_{\pm.09}}$
-S-learner $1.01_{\pm.13}$ $1.02_{\pm.20}$ $0.35_{\pm.13}$ $0.49_{\pm.20}$ $0.71_{\pm.14}$ $0.85_{\pm.19}$ $0.46_{\pm.11}$ $0.55_{\pm.14}$
-T-learner $1.22_{\pm.15}$ $1.48_{\pm.18}$ $1.04_{\pm.13}$ $1.00_{\pm.30}$ $0.83_{\pm.20}$ $0.92_{\pm.28}$ $1.21_{\pm.23}$ $1.67_{\pm.32}$
-CEVAE $0.62_{\pm.17}$ $0.70_{\pm.21}$ $0.88_{\pm.18}$ $1.02_{\pm.23}$ $0.65_{\pm.14}$ $0.76_{\pm.20}$ $0.38_{\pm.07}$ $\mathbf{0.45_{\pm.07}}$
-CFR $0.68_{\pm.13}$ $0.73_{\pm.15}$ $0.88_{\pm.12}$ $0.90_{\pm.15}$ $0.50_{\pm.17}$ $0.56_{\pm.19}$ $0.92_{\pm.20}$ $0.93_{\pm.21}$
-GANITE $0.74_{\pm.20}$ $0.86_{\pm.23}$ $0.89_{\pm.24}$ $1.18_{\pm.25}$ $0.71_{\pm.23}$ $0.82_{\pm.29}$ $1.09_{\pm.27}$ $1.34_{\pm.30}$
-Understanding the distribution of potential outcomes (POs) is essential for uncertainty-aware treatment decisions, as outcomes for the same individual may vary substantially across treatments.
-PO-Flow provides not only point estimates but also full distributional insight, enabling practitioners to assess whether outcomes fall within desired ranges and to evaluate their likelihood (Proposition 4.7).
-5.4 Counterfactual Outcome Estimation
-Evaluation Metrics: We assess counterfactual predictions using the RMSE between model-estimated $\hat{y}_{\text{cf}}^{(1-a)}$ and ground-truth $y_{\text{cf}}^{(1-a)}$. Results are in Table 6.
-Baselines: To our knowledge, very few existing methods are specifically designed for counterfactual outcome prediction conditioned on factuals in general settings—without explicitly modeling structural causal models or relying on temporal information.
-While methods like CFR include “counterfactual” in name, they target unobserved potential outcomes rather than factual-conditioned ones, though we still include them as valuable baselines.
-We also adapted DiffPO and GANITE for this task, as detailed in Appendix B.5.
-Table 6: Root mean squared error (RMSE) of estimated counterfactual outcomes across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets.
-Results are averaged over 10-fold cross-validation.
-ACIC 2016 ACIC 2018 IHDP IBM
-$\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$
-PO-Flow $\mathbf{1.50_{\pm.15}}$ $\mathbf{1.53_{\pm.17}}$ $\mathbf{0.04_{\pm.01}}$ $\mathbf{0.05_{\pm.02}}$ $\mathbf{1.63_{\pm.09}}$ $1.69_{\pm.10}$ $\mathbf{0.04_{\pm.01}}$ $\mathbf{0.04_{\pm.02}}$
-DiffPO $1.77_{\pm.20}$ $1.78_{\pm.20}$ $0.62_{\pm.14}$ $0.65_{\pm.15}$ $1.81_{\pm.22}$ $1.85_{\pm.23}$ $0.62_{\pm.13}$ $0.64_{\pm.16}$
-CFR $15_{\pm.28} 3.3$$38_{\pm.30}$ $0.99_{\pm.05}$ $1.01_{\pm.07}$ $\mathbf{1.63_{\pm.21}}$ $\mathbf{1.65_{\pm.25}}$ $0.86_{\pm.18}$ $0.87_{\pm.16}$
-GANITE $3.41_{\pm.52}$ $3.67_{\pm.55}$ $0.68_{\pm.13}$ $0.70_{\pm.14}$ $2.48_{\pm.53}$ $2.72_{\pm.59}$ $1.84_{\pm.30}$ $2.02_{\pm.37}$
-5.5 Discussions and Comparisons
-Training efficiency: PO-Flow exhibits high training efficiency, converging in approximately 300 gradient steps across multiple causal datasets (Appendix D.3), and demonstrates notable advantages in both training and sampling time compared to other baselines (Table 9).
-Density estimation for principled uncertainty quantification: Our continuous normalizing flow (CNF) backbone enables density learning of potential outcomes, as detailed in Section 5.3.1.
-This property allows us to generate multiple potential outcomes for each individual and to identify the most likely ones based on their log-densities.
-The empirical evaluation of this procedure is presented in Table 4.
-Counterfactual: We take a step further by exploring counterfactual outcomes, enabled by the smooth PO-Flow mapping that preserves factual information during encoding.
-Under (A1)–(A3), we include a supporting recovery result for this architecture (Proposition 12; Corollary 4.5).
-It is among the first to explore counterfactual outcomes in the potential outcomes framework without relying on temporal data or explicitly modeling structural causal models.
-6 Conclusions
-We introduced PO-Flow, a continuous normalizing flow framework for modeling potential and counterfactual outcomes in causal inference.
-PO-Flow provides a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction in potential outcomes—without explicitly modeling structural causal models (SCMs) or relying on auxiliary information.
-To our knowledge, this constitutes one of the earliest attempts to address counterfactual outcome prediction in the potential outcomes literature, a crucial yet underexplored direction.
-Theoretically, we include a supporting recovery result for the proposed architecture (Corollary 4.5).
-Empirically, PO-Flow attains state-of-the-art performance on benchmark datasets across standard causal tasks, while additionally supporting counterfactual outcome prediction and log-density estimation for principled uncertainty quantification.
-In addition, PO-Flow is both computationally efficient and robust compared to modern baselines.
-These results highlight PO-Flow as a unified and scalable solution for potential-outcomes–based causal tasks, supporting reliable decision-making in domains such as healthcare.
+TABLE II: Measuring precision in estimation of the heterogeneous effect using $\sqrt{\epsilon_{\text{PEHE}}}$; Standard deviations omitted and they are reasonably small. I/O denote in-/out-of-sample.
+ACIC16 ACIC18 IHDP IBM
+Method I O I O I O I O
+PO-Flow 0.76 0.82 0.05 0.07 0.41 0.45 0.04 0.04
+DiffPO 0.91 1.12 0.07 0.08 0.79 0.84 0.08 0.09
+INFs 1.37 1.42 0.33 0.37 0.90 0.95 0.16 0.18
+S-learner 3.81 3.87 0.06 0.07 0.92 0.93 0.06 0.07
+T-learner 5.02 5.15 1.21 1.32 1.24 1.31 0.60 0.69
+DR-learner 3.04 3.51 0.43 0.57 1.03 1.28 0.12 0.19
+CEVAE 2.06 2.31 1.19 1.21 1.41 1.43 0.73 0.75
+CFR 2.80 2.74 0.77 0.80 0.73 0.78 1.05 1.08
+GANITE 3.02 3.11 0.55 0.60 1.90 2.40 0.77 0.80
+VI-C Task 3: Learning Potential Outcome Densities
+PO-Flow supports explicit likelihood evaluation of potential outcomes via Proposition V.7.
+We evaluate density accuracy using the KL divergence between the learned density $p_{\theta,Y\mid X,A}(\cdot\mid x,a)$ and the ground-truth density $p_{\mathrm{true}}(\cdot\mid x,a)$.
+For the semi-synthetic benchmarks, $Y^{(a)}\sim\mathcal{N}(\mu_{X}(a),1)$, and we compute
+$\mathrm{KL}\!\left(\hat{Y}^{(a)}\,\|\,Y^{(a)}\right)=\mathbb{E}_{\hat{y}\sim p_{\theta,Y\mid X,A}}\left[\log\frac{p_{\theta,Y\mid X,A}(\hat{y}\mid x,a)}{\mathcal{N}(\hat{y};\mu_{X}(a),1)}\right],$
+where $p_{\theta,Y\mid X,A}$ is evaluated via (9). As shown in Table III, PO-Flow achieves lower KL divergence than GP-ITE and INF on ACIC 2018, IHDP, and IBM.
+TABLE III: KL divergence of learned potential outcome densities (10-fold cross-validation). Parentheses indicate standard deviation.
+Method ACIC16 ACIC18 IHDP IBM
+PO-Flow 0.14(0.04) 0.08(0.02) 0.09(0.03) 0.08(0.03)
+INF 0.22(0.09) 0.15(0.06) 0.14(0.07) 0.10(0.04)
+GP-ITE 1.73(0.34) 0.73(0.19) 1.42(0.30) 0.68(0.17)
+Likelihood evaluation also enables selecting the most likely potential outcomes under the learned density.
+Table IV reports the RMSE of PO-Flow (MAP), where predictions are chosen by maximizing the log-density.
+TABLE IV: RMSE for PO-Flow variants (10-fold cross-validation). I/O denote in-/out-of-sample.
+ACIC18 IHDP IBM
+Method I O I O I O
+PO-Flow 0.53 0.60 0.98 1.19 0.90 0.97
+PO-Flow (MAP) 0.42 0.54 0.96 1.05 0.84 0.90
+VI-D Task 4: Distributional Fit of Potential Outcomes
+To compare estimated potential outcome distributions across methods, we evaluate empirical distribution fit using the Wasserstein-1 distance, following the standard definition
+$\begin{split}&W_{1}\!\left(p_{\theta,Y\mid X,A}(\cdot\mid x,a),\,p_{Y\mid X,A}(\cdot\mid x,a)\right)\\ =&\inf_{\gamma\in\Gamma(p_{\theta},p)}\mathbb{E}_{(\hat{y},y)\sim\gamma}\bigl[\|\hat{y}-y\|_{1}\bigr],\end{split}$
+where $\Gamma$ denotes the “coupling”: the joint distribution between a pair of random variables, whose marginal distributions match $p_{\theta}$ and $p$ respectively.
+Results are reported in Table V.
+PO-Flow consistently achieves strong performance, indicating accurate recovery of the full potential outcome distributions.
+TABLE V: Wasserstein distances (standard deviations omitted, and they are reasonably small). I/O denote in-/out-of-sample.
+ACIC16 ACIC18 IHDP IBM
+Method I O I O I O I O
+PO-Flow 0.36 0.42 0.05 0.12 0.30 0.41 0.33 0.44
+DiffPO 0.47 0.51 0.04 0.13 0.80 0.88 0.69 0.75
+INFs 0.59 0.67 0.08 0.16 0.32 0.40 0.36 0.45
+S-learner 1.01 1.02 0.35 0.49 0.71 0.85 0.46 0.55
+T-learner 1.22 1.48 1.04 1.00 0.83 0.92 1.21 1.67
+CEVAE 0.62 0.70 0.88 1.02 0.65 0.76 0.38 0.45
+CFR 0.68 0.73 0.88 0.90 0.50 0.56 0.92 0.93
+GANITE 0.74 0.86 0.89 1.18 0.71 0.82 1.09 1.34
+VI-E Task 5: Factual-Conditioned Counterfactual Outcome Estimation
+We evaluate factual-conditioned counterfactual prediction using RMSE between the predicted counterfactual $\hat{y}_{\mathrm{cf}}^{(1-a)}$ and the ground-truth $y_{\mathrm{cf}}^{(1-a)}$.
+Results are shown in Table VI, which shows the proposed method achieves the best performance.
+While DiffPO and GANITE do not natively support counterfactual queries, we adapt them for evaluation.
+Details are provided in Appendix B.5 of [35].
+TABLE VI: RMSE of estimated counterfactual outcomes (10-fold cross-validation). DiffPO and GANITE do not natively support this query and are adapted by us for evaluation. I/O denote in-/out-of-sample.
+ACIC16 ACIC18 IHDP IBM
+Method I O I O I O I O
+PO-Flow 1.50 1.53 0.04 0.05 1.63 1.69 0.04 0.04
+DiffPO 1.77 1.78 0.62 0.65 1.81 1.85 0.62 0.64
+GANITE 3.41 3.67 0.68 0.70 2.48 2.72 1.84 2.02
+Compared to DiffPO and INFs: PO-Flow supports factual-conditioned counterfactual prediction beyond standard potential-outcome modeling.
+While DiffPO and INFs primarily model interventional outcome distributions, PO-Flow performs counterfactual inference via an encoding-decoding procedure enabled by a smooth and information-preserving path of continuous normalizing flows.
+Although both baselines can be adapted to this task, they consistently underperform PO-Flow, likely due to INFs’ restrictive discrete-flow parameterization and DiffPO’s forward noising, which limit preservation of factual information.
+VII Conclusions
+We proposed PO-Flow, a continuous normalizing flow framework for causal inference that models potential outcomes and factual-conditioned counterfactuals within the potential outcomes framework.
+By mapping an observed factual outcome to a shared latent representation and decoding it under alternative treatments, PO-Flow enables individual-level factual–counterfactual mapping without requiring a parametric structural causal model.
+This supports individualized “what-if” analysis with likelihood-based uncertainty assessment, and the accompanying theory and experiments show that PO-Flow offers a unified and scalable approach to individualized causal reasoning.
 Acknowledgment
-D.I.
-acknowledges support from NSF (IIS-2212097), ARL (W911NF-2020-221), and ONR (N00014-23-C-1016).
-Y.X.
-acknowledges partial support from NSF DMS-2134037, CMMI-2112533, and the Coca-Cola Foundation.
-Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsor(s).
+The work of D.W. and Y.X. are supported by NSF CMMI-2112533, and the Coca-Cola Foundation.
 Appendix A Proofs
@@
 $\nabla_{\theta}\mathcal{L}_{\text{FM}}(\theta)=\nabla_{\theta}\mathcal{L}_{\text{CFM}}(\theta).$ (17)
-Proposition 12.
-(Encoded $Z$ as a function of the exogenous noise $U$) Let Assumption 4.1 hold.
+Proposition V.4.
+(Encoded $Z$ as a function of the exogenous noise $U$) Let Assumptions III.1 and V.1 hold.
 Assume that the exogenous noise $U\sim\mathrm{Unif}[0,1]$.
@@
 $m_{X,A}(U)\;:=Z=\Phi_{\theta}(Y^{(A)};\,X,A)=\;\Phi_{\theta}\bigl(f_{X}(A,U);\ X,A\bigr),$
-as a function that may depend on the action $A$. Therefore, our target becomes proving that $m_{X,A}(U)$ is functionally invariant to $A$.
-By (A3) in Assumption 4.1, we have $Z\perp\!\!\!\perp A\ |\ X$. Therefore,
-$p_{\theta,Z\mid X,A}(z)=p_{\theta,Z}(z).$ (23)
+as a function that may depend on the treatment $A$. Therefore, our target becomes proving that $m_{X,A}(U)$ is functionally invariant to $A$.
+By (A3) in Assumption V.1, we have $Z\perp\!\!\!\perp A\ |\ X$. Therefore,
+$p_{\theta,Z\mid X,A}(z)=p_{\theta,Z|X}(z).$ (19)
 In continuous normalizing flows, $\Phi_{\theta}$ is invertible with respect to $Y^{(A)}$ due to the invertibility of the underlying neural ODE.
@@
 where the last equation follows from the uniform distribution of $U$ and the fact that $p_{Z}>0$.
-Then based on (23), we have:
+Next, by combining (19) and (20), we have:
 $p_{\theta,Z|X,A=1}(z)=\frac{d}{dz}m_{X,1}^{-1}\!\bigl(z\bigr)=\frac{d}{dz}m_{X,0}^{-1}\!\bigl(z\bigr)=p_{\theta,Z|X,A=0}(z).$ (21)
@@
 is a function that only depends on the exogenous noise $U$ (and covariates $X$), and is invariant to the treatment $A$.
-Corollary 4.5 (Counterfactual recovery) Assume the conditions in Assumption 4.1 hold.
+Corollary V.6 (Counterfactual recovery) Assume the conditions in Assumptions III.1 and V.1 hold.
 Consider a factual sample $\bigl(Y^{(A)},X,A\bigr)$ given by the structural equation $Y^{(A)}=f_{X}\bigl(A,U\bigr)$, and let its encoded representation be $Z=\Phi_{\theta}(Y^{(A)};X,A)$.
@@
 $\Phi_{\theta}^{-1}(\Phi_{\theta}(Y;\,X,A);X,A)=Y.$ (29)
-From Proposition 4.3, we have $Z=\;\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr)\;=\Phi_{\theta}(f_{X}(A,U);\ X,A\bigr)\;=\;\psi_{X}\bigl(U\bigr)$ does not depend on $A$.
+From Proposition V.4, we have $Z=\;\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr)\;=\Phi_{\theta}(f_{X}(A,U);\ X,A\bigr)\;=\;\psi_{X}\bigl(U\bigr)$ does not depend on $A$.
 Therefore, fixing $U$ as the same and plugging in $\mathrm{do}\bigl(A=\gamma\bigr)$, we have:
@@
 $\Phi_{\theta}^{-1}(\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr);\ X,\gamma)=\Phi_{\theta}^{-1}(\Phi_{\theta}(Y_{\mathrm{CF}}^{\gamma};\ X,\gamma);\ X,\gamma).$ (32)
-Noting that the L.H.S. of (36) is exactly the counterfactual procedures (encoding-decoding) of PO-Flow, and by (33), the R.H.S. of (36) equals to $Y_{\mathrm{CF}}^{\gamma}$. Therefore,
+Noting that the L.H.S. of (32) is exactly the counterfactual procedures (encoding-decoding) of PO-Flow, and by (29), the R.H.S. of (32) equals $Y_{\mathrm{CF}}^{\gamma}$. Therefore,
 $\hat{Y}_{\mathrm{CF}}^{\gamma}:=\Phi_{\theta}^{-1}\bigl(Z;\ X,\gamma\bigr)=Y_{\mathrm{CF}}^{\gamma}.$ (33)
-Proposition 4.7. Given a noise sample $z\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
-$\log p_{\theta,Y\mid X,A}\bigl(y(0)\mid x,a,z\bigr)=\log q(z)\;+\;\int_{0}^{1}\nabla_{y}\!\cdot\,v_{\theta}\bigl(y(t),t;\,x,a\bigr)\,dt.$
+Proposition V.7. Given a noise sample $z\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
+$\log p_{\theta,Y\mid X,A}\bigl(y\mid x,a,z\bigr)=\log q(z)\;+\;\int_{0}^{1}\nabla_{y}\!\cdot\,v_{\theta}\bigl(y(t),t;\,x,a\bigr)\,dt.$
 Proof:
@@
 Appendix B Experimental Details
-B.1 Synthetic Causal Datasets
+B-A Synthetic Causal Datasets
 Counterfactual outcomes are unobservable in real-world data but can be simulated using synthetic datasets.
 In our experiments, we evaluate on widely used benchmarks, including the Atlantic Causal Inference Conference datasets (ACIC 2016 and ACIC 2018), the Infant Health and Development Program (IHDP) [11, 21], and the IBM Causal Inference Benchmark [31].
-The ACIC 2016 dataset contains 4,802 samples with 82 covariates each ($d_{X}=82$), ACIC 2018 has 1,000 samples with 177 covariates, IHDP includes 747 samples with 25 covariates, and the IBM dataset consists of 1,000 samples with 177 covariates.
+The ACIC 2016 dataset contains 4,802 samples with 82 covariates each ($d_{X}=82$), ACIC 2018 has 1,000 samples with 177 covariates, The IHDP dataset includes 747 samples with 25 covariates, and the IBM dataset consists of 1,000 samples with 177 covariates.
 Each dataset provides three synthetic values per individual: the factual outcome $y^{(a)}$, the assigned treatment $a$, and the mean potential outcomes under both treatments, $\mu_{0}$ (for $a=0$) and $\mu_{1}$ (for $a=1$).
@@
 where $\beta$, $W$, and $\omega$ are parameters defined by the specific simulation setting.
-B.2 Neural Network Structure and Implementation Information
+B-B Neural Network Structure and Implementation Information
 The neural network used in experiments on causal datasets takes as input $y(t)$, $x$, $t$, and $a$.
@@
 For example, training on the ACIC 2018 dataset with a batch size of 256 typically takes 2–4 seconds per epoch.
-Log-density evaluation, as discussed in Section 4.3, involves divergence integration and requires significantly more time than other test metrics.
-B.3 Runge-Kutta numerical integration method
+Log-density evaluation involves divergence integration and requires significantly more time than other test metrics.
+B-C Runge-Kutta numerical integration method
 We apply the classical 4th-order Runge-Kutta method to solve the Neural ODE in the forward process of PO-Flow.
@@
 with $y(0)=y^{(a)}$ and $y(1)=z$. This produces a numerically integrated trajectory from the observed sample to its latent representation.
-B.4 Hutchinson trace estimator
+B-D Hutchinson trace estimator
 The computations of the log-density of the POs involve the calculation of $\nabla\cdot v_{k}(x,t)$, i.e., the divergence of the velocity field represented by a neural network.
@@
 […]
-B.5 Adapted Implementation Details for DiffPO and GANITE Counterfactuals
+B-E Adapted Implementation Details for DiffPO and GANITE Counterfactuals
 DiffPO was not originally designed for counterfactual tasks.
@@
 Bijective Generation Mechanisms (BGM).
-The BGM framework [43] shows that if the true structural mechanism $f$ is in the BGM class (bijective/strictly monotone in the exogenous noise), then any learned mechanism $\hat{f}$ that (i) is also in the BGM class and (ii) matches the observed distribution—i.e., for the same parents value $X$ and action $A$,
+The BGM framework [25] shows that if the true structural mechanism $f$ is in the BGM class (bijective/strictly monotone in the exogenous noise), then any learned mechanism $\hat{f}$ that (i) is also in the BGM class and (ii) matches the observed distribution—i.e., for the same parents value $X$ and treatment $A$,
 $\hat{f}_{X}(A,U)\ \stackrel{{\scriptstyle d}}{{=}}\ f_{X}(A,U)\quad(\text{equivalently }(\hat{f}(X,\cdot))_{\#}P_{U}=(f(X,\cdot))_{\#}P_{U}),$
@@
 Our findings are specific to a particular methodology and architecture, as this is a methodology paper where we analyze continuous normalizing flows (CNFs).
-With the additional assumption (A3) specific to PO-Flow, Proposition 4.3 shows that the PO-Flow encoded latent $Z$ is bijective in the exogenous noise $U$.
-Consequently, Corollary 4.5 proves that the encode–decode procedure recovers the true counterfactual.
+With the additional assumption (A3) specific to PO-Flow, Proposition V.4 shows that the PO-Flow encoded latent $Z$ is bijective in the exogenous noise $U$.
+Consequently, Corollary V.6 proves that the encode–decode procedure recovers the true counterfactual.
 Relationships.
-Because of Proposition 4.3 (enabled by the CNF), PO-Flow implements a bijective-in-noise mechanism and is thus comparable to BGM in the sense that $\hat{f}$ is bijective in $U$.
-However, BGM additionally requires observational distribution matching to obtain class-level identifiability, whereas the proofs of Proposition 4.3 and Corollary 4.5 do not assume such matching.
-Alternative route. The counterfactual recovery result (Corollary 4.5) can alternatively be established by imposing an additional assumption:
+Because of Proposition V.4 (enabled by the CNF), PO-Flow implements a bijective-in-noise mechanism and is thus comparable to BGM in the sense that $\hat{f}$ is bijective in $U$.
+However, BGM additionally requires observational distribution matching to obtain class-level identifiability, whereas the proofs of Proposition V.4 and Corollary V.6 do not assume such matching.
+Alternative route. The counterfactual recovery result (Corollary V.6) can alternatively be established by imposing an additional assumption:
 (A4) Observational matching. For each $(x,a)$, the PO-Flow–induced observational law matches the true one, i.e.,
 $(\hat{f}_{X}(a,\cdot))_{\#}P_{U}\;=\;(f_{X}(a,\cdot))_{\#}P_{U}.$
-Under (A1)–(A3), Proposition 4.3 establishes a continuously differentiable bijection $\psi_{X}:U\to Z$ (independent of $A$), so the induced mechanism $\hat{f}_{X}$ implemented by PO-Flow is bijective in the exogenous noise, i.e., PO-Flow lies in the BGM class.
+Under (A1)–(A3), Proposition V.4 establishes a continuously differentiable bijection $\psi_{X}:U\to Z$ (independent of $A$), so the induced mechanism $\hat{f}_{X}$ implemented by PO-Flow is bijective in the exogenous noise, i.e., PO-Flow lies in the BGM class.
 With the additional assumption (A4), we can therefore directly invoke the counterfactual recovery result of the BGM framework [25].
 Appendix D Additional Experimental Results
-D.1 Empirical Validation of Assumption (A3)
+D-A Empirical Validation of Assumption (A3)
 We test (A3): $p_{\theta}(Z\mid X,A)=q(Z)=\mathcal{N}(0,1)$, using a two-sample Maximum Mean Discrepancy (MMD).
@@
 The test is conducted on each dataset considered in the paper, and the aggregated results are summarized in the table below.
-Table 7: Empirical validation of Assumption (A3), $p_{\theta}(Z\mid X,A)=q(Z)=\mathcal{N}(0,1)$.
+TABLE VII: Empirical validation of Assumption (A3), $p_{\theta}(Z\mid X,A)=q(Z)=\mathcal{N}(0,1)$.
 We report two-sample MMD values comparing the empirical joint distribution $\{(z_{i},x_{i},a_{i})\}$ with synthetic samples $\{(z_{i}^{\prime},x_{i},a_{i})\}$ across datasets.
@@
 True $4.1\times 10^{-2}$ $5.9\times 10^{-3}$ $2.5\times 10^{-2}$ $5.3\times 10^{-2}$
-As shown in Table 7, PO-Flow achieves small two-sample joint MMD values, which are comparable to those obtained from the ground-truth independent samples $\{(z_{i}^{\prime},x_{i},a_{i})\}$ and $\{(z_{i}^{\prime\prime},x_{i},a_{i})\}$.
+As shown in Table VII, PO-Flow achieves small two-sample joint MMD values, which are comparable to those obtained from the ground-truth independent samples $\{(z_{i}^{\prime},x_{i},a_{i})\}$ and $\{(z_{i}^{\prime\prime},x_{i},a_{i})\}$.
 This provides empirical support for the validity of Assumption (A3).
-D.2 Identifying the Most Likely Potential Outcomes via Log-Density
-As shown in Proposition 4.7 and Section 5.3.1, we can naturally select the most likely potential outcome based on log-density.
-Table 8 reports the RMSE of PO-Flow when using the most likely PO selected by the algorithm.
+D-B Identifying the Most Likely Potential Outcomes via Log-Density
+As shown in Proposition V.7, we can naturally select the most likely potential outcome based on log-density.
+Table VIII reports the RMSE of PO-Flow when using the most likely PO selected by the algorithm.
 Specifically, for each sample we simulate 100 POs with PO-Flow, select the one with the highest log-density, and compute the results accordingly.
-Table 8: Root mean squared error (RMSE) of estimated potential outcomes (POs) across different methods on the ACIC 2018, IHDP, and IBM datasets.
+TABLE VIII: Root mean squared error (RMSE) of estimated potential outcomes (POs) across different methods on the ACIC 2018, IHDP, and IBM datasets.
 For PO-Flow, results are based on the most likely PO selected from 100 simulated outcomes using log-density.
@@
 GANITE $0.88_{\pm.15}$ $0.97_{\pm.15}$ $1.60_{\pm.36}$ $1.67_{\pm.34}$ $2.48_{\pm.39}$ $2.59_{\pm.47}$
-Compared with the averaged RMSE over all POs generated by PO-Flow in Table 1, selecting the most likely POs based on log-density yields smaller RMSEs, providing more confident and closer estimates.
-D.3 Convergence and Computational Costs
+Compared with the averaged RMSE over all POs generated by PO-Flow in Table I, selecting the most likely POs based on log-density yields smaller RMSEs, providing more confident and closer estimates.
+D-C Convergence and Computational Costs
 (a) Convergence of MSE and RMSE for predicted potential outcomes over the training iterations on the ACIC 2018, IHDP, and IBM datasets.
@@
 This highlights the model’s strong training efficiency.
-Table 9: Comparison of model size, training time untill convergence, and sampling cost (for 500 samples).
+TABLE IX: Comparison of model size, training time until convergence, and sampling cost (for 500 samples).
 Methods # Params Training Time Sampling
@@
 CEVAE $97,864$ 4.644.64min $0.47$s
-In addition, Table 9 compares model size, training time to convergence, and sampling time (for 500 samples) across PO-Flow and other deep generative baselines.
+In addition, Table IX compares model size, training time to convergence, and sampling time (for 500 samples) across PO-Flow and other deep generative baselines.
 All experiments were conducted on a single A100 GPU.